### v2

# Benford behavior resulting from stick and box fragmentation processes

Benford behavior resulting from stick and box fragmentation processes
Bruce Fang fangbaojun2002@gmail.com Department of Mathematics and Statistics, Williams College, Williamstown, MA 01267  and  Steven J. Miller sjm1@williams.edu, Steven.Miller.MC.96@aya.yale.edu Department of Mathematics and Statistics, Williams College, Williamstown, MA 01267
(Date: August 19, 2025)
Benford’s law is the statement that in many real-world data set, the probability of having digit $d$ in base $B$, where $1\leq d\leq B$, as the first digit is $\log_{B}\left((d+1)/d\right)$. We sometimes refer to this as weak Benford behavior, and we say that a data set exhibits strong Benford behavior in base $B$ if the probability of having significand at most $s$, where $s\in[1,B)$, is $\log_{B}(s)$. We examine Benford behaviors in two different probabilistic model: stick and box fragmentation. Building on the joint work of [B–] on the 1-dimensional stick fragmentation model, we employ combinatorial identities on multinomial coefficients to reduce the high dimensional stick fragmentation model to the 1-dimensional model and provide a necessary and sufficient condition for the lengths of the stick fragments to converge to strong Benford behavior. Then we answer a conjecture posed by [BDMMM] on the high dimensional box fragmentation model. Using tools from order statistics, we prove that under some mild conditions, faces of any arbitrary dimension of the box have total volume converging to strong Benford’s behavior.
1. Introduction
At the beginning of the $20$th century, astronomer and mathematician Simon Newcomb [Ne] noticed an unusual pattern in the logarithmic tables he used at work. The early pages of the books were far more worn than the later ones, suggesting that numbers starting with smaller digits were consulted more often. From this, Newcomb inferred that people had a tendency to encounter numbers beginning with $1$ more frequently. Specifically, he found that $1$ appeared as the leading digit about $30\%$ of the time, $2$ about $17\%$, with the frequency decreasing for larger digits. Although he formulated a mathematical explanation for this curious phenomenon, his discovery initially went largely overlooked.
It took another 57 years after Newcomb’s discovery for physicist Frank Benford to make the exact same observation as Newcomb: the first pages of logarithmic tables were used far more than others. He formulated this law as follows.
Definition 1.1.
[Ben, Page 554] Data exhibits (weak) Benford behavior base $B$ if the frequency $F_{d}$ of leading digit $d$ is
$F_{d}\ =\ \log_{B}\left(\frac{d+1}{d}\right).$ (1.1)
Nowadays, Benford’s Law is used in detecting many different forms of fraud, and its prevalence in the world fascinates not only mathematicians, but many other scientists as well (to learn more about Benford’s law and its many applications, we recommend [BeHi, Nig, Mil1, NiMi] to name a few).
In 1986, Lemons [Lem] proposed using Benford’s law to analyze the partitioning of a conserved quantity. Since then, driven by the potential application to nuclear fragmentation, mathematicians and physicists have taken an interest in the Benford behavior of various fragmentation processes. Among these processes of interest is stick fragmentation. In a 1-dimensional stick fragmentation model, one begins with a stick of length $L$. Draw $p_{1}$ from a probability distribution on $(0,1)$. This fragments the stick into two sub-sticks of lengths $p_{1}L$ and $(1-p_{1})L$. For each sub-stick, draw two probabilities $p_{2}$ and $p_{3}$ from the same distribution. We perform this fragmentation process for a total of $N$ stages, where at each stage we fragment all the sub-sticks created during the previous stage according to probabilities. Hence, by the end of $N$ stages, there are $2^{N}$ sticks. Of particular interest is whether or not this fragmentation process converges to Benford’s behavior.
1.1. Previous Work on Fragmentation
An important definition when studying a more precise statistical version of Benford behavior is the notion of the significand of a real number, i.e., its leading digits in scientific notation.
Definition 1.2.
Given a positive real number $x$, we say that its significand base $B>1$, denoted $S_{B}(x)$, is the unique real number $S_{B}(x)\in[1,B)$ such that $x=S_{B}(x)\cdot B^{k}$, where $k$ is an integer.
As is common practice with these techniques involving proofs of Benford behavior, we define a more general version of Benford behavior that allows for processes that do not necessarily exhibit Benford behavior at first but converge to it in the limit.
Definition 1.3.
We say that a sequence of random variables $X^{(n)}$ converges to strong Benford behavior base $B$ if and only if
$\mathbb{P}(S_{B}(X^{(n)})\leq s)\to\log_{B}(s)$ (1.2)
for all $s\in[1,B]$. Notice by compactness that this implies uniform convergence of (1.2).
An equivalent formulation to the above is the Uniform Distribution Characterization, which is especially suited for investigation of products of random variables; see [Dia] for a proof.
Proposition 1.4 (Uniform Distribution Characterization).
[Dia] A sequence of random variables $X^{(n)}$ converges to strong Benford behavior in base $B$ if and only if
$\mathbb{P}\left(\log_{B}(X^{(n)})\textup{ mod }1\leq t\right)\rightarrow t,$ (1.3)
for all $t\in[0,1]$. If (1.4) is satisfied, then we say that $\log_{B}(X^{(n)})$ converges to being equidistributed mod 1.
We may now state the previous results on stick fragmentation. Becker et al. [B–] proved a theorem regarding fixed proportion 1-dimensional stick fragmentation (compare with their Theorem 1.11).
Theorem 1.5.
[B–] Choose any $p\in(0,1)$. In Stage 1, cut a given stick at proportion $p$ to create two pieces of length $p$ and 1p1-p. In Stage 2, cut each resulting piece into two pieces at the same proportion $p$. Perform this process for a total of $N$ stages, generating $2^{N}$ sticks with $N+1$ distinct lengths (assuming $p\neq 1/2$) given by
$x_{1}$ $\ =\ Lp^{N}$
$x_{2}$ $\ =\ Lp^{N-1}(1-p)$
$x_{3}$ $\ =\ Lp^{N-2}(1-p)^{2}$
$\quad\vdots$
$x_{N}$ $\ =\ Lp(1-p)^{N-1}$
$x_{N+1}$ $\ =\ L(1-p)^{N},$ (1.4)
where the frequency of $x_{n}$ is $\binom{N}{n}/2^{N}$. Choose $y$ so that $B^{y}=(1-p)/p$, which is the ratio of adjacent lengths (i.e., $x_{i+1}/x_{i}$). The decomposition process results in stick lengths that converge to strong Benford’s behavior base $B$ if and only if $y\not\in\mathbb{Q}$. If $y$ has finite irrationality exponent, the convergence rate can be quantified in terms of that exponent, and there is a power savings.
Theorem 1.5 suggests the possibility of examining another fixed proportion 1-dimensional stick fragmentation model, where we cut a piece into an arbitrary $m\geq 2$ number of pieces according m1m-1 fixed proportions at each stage. Using new techniques involving multinomial coefficients to reduce the problem into the original fixed proportion 1-dimensional stick fragmentation model in [B–], this paper establishes the following result (see Section 2).
Theorem 1.6.
For any integer $m>2$, choose $p_{1},p_{2},\dots,p_{m-1}\in(0,1)$ such that $p_{1}+p_{2}+\dots+p_{m-1}<1$. Set $p_{m}:=1-(p_{1}+p_{2}+\dots+p_{m-1})$. At each stage, we cut all sticks according to proportions $p_{1},p_{2},\dots,p_{m-1}$ to create $m$ pieces. After stage $N$, we have $m^{N}$ sticks in total, of lengths
$A_{k_{1},k_{2},\dots,k_{m}}\ =\ p_{1}^{k_{1}}p_{2}^{k_{2}}\cdots p_{m}^{k_{m}},$ (1.5)
for $0\leq k_{1},k_{2},\dots,k_{m}\leq N$ such that $k_{1}+k_{2}+\cdots+k_{m}=N$. Let $y_{i}=\log_{B}(p_{i}/p_{i+1})$ for $1\leq i\leq m-1$. Then the decomposition process results in stick lengths that converge to strong Benford’s behavior base $B$ if and only if $y_{i}\not\in\mathbb{Q}$ for some $1\leq i\leq m-1$.
Another possibility is to consider fragmentation processes in higher dimensions. This was considered by Durmić and Miller in [DM] and Benford behavior was established in the case of volume of an arbitrary $m$-dimensional box under mild assumptions. This was generalized by Betti et al. in [BDMMM] to any $d$-dimensional volume of $d$-dimensional faces of an arbitrary $m$-dimensional box, where $d\leq m$. To state their results, we start with some definitions.
Definition 1.7.
We say that a set $\mathfrak{B}\subset\mathbb{R}^{m}$ is an $m$-dimensional box if it is a set of the form $a_{1},b_{1}]\times\cdots\times[a_{m},b_{m}]\subset\mathbb{R}^{m}$, where $a_{i}<b_{i}$ are finite numbers.
Definition 1.8.
A linear-fragmentation process is a sequence of random variables $\mathfrak{B}_{0},\mathfrak{B}_{1},\mathfrak{B}_{2},\dots$ such that the following holds.
  1. (1)
    The random variables $\mathfrak{B}_{i}$ are $m$-dimensional boxes.
  2. (2)
    The random variables $\mathfrak{B}_{i}$ form a descending chain $\mathfrak{B}_{0}\supset\mathfrak{B}_{1}\supset\mathfrak{B}_{2}\supset\dots$.
  3. (3)
    The distribution $\mathfrak{B}_{n+1}$ conditioned on $\mathfrak{B}_{n}$ is some fixed distribution of independent proportion cuts $P_{1},\dots,P_{m}$ along each Cartesian axis. These $P_{i}$ are fixed over all $n\geq 0$.
  4. (4)
    The proportion cuts $P_{i}$ are continuous random variables with finite mean, variance, and third moment.
  5. (5)
    $\mathbb{E}[\log_{B}P_{i}]=\mu_{P}\in\mathbb{R}$ and $\textup{Var}[\log_{B}P_{i}]=\sigma_{P}^{2}>0$ for all $1\leq i\leq m$.
Definition 1.9.
Given an $m$-dimensional box $\mathfrak{B}$ and a positive integer $d\leq m$, we say the $d$-volume of $\mathfrak{B}=\prod_{i}[a_{i},b_{i}]$ is the sum of the $d$-dimensional volumes of the $d$-dimensional faces of $\mathfrak{B}$:
$\textup{Vol}_{d}(\mathfrak{B})\ :=\ 2^{m-d}\sum_{|I|=d}\prod_{i\in I}(b_{i}-a_{i}),$ (1.6)
where we are summing over all subsets $I\subset\{1,\dots,m\}$ with cardinality $d$.
[BDMMM] established a sufficient condition for strong Benford behavior to emerge (see their Theorem 1.9), which involves the maximum $d$-dimensional volume of a $d$-dimensional face.
Theorem 1.10 (Maximum Criterion).
[BDMMM] Let $\mathfrak{B}:=\mathfrak{B}_{0}$ be a fixed $m$-dimensional box and $\mathfrak{B}_{0}\supset\mathfrak{B}_{1}\supset\cdots$ be a linear-fragmentation process whose proportion cuts $P_{i}$ have probability density functions $f_{i}:(0,1)\rightarrow(0,\infty)$. Let
$V_{d}^{(N)}\ :=\ \textup{Vol}_{d}(\mathfrak{B}_{N})$ (1.7)
be the sequence of volumes obtained from this process and $\mathfrak{m}_{d}^{(N)}$ denote the maximum product of $d$ sides of $\mathfrak{B}_{N}$. If $\mathfrak{m}_{d}^{(N)}$ converges to strong Benford behavior base $B$ as $N\rightarrow\infty$, then so too does $V_{d}^{(N)}$ converge to strong Benford behavior base $B$ as $N\rightarrow\infty$.
In the same paper, the Maximum Criterion was verified for $d=m$ for any linear fragmentation process, and for $d=1$ in a special case.
Theorem 1.11.
Let $P_{i}^{(j)}$ be i.i.d. log-uniform distributions, i.e., $\log_{B}P_{i}^{(j)}$’s have uniform distributions. In the case of $d=1$, i.e., perimeter, the maximum side length of each box
$\mathfrak{m}_{1}^{(N)}\ =\ \textup{max}_{1\leq i\leq m}P_{i}^{(1)}\cdots P_{i}^{(N)}$ (1.8)
converges to strong Benford behavior base $B$ as $N\rightarrow\infty$.
Hence, it was conjectured in the same paper (see Conjecture 4.1) that for any linear fragmentation process, $m_{d}^{(N)}$ converges to strong Benford behavior.
Conjecture 1.12.
[BDMMM] Every linear fragmentation process satisfies the Maximum Criterion in all dimensions $1\leq d\leq m$.
We prove the conjecture for any $m\geq 2$ and $d\leq m$ under some mild conditions on the linear fragmentation process (see Section 3).
Theorem 1.13.
Suppose that $\mathfrak{B}_{0}\supset\mathfrak{B}_{1}\supset\mathfrak{B}_{2}\supset\cdots$ is a linear fragmentation process on an $m$-dimensional box for an arbitrary $m\geq 2$, such that the proportion cut $P_{i}$ at every stage satisfies the following.
  1. (1)
    We have $\mathbb{E}[\log_{B}(P_{i})]=0$ and $\textup{Var}[\log_{B}(P_{i})]=1$.
  2. (2)
    We have $\log(P_{i})$ is supported on [C,C][-C,C], where $C>0$ is a constant.
  3. (3)
    Suppose that $P^{(k)}_{i}$ is a sequence of i.i.d. random variables $\sim P_{i}$. Define the random variable
    $Z_{i}^{(N)}\ :=\ \frac{\log_{B}(P_{i}^{(1)}\cdots P_{i}^{(N)})}{\sqrt{N}}.$ (1.9)
    Then $Z_{i}^{(N)}$ has probability density function $f_{Z_{i}^{(N)}}(z_{i})=\varphi(z_{i})+A(z_{i})$ and cumulative density function $F_{Z_{i}^{(N)}}(z_{i})=\Phi(z_{i})+B(z_{i})$, where $\varphi(x)$ and $\Phi(x)$ are the probability density function and cumulative density function of the standard normal $N(0,1)$, $A(x)=O(N^{-1/2-\delta})$ and $B(x)=O(N^{-\delta})$ for some $\delta>0$.
Then this linear fragmentation process satisfies the condition of the Maximum Criterion for any $d\leq m$, i.e., $\mathfrak{m}_{d}^{(N)}$ converges to strong Benford behavior base $B$ as $N\rightarrow\infty$ for any $d\leq m$, thus by Theorem 1.10, $V_{d}^{(N)}$ converges to strong Benford behavior base $B$ as $N\rightarrow\infty$ for any $d\leq m$.
2. Proof of Theorem 1.6
2.1. Preliminary
In [FILM], Fang, Irons, Lippelman and Miller proved that if $y_{i}\in\mathbb{Q}$ for all $1\leq i\leq m-1$, then the distribution of the stick lengths does not follow Benford behavior as $N\rightarrow\infty$. To be exact, they were able to show that the distribution of the logarithms of the stick lengths converges to a discrete distribution, contrary to a uniform distribution on $[0,1]$ which is an equivalent condition for convergence to strong Benford behavior by Uniform Distribution Characterization 1.4. Hence, having $y_{i}\not\in\mathbb{Q}$ for some $1\leq i\leq m-1$ is a necessary condition for the distribution of the stick lengths to follow Benford behavior.
For the remainder of this section, we show that having $y_{i}\not\in\mathbb{Q}$ for some $1\leq i\leq m-1$ is also a sufficient condition for the distribution of the stick lengths to converge to Benford’s behavior. Since we can always reorder the $p_{i}^{k_{i}}$’s in the stick length $A_{k_{1},k_{2},\dots,k_{m}}:=p_{1}^{k_{1}}p_{2}^{k_{2}}\cdots p_{m}^{k_{m}}$ and thus the order of $y_{i}$’s in the factorization, then it suffices to show that the first exponent $y_{1}\not\in\mathbb{Q}$ is a sufficient condition for the distribution of the stick lengths to converge to Benford’s behavior.
We know that after stage $N$, the number of sticks of length $A_{k_{1},k_{2},\dots,k_{m}}$ is equal to the multinomial coefficient
$\binom{k_{1}+k_{2}+\cdots+k_{m}}{k_{1},k_{2},\dots,k_{m}}\ =\ \binom{N}{k_{1},k_{2},\dots,k_{m}}$. (2.1)
From the proof of Theorem , we know that the stick length $A_{k_{1},k_{2},\dots,k_{m}}$ has the factorization
$A_{k_{1},k_{2},\dots,k_{m}}\ =\ p_{1}^{k_{1}}p_{2}^{k_{2}}\cdots p_{m}^{k_{m}}\ =\ \left(\frac{p_{1}}{p_{2}}\right)^{k_{1}}\left(\frac{p_{2}}{p_{3}}\right)^{k_{1}+k_{2}}\cdots\left(\frac{p_{m-1}}{p_{m}}\right)^{\sum_{j=1}^{m-1}k_{j}}(p_{m})^{N},$ (2.2)
which motivates us to define the exponents $y_{i}=\log(p_{i}/p_{i+1})$ for $1\leq i\leq m-1$. Now, fix an interval $(a,b)\subset(0,1)$. Let
$\chi_{(a,b)}(x)\ :=\ \mathbbm{1}\Big{(}\log(x)\textup{ mod 1}\in(a,b)\Big{)}$ (2.3)
be the indicator function for $\log(x)\textup{ mod 1}\in(a,b)$. Then after stage $N$, the probability that a stick length mod 1 is in $(a,b)$ equals
$F_{N}(a,b)\ :=\ \frac{1}{m^{N}}\sum_{\begin{subarray}{c}0\leq k_{1},k_{2}\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}\chi_{(a,b)}\left(\prod_{i=1}^{m}p_{i}^{k_{i}}\right).$ (2.4)
Our goal is to prove that as $N\rightarrow\infty$,
$F(a,b)\ :=\ \lim_{N\rightarrow\infty}F_{N}(a,b)\ =\ b-a,$ (2.5)
which is a sufficient condition for Benford behavior.
2.2. Combinatorial identities
The key steps to our proof are two combinatorial identities related to multinomial coefficients, which we present and prove below. We first review the multinomial coefficients.
Definition 2.1.
Suppose that $n,k_{1},k_{2},\dots,k_{m}\geq 0$ and $k_{1}+\cdots+k_{m}=n$, then we define the multinomial coefficients by
$\binom{n}{k_{1},k_{2},\dots,k_{m}}\ :=\ \frac{n!}{k_{1}!k_{2}!\cdots k_{m}!}.$ (2.6)
The following identity on the factorization of multinomial coefficients allows us to express multinomial coefficients in terms of product of binomial coefficients. As we shall see, this forms the key ingredient that reduces our multi-proportion problem into the original single-proportion problem in [B–].
Lemma 2.2.
[Stan] For $m\geq 1$,
$\binom{k_{1}+k_{2}+\cdots+k_{m}}{k_{1},k_{2},\dots,k_{m}}\ =\ \prod_{i=1}^{m}\binom{k_{1}+\cdots+k_{i}}{k_{i}}$. (2.7)
The proof for the identity above is a straightforward induction through the expansion of the multinomial and binomial coefficients. Below is another identity that is especially useful later for evaluating sums of multinomial coefficients.
Lemma 2.3.
[Stan] For $N\in\mathbb{Z}_{\geq 0}$ and $m\in\mathbb{Z}_{>0}$,
$m^{N}\ =\ \sum_{\begin{subarray}{c}0\leq k_{1},k_{2},\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}\ =\ \sum_{\begin{subarray}{c}0\leq k_{1},\dots,k_{m-1}\leq N\\ k_{1}+k_{2}+\cdots+k_{m-1}=N\end{subarray}}2^{k_{1}}\binom{N}{k_{1},k_{2},\cdots,k_{m-1}}$. (2.8)
The most straightforward proof of the identity above relies on two different ways of representing and expanding $m^{N}$: one as $(\underbrace{1+\cdots+1}_{\textrm{$N$}})^{N}$ and the other as $(2+\underbrace{1+\cdots+1}_{\textrm{$N-2$}})^{N}$. See [Stan] for more.
2.3. Truncation
We prove that for any $\epsilon\in(0,1)$, the contribution to $F_{N}(a,b)$ of multinomial coefficients with $k_{1}+k_{2}<N^{\epsilon}$ is negligible. This allows us to only consider $k_{1}+k_{2}$ that are sufficiently large for these stick lengths $A_{k_{1},k_{2},\dots,k_{m}}$.
Proposition 2.4.
For any $\epsilon\in(0,1)$,
$\sum_{\begin{subarray}{c}0\leq k_{1},k_{2}\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}\ \leq\ (m-2)^{N}N^{N^{\epsilon}+2}.$ (2.9)
Proof.
We know that if $k_{1}+k_{2}+\cdots+k_{m}=N$, then
$\binom{N}{k_{1},k_{2},\dots,k_{m}}$ $\ =\ \frac{N(N-1)\cdots(N-(k_{1}+k_{2})+1)}{k_{1}!k_{2}!}\cdot\frac{(N-(k_{1}+k_{2}))!}{k_{3}!k_{4}!\cdots k_{m}!}$
$\ \leq\ N^{k_{1}+k_{2}}\cdot\binom{N-(k_{1}+k_{2})}{k_{3},k_{4},\dots,k_{m}}$. (2.10)
Hence,
$\sum_{\begin{subarray}{c}0\leq k_{1},k_{2},\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}$ $\ =\ \sum_{\begin{subarray}{c}0\leq k_{1},k_{2}<N^{\epsilon}\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}N^{k_{1}+k_{2}}\sum_{\begin{subarray}{c}0\leq k_{3},k_{4},\dots,k_{m}\leq N-(k_{1}+k_{2})\\ k_{3}+k_{4}+\cdots+k_{m}=N-(k_{1}+k_{2})\end{subarray}}\binom{N-(k_{1}+k_{2})}{k_{3},k_{4},\dots,k_{m}}$
$\ \leq\ (m-2)^{N}\sum_{\begin{subarray}{c}0\leq k_{1},k_{2}<N^{\epsilon}\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}N^{k_{1}+k_{2}}$
$\ \leq\ (m-2)^{N}(N^{\epsilon})^{2}N^{N^{\epsilon}}\ \leq\ (m-2)^{N}N^{N^{\epsilon}+2}.$ (2.11)
We know the probability that a stick length $p_{1}^{k_{1}}p_{2}^{k_{2}}\cdots p_{m}^{k_{m}}$ satisfies $k_{1}+k_{2}<N^{\epsilon}$ is
$\frac{1}{m^{N}}\sum_{\begin{subarray}{c}0\leq k_{1},k_{2}\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}$. (2.12)
According to Proposition 2.4, we have that (2.12) is bounded above by $((m-2)/m)^{N}N^{N^{\epsilon}+2}$. Hence, the logarithm of the probability is bounded above by
$\log\left(\left(\frac{m-2}{m}\right)^{N}N^{N^{\epsilon}+2}\right)\ =\ N\log\left(\frac{m-2}{m}\right)+(N^{\epsilon}+2)\log(N).$ (2.13)
It is clear that (2.13) goes to -$\infty$ as $N\rightarrow\infty$. Thus the probability goes to 0 as $N\rightarrow\infty$. So it suffices to consider the case where $k_{1}+k_{2}\geq N^{\epsilon}$. Now, fix $k(N)\geq N^{\epsilon}$ and let $0\leq k_{1},k_{2}\leq k(N)$ with $k_{1}+k_{2}=k(N)$ and $0\leq k_{3},k_{4},\dots,k_{m}\leq N-k(N)$ with $k_{3}+k_{4}+\dots+k_{m}=N-k(N)$. By [B–], we know that the frequency $k_{1}$ of stick length $A_{k_{1},k_{2},\dots,k_{m}}=p_{1}^{k_{1}}p_{2}^{k_{2}}\dots p_{m}^{k_{m}}$ follows a binomial distribution with mean $k(N)/2$ and standard deviation $\sqrt{k(N)}/2$. Pick some $\delta\in(0,\epsilon/10)$. We see that it suffices to consider cases where $|k_{1}-k(N)/2|<(\lceil N^{\delta}\rceil\sqrt{k(N)})/2$, because the probability that $k_{1}$ is outside this range is asymptotically small by Chebyshev’s inequality:
$\mathbb{P}\left(\left|k_{1}-\frac{k(N)}{2}\right|\ \geq\ \frac{\lceil N^{\delta}\rceil\sqrt{k(N)}}{2}\right)\ \leq\ \frac{1}{\lceil N^{\delta}\rceil^{2}}.$ (2.14)
2.4. Near uniform probability within small intervals
[…]
$k_{1,\ell}\ :=\ f_{\ell}\left(\frac{k(N)}{2}\right)+\ell\lceil N^{\delta}\rceil,\quad k_{1,\ell,i}\ :=\ f_{\ell}\left(\frac{k(N)}{2}\right)+\ell\lceil N^{\delta}\rceil+i,\quad 0\leq i\leq N^{\delta}-1,$ (2.15)
[…]
We follow Section (5.52) of [B–] to obtain a bound for the difference.
Proposition 2.5.
For $\ell\leq\sqrt{k(N)}/2$,
$\left|\binom{k(N)}{k_{1,\ell}}-\binom{k(N)}{k_{1,\ell+1}}\right|\ \leq\ O\left(\binom{k(N)}{k_{1,\ell}}\cdot N^{-\frac{3\epsilon}{10}}\right).$ (2.17)
Since the proof follows similar argument to Section (5.52) of [B–], we leave the proof details to Appendix A.
2.5. Equidistribution within small intervals
For fixed $k(N):=k_{1}+k_{2},k_{3},k_{4}\dots,k_{m}$, and $\ell$, the stick length
$A_{k_{1,\ell,i},k_{2},\dots,k_{m}}\ =\ \left(\frac{p_{1}}{p_{2}}\right)^{k_{1,\ell,i}}\left(\frac{p_{2}}{p_{3}}\right)^{k(N)}\left(\frac{p_{3}}{p_{4}}\right)^{k(N)+k_{3}}\cdots\left(\frac{p_{m-1}}{p_{m}}\right)^{k(N)+\sum_{j=3}^{m-1}k_{j}}(p_{m})^{N}$ (2.18)
[…]
Definition 2.6.
[KN] Suppose that $x$ is a real number. The irrationality exponent $\mu$ of $x$ is the supremum of the set of $\mu$ such that $0<|x-p/q|<1/q^{\mu}$ is satisfied by an infinite number of coprime integer pairs $(p,q)$ with $q>0$.
Theorem 2.7.
[KN, Theorem 3.2, Example 3.2] Let $\kappa$ be the irrationality exponent of $\log_{B}(p_{1}/p_{2})$. If $\kappa<\infty$, then
$|J_{\ell}(a,b)|\ =\ (b-a)N^{\delta}+O\left(N^{\delta\left(1-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\right),$ (2.19)
for some $\epsilon^{\prime}>0$ such that $\frac{1}{\kappa}+\epsilon^{\prime}<0$ and there is a power saving with the error term. If $\kappa=\infty$, then
$|J_{\ell}(a,b)|\ =\ (b-a)N^{\delta}+o(N^{\delta}).$ (2.20)
In either case, the error term is asymptotically much smaller than the main term $(b-a)N^{\delta}$. Thus, if $\log_{B}(p_{1}/p_{2})\not\in\mathbb{Q}$, then $\log_{B}(A_{k_{1,\ell,i},k_{2},\dots,k_{m}})$ converges to being equidistributed mod 1.
2.6. Evaluation of sums within and across intervals
[…]
$\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$. (2.21)
By Proposition 2.5 and Section 2.5, we have
$\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$ $\ =\ \sum_{i\in J_{\ell}(a,b)}\Bigg{\{}\binom{k(N)}{k_{1,\ell}}+O\left(\binom{k(N)}{k_{1,\ell}}N^{-\frac{3\epsilon}{10}}\right)\Bigg{\}}$
$\ =\ \Bigg{\{}\binom{k(N)}{k_{1,\ell}}\sum_{i\in J_{\ell}(a,b)}1\Bigg{\}}+O\left(\binom{k(N)}{k_{1,\ell}}N^{-\frac{3\epsilon}{10}}\sum_{i\in J_{\ell}(a,b)}1\right)$
$\ =\ (b-a)N^{\delta}\binom{k(N)}{k_{1,\ell}}+O\left(\binom{k(N)}{k_{1,\ell}}N^{\delta\left(1-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\right).$ (2.22)
We justify the error term in the last line of (2.6). Since $\kappa$ is the irrationality exponent of an irrational number, then $\kappa>1$ (see [BBS]). Also, $\epsilon^{\prime}>0$, so $1<-1/\kappa+\epsilon^{\prime}$. Moreover, $1/\kappa+\epsilon^{\prime}<0$ by assumption. Combining these with the fact that $\delta\in(0,\epsilon/10)$ gives the dominant error term in the last line of (2.6).
Now, we count the number of stick lengths $A_{k_{1,\ell,i},k_{2},\dots,k_{m}}$ such that $\log_{B}(A_{k_{1,\ell,i},k_{2},\dots,k_{m}})$ mod 1 is in $(a,b)$ over all the intervals. By the truncation in Section 2.3 the main term comes from the sum over $\ell$ from $\sqrt{k(N)}/2$ to $\sqrt{k(N)}/2$. The number is given by
$\sum_{\ell=-k(N)/(2N^{\delta})}^{k(N)/(2N^{\delta})}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}\ =\ \sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}+\sum_{|\ell|>\sqrt{k(N)}/2}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$. (2.23)
Using (2.14), we can provide an upper bound on the second sum on the RHS of (2.23):
$\sum_{|\ell|>\sqrt{k(N)}/2}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$ $\ \leq\ \mathbb{P}\left(\left|k_{1}-\frac{k(N)}{2}\right|\ \geq\ \frac{N^{\delta}\sqrt{k(N)}}{2}\right)\cdot\sum_{0\leq k_{1}\leq N}\binom{k(N)}{k_{1}}$
$\ =\ \mathbb{P}\left(\left|k_{1}-\frac{k(N)}{2}\right|\ \geq\ \frac{N^{\delta}\sqrt{k(N)}}{2}\right)\cdot 2^{k(N)}\ =\ O\left(\frac{2^{k(N)}}{N^{2\delta}}\right).$ (2.24)
On the other hand, we apply (2.6) to the first sum on the RHS of (2.23):
$\sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}\ =\ \sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}(b-a)N^{\delta}\binom{k(N)}{k_{1,\ell}}+O\left(\sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}\binom{k(N)}{k_{1,\ell}}N^{\delta\left(1-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\right).$ (2.25)
By (2.6) and (5.17) and (5.18) of [B–], by choosing $q=N^{\delta}$, we have
$\sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$ $\ =\ (b-a)2^{k(N)}+O\left(2^{k(N)}N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\right).$ (2.26)
Substituting (2.6) and (2.26) into (2.23) yields
$\sum_{\ell=-k(N)/(2N^{\delta})}^{k(N)/(2N^{\delta})}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}\ =\ (b-a)2^{k(N)}+O\left(2^{k(N)}N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\right).$ (2.27)
Finally, we sum over all $k(N),k_{3},k_{4},\dots,k_{m}$, which gives us the total number of lengths $A_{k_{1},k_{2},\dots,k_{m}}$ such that $\log_{B}(A_{k_{1},k_{2},\dots,k_{m}})$ mod 1 is in $(a,b)$. We break the sum based on whether $k(N)\geq N^{\epsilon}$ or not. For $k(N)\geq N^{\epsilon}$, we apply (2.27):
$\sum_{\begin{subarray}{c}0\leq k(N),k_{3},k_{4},\dots,k_{m}\leq N\\ k(N)+k_{3}+k_{4}+\cdots+k_{m}=N\\ k(N)\geq N^{\epsilon}\end{subarray}}\binom{N}{k(N),k_{3},k_{4},\dots,k_{m}}\sum_{\ell=-k(N)/(2N^{\delta})}^{k(N)/(2N^{\delta})}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$
$\ =\ (b-a)\sum_{\begin{subarray}{c}0\leq k(N),k_{3},k_{4},\dots,k_{m}\leq N\\ k(N)+k_{3}+k_{4}+\cdots+k_{m}=N\\ k(N)\geq N^{\epsilon}\end{subarray}}2^{k(N)}\binom{N}{k(N),k_{3},k_{4},\dots,k_{m}}$
$\quad+O\left(N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\sum_{\begin{subarray}{c}0\leq k(N),k_{3},k_{4},\dots,k_{m}\leq N\\ k(N)+k_{3}+k_{4}+\cdots+k_{m}=N\\ k(N)\geq N^{\epsilon}\end{subarray}}2^{k(N)}\binom{N}{k(N),k_{3},k_{4},\dots,k_{m}}\right)$. (2.28)
Note that by Lemma 2.3
$\sum_{\begin{subarray}{c}0\leq k(N),k_{3},k_{4},\dots,k_{m}\leq N\\ k(N)+k_{3}+k_{4}+\cdots+k_{m}=N\\ k(N)\geq N^{\epsilon}\end{subarray}}2^{k(N)}\binom{N}{k(N),k_{3},k_{4},\dots,k_{m}}$
$\ =\ m^{N}-\sum_{\begin{subarray}{c}0\leq k_{1},k_{2},\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}\sum_{k_{1}=0}^{k_{1}+k_{2}}\binom{N}{k_{1},k_{2},\dots,k_{m}}$
$\ =\ m^{N}+O((m-2)^{N}N^{N^{\epsilon}+2}),$ (2.29)
where in the last line, we apply Proposition 2.4. Substituting (2.6) into (2.6) gives
$\sum_{\begin{subarray}{c}0\leq k(N),k_{3},k_{4},\dots,k_{m}\leq N\\ k(N)+k_{3}+k_{4}+\cdots+k_{m}=N\\ k(N)\geq N^{\epsilon}\end{subarray}}\binom{N}{k(N),k_{3},k_{4},\dots,k_{m}}\sum_{\ell=-k(N)/(2N^{\delta})}^{k(N)/(2N^{\delta})}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$
$\ =\ (b-a)m^{N}+O\left((m-2)^{N}N^{N^{\epsilon}+2}+m^{N}N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}+(m-2)^{N}N^{N^{\epsilon}+2}N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\right)$
$\ =\ (b-a)m^{N}+O\left(m^{N}N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\right).$ (2.30)
For $k(N)<N^{\epsilon}$, by Lemma 2.7 and Proposition 2.4 we have
$\sum_{\begin{subarray}{c}0\leq k(N),k_{3},k_{4},\dots,k_{m}\leq N\\ k(N)+k_{3}+k_{4}+\cdots+k_{m}=N\\ k(N)<N^{\epsilon}\end{subarray}}\binom{N}{k(N),k_{3},k_{4},\dots,k_{m}}\sum_{\ell=-k(N)/(2N^{\delta})}^{k(N)/(2N^{\delta})}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$
$\ \leq\ \sum_{\begin{subarray}{c}0\leq k(N),k_{3},k_{4},\dots,k_{m}\leq N\\ k(N)+k_{3}+k_{4}+\cdots+k_{m}=N\\ k(N)<N^{\epsilon}\end{subarray}}\binom{N}{k(N),k_{3},k_{4},\dots,k_{m}}\sum_{k_{1}=0}^{k(N)}\binom{k(N)}{k_{1}}$
$\ =\ \sum_{\begin{subarray}{c}0\leq k_{1},k_{2},\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}\ =\ O\left((m-2)^{N}N^{N^{\epsilon}+2}\right).$ (2.31)
Combining (2.6) and (2.6), we have
$\sum_{\begin{subarray}{c}0\leq k(N),k_{3},k_{4},\dots,k_{m}\leq N\\ k(N)+k_{3}+k_{4}+\cdots+k_{m}=N\end{subarray}}\binom{N}{k(N),k_{3},k_{4},\dots,k_{m}}\sum_{\ell=-k(N)/(2N^{\delta})}^{k(N)/(2N^{\delta})}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$
$\ =\ (b-a)m^{N}+O\left(N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}m^{N}+N^{N^{\epsilon}+2}(m-2)^{N}\right)$
$\ =\ (b-a)m^{N}+O\left(N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}m^{N}\right),$ (2.32)
After dividing by $m^{N}$, the total number of sticks after stage $N$, we arrive at
$F_{N}(a,b)\ =\ b-a+O\left(N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\right).$ (2.33)
We have shown that if $\log(p_{1}/p_{2})$ has a finite irrationality exponent $\kappa$, then the logarithm of the stick lengths is equidistributed mod 1, and thus the stick lengths converge to the Benford distribution. If $\log(p_{1}/p_{2})$ has infinite irrationality exponent, the calculation is exactly the same, except that we start with the error term $o(N^{\delta})$ instead of $O(N^{\delta\left(1-\frac{1}{\kappa}+\epsilon^{\prime}\right)})$ in Section 2.5, and so we end up with the $F_{N}(a,b)\ =\ b-a+o(1)$. Thus, this establishes the strong Benford behavior of fixed multi-proportion 1-dimensional stick fragmentation process when $y_{i}\not\in\mathbb{Q}$ for some $1\leq i\leq m-1$. ∎
3. Proof of Theorem 1.13
3.1. Preliminary
In this section, we prove Theorem 1.13, i.e., resolve Conjecture 4.1 of [BDMMM] that we mentioned in Section 1 under some mild conditions. Let us first recall the conjecture.
Conjecture 3.1.
Every linear fragmentation process satisfies the maximum criterion in all dimensions $1\leq d\leq m$.
[BDMMM] has verified the conjecture in the case of $d=1$ and arbitrary $m$ for i.i.d. log-uniform distributions, i.e., the proportion at which each dimension is cut is $P_{i}$, where $\log P_{i}\sim\textup{Uniform}(a,b)$ and $a<b\leq 0$. They decide to normalize $P_{i}$ to $\textup{Uniform}(-\sqrt{3},\sqrt{3})$ with mean 0 and variance 1, which simplifies the calculation later on and does not affect the nature or the solution to the problem. After stage $N$, the side length of the $i$th dimension of the box, after taking logarithm and normalizing, is
$Z_{i}^{(N)}\ :=\ \frac{\log_{B}(P_{i}^{(1)}\cdots P_{i}^{(N)})-N\mu_{P}}{\sqrt{N}\sigma_{P}}\ =\ \frac{\log_{B}(P_{i}^{(1)}\cdots P_{i}^{(N)})}{\sqrt{N}}.$ (3.1)
By the CLT, $Z_{i}^{(N)}$ converges to $N(0,1)$. Moreover, Lemma 3.1 states that if $f_{Z_{i}^{(N)}}(z)$ is the probability density function of $Z_{i}^{(N)}$ and $\varphi(z)$ is the probability density function of $N(0,1)$, then for all $z\in[-\sqrt{3}N,\sqrt{3}N]$ and any $\epsilon>0$, we have
$f_{Z_{i}^{(N)}}(z_{i})\ =\ \varphi(z_{i})+O(N^{-1+4\epsilon}),$ (3.2)
where
$\varphi(x)\ :=\ \frac{1}{\sqrt{2\pi}}e^{-x^{2}/2}.$ (3.3)
Due to the presence of the error term, to establish the strong Benford behavior of the process, they need to rely on the fact that the distributions for the cuts are finitely supported with uniformly bounded supports. They also need to rely on the fact that the exponent of the error term is less than 1/2-1/2 for sufficiently small $\epsilon>0$, which allows the error term from integrating $f_{Z^{(N)}_{i}}(z)$ over the support to be asymptotically 0.
For the rest of this paper, we prove Conjecture 4.1 of [BDMMM] for any $m\geq 2$ and $d\leq m$. Physically, this means that we consider any arbitrary dimensional volume of faces of that dimension of the $m$-dimensional box. We assume that the $\log(P_{i})$ are i.i.d., have mean 0 and variance 1, are finitely supported with support [C,C][-C,C], and that the probability density function (PDF) of $Z_{i}^{(N)}$ is $f_{Z_{i}^{(N)}}(z_{i})=\varphi(z_{i})+A(z_{i})$, where $A(z_{i})=O(N^{-1/2-\delta})$ for some $\delta>0$. We also assume that the cumulative density function (CDF) of $Z_{i}^{(N)}$ is given by $F_{Z_{i}^{(N)}}(z_{i})=\Phi(z_{i})+B(z_{i})$, where $\Phi$ is the CDF of $N(0,1)$ and $B(z_{i})=O(N^{-\delta})$. We prove that the maximum $d$-dimensional volume $\mathfrak{m}_{d}^{(N)}$ of a $d$-dimensional face of the $m$-dimensional box converges to strong Benford behavior. Once we prove this statement, then by the maximum criterion established in [BDMMM], we have that the $d$-dimensional volume $V_{d}^{(N)}$ of the $d$-dimensional faces of the $m$-dimensional box also converges to strong Benford behavior.
The key idea behind our proof is the use of the Mean Value Theorem and a recursive method of change of variables and consistent cancellation that provide upper bounds on the error terms in terms of Gaussian densities. We shall explore this more in Section 3.6.
To begin, we use knowledge from order statistics to find the PDF of $\mathfrak{m}_{d}^{(N)}$.
Definition 3.2.
Suppose that $X_{1},X_{2},\dots,X_{k}$ are random variables, and $_{i}$ returns the $i$th largest number among a list $(a_{1},a_{2},\dots,a_{k})$ of real numbers. For each outcome $\omega$ in the sample space $\Omega$, we define
$X_{(k-i+1)}(w)\ :=\ [(X_{1}(w),X_{2}(w),\dots,X_{k}(w))]_{k-i+1}.$ (3.4)
We say that the random variable $X_{(k-i+1)}$ is the $i$th order statistics, or the $i$th largest random variable among $X_{1},X_{2},\dots,X_{k}$. Hence, as random variables, $X_{(1)}\leq X_{(2)}\leq\dots\leq X_{(k)}$.
Let $S_{i}^{(N)}$ be the side length of the $i\textsuperscript{th}$ dimension of the $m$-dimensional box after stage $N$. Then the maximum volume of a $d$-dimensional face is the product of the longest $d$ sides of the box, i.e., $\mathfrak{m}_{d}^{(N)}=\prod_{i=1}^{d}S_{(m+1-i)}^{(N)}$. We define
$Y^{(N)}\ :=\ \frac{\log_{B}(\mathfrak{m}_{d}^{(N)})}{\sqrt{N}}\ =\ \sum_{i=1}^{d}\frac{\log_{B}(S^{(N)}_{(m+1-i)})}{\sqrt{N}}\ =\ \sum_{i=1}^{d}Z_{(m+1-i)}^{(N)},$ (3.5)
To find the PDF of $Y^{(N)}$, we need the joint PDF of $Z_{(m+1-i)}^{(N)}$ for all $1\leq i\leq d$. This is a standard question in order statistics.
Proposition 3.3.
[DL] For i.i.d. random variables $X_{1},\dots,X_{k}$ with PDF $f(x)$ and CDF $F(x)$, the joint PDF of $X_{(i)},\dots,X_{(k)}$ where $1\leq i\leq k$ is given by
$f_{X_{(i)},\dots,X_{(k)}}(x_{i},\dots,x_{k})\ =\ C_{k}^{i}f(x_{i})\cdots f(x_{k})(F(x_{i}))^{i-1},$ (3.6)
where $C_{k}^{i}:=k!/(i-1)!$.
Thus, in our case, the joint PDF of $Z_{(m+1-i)}^{(N)}$ for all $1\leq i\leq d$ is
$f_{Z_{(m+1-d)},\dots,Z_{(m)}}(z_{m+1-d},\dots,z_{m})$
$\ =\ C_{m}^{m+1-d}\left(\Phi(z_{m+1-d})+B(z_{m+1-d})\right)^{m-d}\prod_{i=1}^{d}\left(\varphi(z_{m+i-d})+A(z_{m+i-d})\right).$ (3.7)
By binomial expansion,
$\left(\Phi(z_{m+1-d})+B(z_{m+1-d})\right)^{m-d}\ =\ \sum_{j=0}^{m-d}\binom{m-d}{j}(\Phi(z_{m+1-d}))^{m-d-j}(B(z_{m+1-d}))^{j}.$ (3.8)
Note that $B(z_{m+1-d})=O(N^{-\delta})$ and $0\leq\Phi(x)\leq 1$, which we shall constantly use without reference for the remainder of the paper. Then any term where $j>1$ is at most of the same magnitude as $B(z_{m+1-d})$. We have
$\left(\Phi(z_{m+1-d})+B(z_{m+1-d})\right)^{m-d}\ =\ \left(\Phi(z_{m+1-d})\right)^{m-d}+O(B(z_{m+1-d})).$ (3.9)
We make a slight abuse of notation here to write the error term above as $B(z_{m+1-d})$, which does not change the order of magnitude of the error term. Hence,
$f_{Z_{(m+1-d)},\dots,Z_{(m)}}(z_{m+1-d},\dots,z_{m})$
$\ =\ C_{m}^{m+1-d}\left(\Phi(z_{m+1-d})^{m-d}+B(z_{m+1-d})\right)\prod_{i=1}^{d}\left(\varphi(z_{m+i-d})+A(z_{m+i-d})\right).$ (3.10)
3.2. PDF of $Y^{(N)}$
In this section, we find the PDF of $Y^{(N)}$ as defined in (3.1). The CDF $F_{Y^{(N)}}(y)$ of $Y^{(N)}$ is given by integrating over the appropriate region in $\mathbb{R}^{d}$, that is, over all the values $(z_{m+1-d},\dots,z_{m})$ of $(Z_{(m+1-d)}^{(N)},\dots,Z_{(m)}^{(N)})$ that sums to $y$. We know that $Z_{(m+1-d)}^{(N)},\dots Z_{(m)}^{(N)}$ are not independent, i.e., they have to satisfy $Z_{(m+1-d)}^{(N)}\leq\cdots\leq Z_{(m)}^{(N)}$. In general, we know that any $1\leq j\leq d$, $z_{m+j-d}$ can take value from $z_{m+(j-1)-d}$ to $(y-\sum_{i=1}^{j-1}z_{m+i-d})/(d-j+1)$. The only exception is $z_{m+1-d}$, whose upper bound is already correctly stated but the lower bound is $C\sqrt{N}$. Hence, we have
$F_{Y^{(N)}}(y)$
$\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z_{(m+1-d)},\dots,Z_{(m)}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d},$ (3.11)
where we use the product of integrals to denote
$\prod_{j=k}^{d}\int_{z_{j}}^{y_{j}}\ :=\ \int_{z_{k}}^{y_{k}}\cdots\int_{z_{j}}^{y_{j}}\cdots\int_{z_{d}}^{y_{d}}.$ (3.12)
We separate out the main term $C_{m}^{m+1-d}\Phi(z_{m+1-d})^{m-d}\prod_{i=1}^{d}\varphi(z_{m+i-d})$ of the joint PDF (3.1) and denote it by
$m_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})\ :=\ C_{m}^{m+1-d}\Phi(z_{m+1-d})^{m-d}\prod_{i=1}^{d}\varphi(z_{m+i-d}).$ (3.13)
We also denote the error term of the joint PDF (3.1) by $e_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})$. Let $\mathcal{M}_{Y^{(N)}}(y)$ and $\mathcal{E}_{Y^{(N)}}(y)$ be the main term and the error term of $F_{Y^{(N)}}(y)$. Then
$F_{Y^{(N)}}(y)$ $\ =\ \mathcal{M}_{Y^{(N)}}(y)+\mathcal{E}_{Y^{(N)}}(y)$
$\mathcal{M}_{Y^{(N)}}(y)$ $\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}m_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}$
$\mathcal{E}_{Y^{(N)}}(y)$ $\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}e_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}.$ (3.14)
We are now ready to find the PDF $f_{Y^{(N)}}$ of $Y^{(N)}$. We know that
$f_{Y^{(N)}}(y)\ =\ \frac{d}{dy}F_{Y^{(N)}}(y)\ =\ \frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)+\frac{d}{dy}\mathcal{E}_{Y^{(N)}}(y).$ (3.15)
To differentiate $\mathcal{M}_{Y^{(N)}}(y)$ and $\mathcal{E}_{Y^{(N)}}(y)$, we apply the Leibniz integral rule for differentiate under the integral sign.
Lemma 3.4.
[F] Suppose that $f(x,t)$ is a function such that both $f(x,t)$ and $df(x,t)/dx$ are continuous in $t$ and $x$ in the $xt$-plane, and $a(x)$ and $b(x)$ are also continuously differentiable. Then
$\frac{d}{dx}\int_{a(x)}^{b(x)}f(x,t)dt\ =\ f(x,b(x))\cdot\frac{d}{dx}b(x)-f(x,a(x))\cdot\frac{d}{dx}a(x)+\int_{a(x)}^{b(x)}\frac{\partial}{\partial x}f(x,t)dt.$ (3.16)
For the rest of this paper, one can easily check the regularity conditions on $f(x,t)$, $a(x)$, and $b(x)$ each time we apply the Leibniz integral rule, so there will not be explicit mention of the regularity conditions again. We first make some observations on the effect of $d/dy$ on
$\int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}g(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}$ (3.17)
for any Riemann integrable function $g(z_{m+1-d},\dots,z_{m})$. One can check that each application of the Leibniz integral rule to (3.17) gives two integrals, one of which has the lower bound of the integration coinciding with the upper bound of the integration and thus equals to 0. As a result, we are always left with one integral that interchanges the order of the differentiation and the integration signs. Eventually, (3.17) becomes
$\int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\frac{\partial}{\partial y}\left(\int_{z_{m-1}}^{y-\sum_{i=1}^{d-1}z_{m+i-d}}g(z_{m+1-d},\dots,z_{m})dz_{m}\right)dz_{m-1}\cdots dz_{m+1-d}$
$\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}g\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\cdots dz_{m+1-d}.$ (3.18)
Applying (3.2) to $d\mathcal{M}_{Y^{(N)}}(y)/dy$ and $d\mathcal{E}_{Y^{(N)}}(y)/dy$, where $\mathcal{M}_{Y^{(N)}}(y)$ and $\mathcal{E}_{Y^{(N)}}(y)$ are as defined in (3.2), we have
$\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)$ $\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}m_{Y^{(N)}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}$
$\quad\quad\cdots dz_{m+1-d}$
$\frac{d}{dy}\mathcal{E}_{Y^{(N)}}(y)$ $\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}e_{Y^{(N)}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}$
$\quad\quad\cdots dz_{m+1-d}.$ (3.19)
3.3. Outline of problem
We are ready to formulate our problem. We want to show that $\mathfrak{m}_{d}^{(N)}$ converges to strong Benford behavior, which is equivalent to showing that $\log_{B}(\mathfrak{m}_{d}^{(N)})$ converges to being equidistributed mod 1. Hence, in terms of the probability density function $f_{Y^{(N)}}(y)$ of $Y^{(N)}:=\log_{B}(\mathfrak{m}_{d}^{(N)})/\sqrt{N}$, we can formulate our problem as to showing
$F_{N}(a,b)\ :=\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}f_{Y^{(N)}}(y)dy\ \approx\ b-a,$ (3.20)
for all $(a,b)\subset(0,1)$ for sufficiently large $N$. Let the main term and the error term of $F_{N}(a,b)$ be defined respectively as
$\mathcal{M}_{N}(a,b)$ $\ :=\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)dy;$ (3.21)
$\mathcal{E}_{N}(a,b)$ $\ :=\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}\mathcal{E}_{Y^{(N)}}(y)dy.$ (3.22)
Our goal is to show
$\mathcal{M}_{N}(a,b)\ \approx\ b-a,\quad\mathcal{E}_{N}(a,b)\ \approx\ 0.$ (3.23)
3.4. Upper bound on error term $\mathcal{E}_{N}(a,b)$
We want to show that the error term $\mathcal{E}_{N}(a,b)$ is negligible. We start by giving some bounds on $d\mathcal{E}_{Y^{(N)}}(y)/dy$ from (3.2). Recall that $e_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})$ is the error term of the joint density function (3.1). Let $\ell$] be short hand for $\{1,\dots,\ell\}$. Then
$\frac{1}{C_{m}^{m+1-d}}e_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})$
$\quad\ =\ \Phi(z_{m+1-d})^{m-d}\left(\prod_{i=1}^{d}(\varphi(z_{m+i-d})+A(z_{m+i-d}))-\prod_{i=1}^{d}\varphi(z_{m+i-d})\right)$
$\quad\quad\quad+B(z_{m+1-d})\prod_{i=1}^{d}(\varphi(z_{m+i-d})+A(z_{m+i-d}))$
$\quad\ =\ \left(\Phi(z_{m+1-d})^{m-d}+B(z_{m+1-d})\right)\left(\sum_{\begin{subarray}{c}S\subset[d]\\ S\neq d\end{subarray}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\right)$
$\quad\quad\quad+B(z_{m+1-d})\prod_{i=1}^{d}\varphi(z_{m+i-d})$
$\ \ll\ \sum_{\begin{subarray}{c}S\subset[d]\\ S\neq d\end{subarray}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)+N^{-\delta}\prod_{i=1}^{d}\varphi(z_{m+i-d}).$ (3.24)
We know from (3.2) that to bound $d\mathcal{E}_{Y^{(N)}}(y)/dy$, it suffices to bound
$D_{1}(y)$ $\ :=\ \sum_{\begin{subarray}{c}S\subset[d]\\ S\neq[d]\end{subarray}}\int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\Bigg{|}_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}$
$\quad\quad dz_{m-1}\cdots dz_{m+1-d}$ (3.25)
as well as
$D_{2}(y)\ :=\ N^{-\delta}\int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\prod_{i=1}^{d}\varphi(z_{m+i-d})\Bigg{|}_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}dz_{m-1}\cdots dz_{m+1-d},$ (3.26)
because then we have $d\mathcal{E}_{Y^{(N)}}dy\ll D_{1}(y)+D_{2}(y)$. We first look at $D_{1}(y)$. Suppose that $k_{S}$ is the largest index such that $k\in[d]\setminus S$. Such an index $k$ exists, since $S\neq[d]$. First, assume that we have $S$’s where $k_{S}=d$. Then the integrand of (3.4) becomes
$\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\Bigg{|}_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}$
$\quad\ =\ \left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d-1]\setminus S}A(z_{m+i-d})\right)A\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
$\quad\ \ll\ N^{-\frac{1}{2}-\delta}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d-1]\setminus S}A(z_{m+i-d})\right),$ (3.27)
where in the last line we use the fact that $A(x)=O(N^{-1/2-\delta})$. Note that the lower and upper bounds on the interval of integration for each integral in (3.4) are $O(\sqrt{N})$, since $y,z_{m+1-d},\dots,z_{m-1}$ are all $O(\sqrt{N})$. Choose a constant $C^{\prime}$ such that $C^{\prime}\sqrt{N}$ is an upper bound on the upper bounds on the interval of integration for all the integrals in (3.4) and $C^{\prime}\sqrt{N}$ is a lower bound on the interval of integration for all the integrals in (3.4). Hence,
$\int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\Bigg{|}_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}$
$dz_{m-1}\cdots dz_{m+1-d}$
$\quad\ \ll\ N^{-1/2-\delta}\underbrace{\int_{-C^{\prime}\sqrt{N}}^{C^{\prime}\sqrt{N}}\cdots\int_{-C^{\prime}\sqrt{N}}^{C\sqrt{N}}}_{d-1}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d-1]\setminus S}N^{-1/2-\delta}\right)dz_{m-1}\cdots dz_{m+1-d}$
$\quad\ \ll\ N^{-1/2-\delta-|[d-1]\setminus S|\delta}\prod_{i\in S}\left(\int_{-C^{\prime}\sqrt{N}}^{C^{\prime}\sqrt{N}}\varphi(z_{m+i-d})dz_{m+i-d}\right)\ \ll\ N^{-1/2-(1+|[d-1]\setminus S|)\delta}\ \ll\ N^{-1/2-\delta},$ (3.28)
where in the last line we use the fact that
$\int_{-C^{\prime}\sqrt{N}}^{C^{\prime}\sqrt{N}}\varphi(x)dx\ \ll\ \int_{-\infty}^{\infty}\varphi(x)dx\ =\ 1,$ (3.29)
which we shall use frequently without reference from now on. Since the number of subsets $S$ of $[d]$ is finite, then (3.4) gives us
$D_{1}(y)\ \ll\ N^{-1/2-\delta}.$ (3.30)
The case when $k_{S}<d$ is similar, though more care is needed to identify the $O(N^{-1/2-\delta})$ decay, which essentially comes from the $A(z_{m+k_{S}-d})$ term. With this in mind, using the same method for the case when $k_{S}=d$ as well as the convolution formula for Gaussian PDF, we can again show that $D_{1}(y)\ll N^{-1/2-\delta}$. We leave the details to Appendix B.
We now turn to $D_{2}(y)$ as defined in (3.26), which is relatively more straightforward to bound. We have
$D_{2}(y)\ \ll\ N^{-\delta}\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-1}\prod_{i=1}^{d-1}\varphi(z_{m+i-d})\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\cdots dz_{m+1-d}.$ (3.31)
We recognize that the integral is the convolution of $d$ standard Gaussian density functions. We know that the convolution is the probability density function of sum of $d$ i.i.d. standard Gaussian random variables, which itself is also with mean 0 and variance $d$. Hence,
$D_{2}(y)\ \ll\ N^{-\delta}\frac{1}{\sqrt{2\pi d}}e^{-y^{2}/(2d)}.$ (3.32)
Thus,
$\frac{d}{dy}\mathcal{E}_{Y^{(N)}}(y)$ $\ \ll\ D_{1}(y)+D_{2}(y)\ \ll\ N^{-\delta}\frac{1}{\sqrt{2\pi d}}e^{-y^{2}/(2d)}+N^{-1/2-\delta}.$ (3.33)
We are now ready to bound the error term $\mathcal{E}_{N}(a,b)$ from (3.22):
$\mathcal{E}_{N}(a,b)$ $\ =\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}\mathcal{E}_{Y^{(N)}}(y)dy$
$\ \ll\ N^{-\delta}\int_{-dC\sqrt{N}}^{dC\sqrt{N}}\frac{1}{\sqrt{2\pi d}}e^{-y^{2}/(2d)}dy+\int_{-dC\cdot\sqrt{N}}^{dC\cdot\sqrt{N}}N^{-1/2-\delta}dy\ \ll\ N^{-\delta}.$ (3.34)
3.5. Strategy for remainder of proof
Our next task is to show that the main term $\mathcal{M}_{N}(a,b)$ defined in (3.21) satisfies
$\mathcal{M}_{N}(a,b)\ :=\ \sum_{-dC\cdot N}^{dC\cdot N-1}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)dy\ \approx\ b-a.$ (3.35)
Our strategy is the following. First, we want to write the integral of $d\mathcal{M}_{Y^{(N)}}(y)/dy$ over each interval $(a+n)/\sqrt{N},(b+n)/\sqrt{N}]$ as the sum of a main term $m_{n}(a,b)$ that is constant with respect to $y$ within the interval and an error term $e_{n}(a,b)$ for each $n$. Essentially, this shows that $d\mathcal{M}_{Y^{(N)}}(y)/dy$ is almost constant within a small interval, incurring an negligible cost that, as we shall prove, does not accumulate, i.e.,
$\mathcal{M}_{\textup{err},N}(a,b)\ =\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}e_{n}(a,b)\ \approx\ 0.$ (3.36)
Finally, we prove that the sum of the main term
$\mathcal{M}_{\textup{main},N}(a,b)\ =\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}m_{n}(a,b)$ (3.37)
is a Riemann sum that converges to bab-a. It then follows that
$\mathcal{M}_{N}(a,b)\ =\ \mathcal{M}_{\textup{main},N}(a,b)+\mathcal{M}_{\textup{err},N}(a,b)\ \approx\ b-a.$ (3.38)
Thus, based on the estimate on the error term in (3.22),
$F_{N}(a,b)\ =\ \mathcal{M}_{N}(a,b)+\mathcal{E}_{N}(a,b)\ \approx\ b-a,$ (3.39)
which establishes the equidistribution result.
3.6. Equidistribution within small interval
In this section, we show that $d\mathcal{M}_{Y^{(N)}}(y)/dy$ is equidistributed within each small interval $(a+n)/\sqrt{N},(b+n)/\sqrt{N}]$. For each $n$, we write the integral over $(a+n)/\sqrt{N},(b+n)/\sqrt{N}]$ in terms of a main term $m_{n}(a,b)$ that is constant with respect to $y$ and an error term $e_{n}(a,b)$, i.e.,
$\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)dy\ =\ m_{n}(a,b)+e_{n}(a,b),$ (3.40)
where
$m_{n}(a,b)$ $\ :=\ \int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}\left(\frac{n}{\sqrt{N}}\right)dy,$
$e_{n}(a,b)$ $\ :=\ \int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)-\frac{d}{dy}\mathcal{M}_{Y^{(N)}}\left(\frac{n}{\sqrt{N}}\right)dy.$ (3.41)
Since $d\mathcal{M}_{Y^{(N)}}(n/\sqrt{N})/dy$ is constant, then the main term $m_{n}(a,b)$ is
$m_{n}(a,b)\ =\ \frac{b-a}{\sqrt{N}}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}\left(\frac{n}{\sqrt{N}}\right).$ (3.42)
We now quantify the error term. We start by providing an upper bound on
$\left|\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)-\frac{d}{dy}\mathcal{M}_{Y^{(N)}}\left(\frac{n}{\sqrt{N}}\right)\right|,$ (3.43)
for $y\in[n/\sqrt{N},(n+1)/\sqrt{N}]\supset[(a+n)/\sqrt{N},(b+n)/\sqrt{N}]$. As we shall see, $d\mathcal{M}_{Y^{(N)}}(y)/dy$ is a continuous function over $n/\sqrt{N},(n+1)/\sqrt{N}]$ and a differentiable function over $(n/\sqrt{N},(n+1)/\sqrt{N})$. By the Mean Value Theorem, for any $y_{1},y_{2}\in[n/\sqrt{N},(n+1)/\sqrt{N}]$ with $y_{1}<y_{2}$, we have
$\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y_{1})-\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y_{2})\ =\ \frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(c_{n})(y_{1}-y_{2})$ (3.44)
for some $c_{n}\in(y_{1},y_{2})$. Hence
$\left|\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y_{1})-\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y_{2})\right|\ \leq\ \left|\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(c_{n})\right||y_{1}-y_{2}|\ \leq\ \frac{1}{\sqrt{N}}\left|\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(c_{n})\right|.$ (3.45)
We want to provide an upper on $|d^{2}\mathcal{M}_{Y^{(N)}}(y)/dy^{2}|$ over the interval $n/\sqrt{N},(n+1)/\sqrt{N}]$. Recall from (3.2) that
$\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)$ $\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}m_{Y^{(N)}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}$
$\quad\quad\cdots dz_{m+1-d}$
$\ =\ \ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\Phi(z_{m+1-d})\right)^{m-d}\prod_{i=1}^{d-1}\varphi(z_{m+i-d})\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
$\quad\quad dz_{m-1}\cdots dz_{m+1-d}.$ (3.46)
Following the same procedure as shown in Section 3.2, we can again exchange the order of differentiation and integration. Since the product in the integrand in (3.6) goes from $j=2$ to d1d-1, we need to discuss the cases when $d=2$ and $d>2$ separately. If $d=2$,
$\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)\ =\ \int_{-C\sqrt{N}}^{\frac{y}{2}}m_{Y^{{(N)}}}\left(z_{m-1},y-z_{m-1}\right)dz_{m-1}.$ (3.47)
By the Leibniz integral rule and (3.13),
$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)$
$\ =\ \frac{1}{2}m_{Y^{(N)}}\left(\frac{y}{2},\frac{y}{2}\right)+\int_{-C\sqrt{N}}^{\frac{y}{2}}\frac{\partial}{\partial y}\left(m_{Y^{(N)}}(z_{m-1},y-z_{m-1})\right)dz_{m-1}$
$\ \ll\ \varphi\left(\frac{y}{2}\right)+\int_{-C\sqrt{N}}^{\frac{y}{2}}(\Phi(z_{m-1}))^{m-2}\varphi(z_{m-1})\cdot(-1)\cdot(y-z_{m-1})\varphi(y-z_{m-1})dz_{m-1}.$ (3.48)
We know that $y$ is an upper bound on the value of $z_{m-1}+z_{m}$, then $y\geq-2C\sqrt{N}$, and $y/2\geq-C\sqrt{N}$. We now break into cases when $y\leq 0$ and $y\geq 0$. When $y\leq 0$, $\varphi(z_{m-1})$ is at most $\varphi(y/2)$ on $C\sqrt{N},y/2]$. So (3.6) becomes
$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)$ $\ \ll\ \varphi\left(\frac{y}{2}\right)+\varphi\left(\frac{y}{2}\right)\int_{-\infty}^{\infty}|y-z_{m-1}|\varphi(y-z_{m-1})dz_{m-1}$
$\ =\ \varphi\left(\frac{y}{2}\right)+\varphi\left(\frac{y}{2}\right)\int_{-\infty}^{\infty}|z_{m-1}|\varphi(z_{m-1})dz_{m-1}\ \ll\ \varphi\left(\frac{y}{2}\right),$ (3.49)
where in the last line we use the fact that the integral $\int_{-\infty}^{\infty}|x|\varphi(x)dx$ is the expected value of the absolute value of a standard normal random variable, which is finite. When $y\geq 0$, we have
$y-z_{m-1}$ $\ \geq\ y-\frac{y}{2}\ =\ \frac{y}{2}\ \geq\ 0.$ (3.50)
Hence, (3.6) becomes
$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)$ $\ \ll\ \varphi\left(\frac{y}{2}\right)+\int_{-C\sqrt{N}}^{\frac{y}{2}}(y-z_{m-1})\varphi(y-z_{m-1})dz_{m-1}$
$\ =\ \varphi\left(\frac{y}{2}\right)+\varphi(y-z_{m-1})\Bigg{|}_{z_{m-1}=-C\sqrt{N}}^{z_{m-1}=\frac{y}{2}}\ \ll\ \varphi\left(\frac{y}{2}\right)+\varphi\left(y+C\sqrt{N}\right),$ (3.51)
where in the second line we employ the fact that $\int(y-x)\varphi(y-x)dx=\varphi(y-x)$. Thus, if $d=2$, regardless of whether $y\leq 0$ or $y\geq 0$, $d^{2}\mathcal{M}_{Y^{(N)}}(y)/dy^{2}$ is on the order of $\varphi(y/2)+\varphi(y+C\sqrt{N})$. This concludes the estimate of $d^{2}\mathcal{M}_{Y^{(N)}}(y)/dy^{2}$ for $d=2$.
If $d>2$, we repeatedly apply Leibniz integral rule as we have demonstrated in Subsection 3.4 to obtain
$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)$
$\quad\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-2}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\frac{\partial}{\partial y}\Bigg{(}\int_{z_{m-2}}^{\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}}m_{Y^{(N)}}\Bigg{(}z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\Bigg{)}$
$\quad\quad\quad\quad dz_{m-1}\Bigg{)}dz_{m-2}\cdots dz_{m+1-d}$
$\quad\ \ll\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-2}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\Phi(z_{m+1-d})\right)^{m-d}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\right)^{2}$
$\quad\quad+\Bigg{(}\int_{z_{m-2}}^{\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}}(\Phi(z_{m+1-d}))^{m-d}\left(\prod_{i=1}^{d-1}\varphi(z_{m+i-d})\right)(-1)\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
$\quad\quad\quad\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\Bigg{)}dz_{m-2}\cdots dz_{m+1-d}.$ (3.52)
Now, similar to the case when $d=2$, we estimate the second integrand in (3.6) by breaking down into cases when $(y-\sum_{i=1}^{d-2}z_{m+i-d})/2\leq 0$ and $(y-\sum_{i=1}^{d-2}z_{m+i-d})\geq 0$. We leave the details to Appendix C. Hence, we have $d^{2}\mathcal{M}_{Y^{(N)}}(y)/dy^{2}$ in (3.6) is bounded above by
$\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)dz_{m-2}\cdots dz_{m+1-d}$
$\quad\quad+\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)dz_{m-2}\cdots dz_{m+1-d}.$ (3.53)
We first obtain some estimate on the first integral of (3.6). The idea behind estimating this integral is recursive uses of change of variables that lead to consistent cancellation. We first do the change of variable $z_{m-2}\mapsto z_{m-2}+(y-\sum_{i=1}^{d-3}z_{m+i-d})/5$. We have
$\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)dz_{m-2}\cdots dz_{m+1-d}$
$\quad\ \ll\ \underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}e^{-\frac{\sum_{i=1}^{d-3}z^{2}_{m+i-d}}{2}}\cdot e^{-\frac{\left(z_{m-2}+\left(y-\sum_{i=1}^{d-3}z_{m+i-d}\right)/5\right)^{2}}{2}}$
$\quad\quad\quad\cdot e^{-\frac{\frac{1}{4}\left(-z_{m-2}+4\left(y-\sum_{i=1}^{d-3}z_{m+i-d}\right)/5\right)^{2}}{2}}dz_{m-2}\cdots dz_{m+1-d}.$ (3.54)
Notice that in the expansion of the exponents, we get a cancellation of the term $2z_{m-2}((y-\sum_{i=1}^{d-3}z_{m+i-d})/5)$. Hence, (3.6) is bounded above by
$\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-3}e^{-\frac{\sum_{i=1}^{d-3}z_{m+i-d}^{2}}{2}}\cdot e^{-\frac{\frac{1}{5}\left(y-\sum_{i=1}^{d-3}z_{m+i-d}\right)^{2}}{2}}\left(\int_{-\infty}^{\infty}e^{-\frac{\frac{5}{4}z_{m-2}^{2}}{2}}dz_{m-2}\right)dz_{m-3}\cdots dz_{m+1-d}$
$\quad\ \ll\ \underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-3}e^{-\frac{\sum_{i=1}^{d-3}z_{m+i-d}^{2}}{2}}\cdot e^{-\frac{\frac{1}{5}\left(y-\sum_{i=1}^{d-3}z_{m+i-d}\right)^{2}}{2}}dz_{m-3}\cdots dz_{m+1-d}$ (3.55)
We see that the estimate in (3.6) has a similar structure, which allows us to use induction to estimate the interval. For an arbitrary $k$ such that $1\leq k\leq d-2$, we want to look at
$\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{k}e^{-\frac{\sum_{i=1}^{k}z^{2}_{m+i-d}}{2}}\cdot e^{-\frac{A_{k}(y-\sum_{i=1}^{k}z_{m+i-d})^{2}}{2}}dz_{m+k-d}\cdots dz_{m+1-d},$ (3.56)
where $A_{k}>0$. Using the change of variable $z_{m+k-d}\mapsto z_{m+k-d}+\frac{A_{k}}{1+A_{k}}(y-\sum_{i=1}^{k-1}z_{m+i-d})$, (3.56) is bounded above by
$\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{k}e^{-\frac{\sum_{i=1}^{k-1}z^{2}_{m+i-d}}{2}}\cdot e^{-\frac{\left(z_{m+k-d}+\frac{A_{k}}{1+A_{k}}\left(y-\sum_{i=1}^{k-1}z_{m+i-d}\right)\right)^{2}}{2}}$
$\quad\quad\quad\cdot e^{-\frac{A_{k}\left(-z_{m+k-d}+\frac{1}{1+A_{k}}\left(y-\sum_{i=1}^{k-1}z_{m+i-d}\right)\right)^{2}}{2}}dz_{m+k-d}\cdots dz_{m+1-d}.$ (3.57)
Expanding the exponents also gives a cancellation of the term
$\frac{2A_{k}}{1+A_{k}}z_{m+k-d}\left(y-\sum_{i=1}^{k-1}z_{m+i-d}\right).$ (3.58)
Hence, (3.56) is bounded above by
$\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{k}e^{-\frac{\sum_{i=1}^{k-1}z^{2}_{m+i-d}}{2}}\cdot e^{-\frac{A_{k-1}\left(y-\sum_{i=1}^{k-1}z_{m+i-d}\right)^{2}}{2}}e^{-\frac{(1+A_{k})z^{2}_{m+k-d}}{2}}dz_{m+k-d}\cdots dz_{m+1-d},$ (3.59)
where
$A_{k-1}\ :=\ \frac{A_{k}^{2}}{1+A_{k}^{2}}+\frac{A_{k}}{(1+A_{k})^{2}}.$ (3.60)
Note that since $A_{k}>0$, then $A_{k+1}>0$. Hence, (3.56) becomes bounded above by
$\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{k-1}e^{-\frac{\sum_{i=1}^{k-1}z^{2}_{m+i-d}}{2}}\cdot e^{-\frac{A_{k-1}\left(y-\sum_{i=1}^{k-1}z_{m+i-d}\right)^{2}}{2}}\left(\int_{-\infty}^{\infty}e^{-\frac{(1+A_{k})z_{m+k-d}^{2}}{2}}dz_{m+k-d}\right)$
$\quad\quad\quad dz_{m+(k-1)-d}\cdots dz_{m+1-d}$
$\quad\ \ll\ \underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{k-1}e^{-\frac{\sum_{i=1}^{k-1}z^{2}_{m+i-d}}{2}}\cdot e^{-\frac{A_{k-1}\left(y-\sum_{i=1}^{k-1}z_{m+i-d}\right)^{2}}{2}}dz_{m+(k-1)-d}\cdots dz_{m+1-d},$ (3.61)
Thus, by induction, (3.6) becomes bounded above by
$e^{-(A_{0}y^{2})/2},$ (3.62)
where $A_{0}>0$ is recursively defined by (3.60) with $A_{d-2}=1/4$. Similarly, we can show by induction that the second integral in (3.6) has the estimate
$\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)\right)dz_{m-2}\cdots dz_{m+1-d}\ \ll\ e^{-(B_{0}y^{2})/2}$ (3.63)
for some $B_{0}>0$. Hence, if $d>2$, (3.6) becomes
$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)$ $\ \ll\ e^{-(A_{0}y^{2})/2}+e^{-(B_{0}y^{2})/2}\ \ll\ e^{-(C_{0}y^{2})/2},$ (3.64)
where $C_{0}:=\min\{A_{0},B_{0}\}$. Recall from (3.6) that when $d=2$,
$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)\ \ll\ \varphi\left(\frac{y}{2}\right)+\varphi\left(y+C\sqrt{N}\right).$ (3.65)
Thus, regardless of whether $d=2$ or $d>2$,
$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)\ \ll\ e^{-\frac{D_{0}y^{2}}{2}}+e^{-\frac{D_{0}\left(y+C\sqrt{N}\right)^{2}}{2}},$ (3.66)
where $D_{0}:=\min\{C_{0},1/4\}$. Note that $e^{-D_{0}y^{2}/2}$ and $\varphi(\sqrt{D_{0}}(y+C\sqrt{N}))$ each has only one global extreme, at $y=0$ and $y=-C\sqrt{N}$ respectively. Hence, locally on $n/\sqrt{N},(n+1)/\sqrt{N}]$ the functions are monotonic and can be bounded above by the sum of its values at the two end points. Hence, for all $n$,
$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)\ \ll\ e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}+C\sqrt{N}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}+C\sqrt{N}\right)^{2}}{2}},$ (3.67)
for all $y\in[n/\sqrt{N},(n+1)/\sqrt{N}]$. Thus, returning to (3.45), for all $y_{1},y_{2}\in[n/\sqrt{N},(n+1)/\sqrt{N}]$,
$\left|\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y_{1})-\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y_{2})\right|$
$\quad\ \ll\ \frac{1}{\sqrt{N}}\left(e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}+C\sqrt{N}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}+C\sqrt{N}\right)^{2}}{2}}\right).$ (3.68)
Thus, the error term $e_{n}(a,b)$ as defined in (3.6) is bounded by
$e_{n}(a,b)\ \leq\ \int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\left|\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)-\frac{d}{dy}\mathcal{M}_{Y^{(N)}}\left(\frac{n}{\sqrt{N}}\right)\right|dy$
$\quad\ \ll\ \left(\frac{1}{\sqrt{N}}\right)^{2}\left(e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}+C\sqrt{N}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}+C\sqrt{N}\right)^{2}}{2}}\right).$ (3.69)
3.7. Upper bound on error term $\mathcal{M}_{\textup{err},N}(a,b)$
In this section, we want to prove that the error term of the main term $\mathcal{M}_{N}(a,b)$ is small, i.e.,
$\mathcal{M}_{\textup{err,N}}(a,b)\ =\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}e_{n}(a,b)\ \approx\ 0.$ (3.70)
Based on the estimate $e_{n}(a,b)$ in (3.6), we can pull out one of the $1/\sqrt{N}$ factors in the bracket and obtain
$\mathcal{M}_{\textup{err},N}(a,b)\ =\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}e_{n}(a,b)$
$\ \ll\ \frac{1}{\sqrt{N}}\cdot\left(\frac{1}{\sqrt{N}}\sum_{n=-dC\cdot N}^{dC\cdot N-1}\left(e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}+C\sqrt{N}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}+C\sqrt{N}\right)^{2}}{2}}\right)\right).$ (3.71)
We observe that the above in the parentheses are Riemann sums for $e^{-D_{0}x^{2}/2}$ and $e^{-D_{0}(x+C\sqrt{N})^{2}/2}$ on $(-\infty,\infty)$, which are both finite. Hence
$\mathcal{M}_{\textup{err},N}(a,b)\ \ll\ \frac{1}{\sqrt{N}}.$ (3.72)
We have thus established that the error term $\mathcal{M}_{\textup{err},N}(a,b)$ of the main term $\mathcal{M}_{N}(a,b)$ is negligible.
3.8. Evaluation of main term $\mathcal{M}_{\textup{main},N}(a,b)$
Finally, we establish the main term $\mathcal{M}_{\textup{main},N}(a,b)$ of the main term $\mathcal{M}_{N}(a,b)$. From (3.37) and (3.42), we see that
$\mathcal{M}_{\textup{main},N}(a,b)$ $\ =\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}m_{n}(a,b)$
$\ =\ (b-a)\cdot\left(\frac{1}{\sqrt{N}}\sum_{n=-dC\cdot N}^{dC\cdot N-1}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}\left(\frac{n}{\sqrt{N}}\right)\right).$ (3.73)
Since the term in the parentheses above is a Riemann sum for $\int_{-\infty}^{\infty}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)dy$, we have
$\frac{1}{\sqrt{N}}\sum_{n=-2C\cdot N}^{2C\cdot N-1}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}\left(\frac{n}{\sqrt{N}}\right)$ $\ =\ \int_{-\infty}^{\infty}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)dy+o(1)$
$\ =\ \mathcal{M}_{Y^{(N)}}(\infty)-\mathcal{M}_{Y^{(N)}}(-\infty)+o(1).$ (3.74)
Hence the main term $\mathcal{M}_{\textup{main},N}(a,b)$ becomes
$\mathcal{M}_{\textup{main},N}(a,b)\ =\ (b-a)(\mathcal{M}_{Y^{(N)}}(\infty)-\mathcal{M}_{Y^{(N)}}(-\infty))+o(1).$ (3.75)
Recall that
$\mathcal{M}_{Y^{(N)}}(y)$ $\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}m_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}.$ (3.76)
To simplify our calculation, we want to extend the interval of integration of the outer integral to -$\infty$. We do so by showing that the tail of the integral is small, i.e.,
$T_{N}\ :=\ \int_{-\infty}^{-C\sqrt{N}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}m_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}\ \approx\ 0.$ (3.77)
We have
$T_{N}$ $\ \ll\ \int_{-\infty}^{-C\sqrt{N}}\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-1}(\Phi(z_{m+1-d}))^{m-d}\prod_{i=1}^{d}\varphi(z_{m+i-d})dz_{m}\cdots dz_{m+1-d}$
$\ \ll\ \int_{-\infty}^{-C\sqrt{N}}\varphi(z_{m+1-d})dz_{m+1-d},$ (3.78)
Hence, we have reduced the problem to showing the smallness of the Gaussian tail. We first make a definition.
Definition 3.5.
For functions $f(x)$ and $g(x)$, we say that $f(x)=\Theta(g(x))$ if $f(x)=O(g(x))$ and $g(x)=O(f(x))$.
The following result is a straightforward calculation that provides an upper bound on the Gaussian tail.
Proposition 3.6.
[Mil1] Suppose that $g:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ is a function such that $g(N)=\Theta(N^{\epsilon^{\prime}})$ for some $\epsilon^{\prime}>0$. Then
$\int_{-\infty}^{-g(N)}\frac{1}{\sqrt{2\pi}}e^{-x^{2}/2}dx\ =\ \int_{g(N)}^{\infty}\frac{1}{\sqrt{2\pi}}e^{-x^{2}/2}dx\ \ll\ e^{-g(N)/2}.$ (3.79)
Back to (3.8), by Proposition 3.6
$T_{N}\ \ll\ e^{-C\sqrt{N}}.$ (3.80)
Thus, (3.76) becomes
$\mathcal{M}_{Y^{(N)}}(y)$ $\ =\ \int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}m_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}+O\left(e^{-C\sqrt{N}}\right).$ (3.81)
Now, by Proposition 3.3,
$m_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})\ :=\ C_{m}^{m+1-d}\left(\Phi(z_{m+1-d})\right)^{m-d}\prod_{i=1}^{d}\varphi(z_{m+i-d})$ (3.82)
is the joint PDF of $W_{(m+1-d)},\dots,W_{(m)}$, where $W_{1},\dots,W_{m}$ are i.i.d. $\sim N(0,1)$. Thus, following our derivation in Subsection 3.2,
$H(y)\ :=\ \int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}m_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}$ (3.83)
is the CDF of $\sum_{i=1}^{d}W_{(m+i-d)}$. We know that for a cumulative density function $H(y)$, $H(-\infty)=0$ and $H(\infty)=1$. Hence,
$\mathcal{M}_{Y^{(N)}}(-\infty)$ $\ =\ H(-\infty)+O(e^{-C\sqrt{N}})\ =\ O\left(e^{-C\sqrt{N}}\right)$
$\mathcal{M}_{Y^{(N)}}(\infty)$ $\ =\ H(\infty)+O(e^{-C\sqrt{N}})\ =\ 1+O\left(e^{-C\sqrt{N}}\right).$ (3.84)
Thus, substituting these two estimates into (3.75) yields
$\mathcal{M}_{\textup{main},N}(a,b)\ =\ (b-a)\left(1+O\left(e^{-C\sqrt{N}}\right)\right)+o(1))\ =\ (b-a)+o(1).$ (3.85)
Returning to (3.39) and (3.38), combined with the estimate for the error term $\mathcal{E}_{N}(a,b)$ in (3.4) and the error term $\mathcal{M}_{\textup{err},N}(a,b)$ of $\mathcal{M}_{N}(a,b)$ in (3.72), we have
$F_{N}(a,b)$ $\ =\ \mathcal{M}_{N}(a,b)+\mathcal{E}_{N}(a,b)$
$\ =\ \mathcal{M}_{\textup{main},N}(a,b)+\mathcal{M}_{\textup{err},N}(a,b)+\mathcal{E}_{N}(a,b)$
$\ =\ (b-a)+o(1)+O\left(\frac{1}{\sqrt{N}}\right)+O(N^{-\delta})$
$\ =\ (b-a)+o(1).$ (3.86)
We conclude that $\log_{B}(\mathfrak{m}_{d}^{(N)})$ converges to being equidistributed mod 1, and therefore by Uniform Distribution Characterization $\mathfrak{m}_{d}^{(N)}$ converges to strong Benford behavior. ∎
Appendix A Proof of Proposition 2.5
In this section, we prove Proposition 2.5 which provides a quantitative bound on the difference among probabilities within an interval. Let us first recall the statement of Proposition 2.5.
Proposition 2.5.
For $\ell\leq\sqrt{k(N)}/2$,
$\left|\binom{k(N)}{k_{1,\ell}}-\binom{k(N)}{k_{1,\ell+1}}\right|\ \leq\ O\left(\binom{k(N)}{k_{1,\ell}}\cdot N^{-\frac{3\epsilon}{10}}\right).$ (A.1)
Proof.
We first factor out $\binom{k(N)}{k_{1,\ell}}$ from the difference:
$\binom{k(N)}{k_{1,\ell}}-\binom{k(N)}{k_{1,\ell+1}}\ =\ \binom{k(N)}{\frac{k(N)}{2}+\ell N^{\delta}}-\binom{k(N)}{\frac{k(N)}{2}+(\ell+1)N^{\delta}}$
$\ =\ \frac{k(N)!}{\left(\frac{k(N)}{2}+\ell N^{\delta}\right)!\left(\frac{k(N)}{2}-\ell N^{\delta}\right)!}-\frac{k(N)!}{\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)!\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)}$
$\ =\ \frac{k(N)!\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)!\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!-k(N)!\left(\frac{k(N)}{2}+\ell N^{\delta}\right)!\left(\frac{k(N)}{2}-\ell N^{\delta}\right)!}{\left(\frac{k(N)}{2}+\ell N^{\delta}\right)!\left(\frac{k(N)}{2}-\ell N^{\delta}\right)!\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)!\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!}$
$\ =\ \frac{k(N)!}{\left(\frac{k(N)}{2}+\ell N^{\delta}\right)!\left(\frac{k(N)}{2}-\ell N^{\delta}\right)!}$
$\cdot\frac{\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!-\left(\frac{k(N)}{2}+\ell N^{\delta}\right)!\left(\frac{k(N)}{2}-\ell N^{\delta}\right)!}{\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)!\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!}$
$\ =\ \binom{k(N)}{k_{1,\ell}}\left(1-\frac{\left(\frac{k(N)}{2}+\ell N^{\delta}\right)!\left(\frac{k(N)}{2}-\ell N^{\delta}\right)!}{\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)!\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!}\right).$ (A.2)
Now, we analyze the term
$\alpha_{\ell,N}\ :=\ \frac{\left(\frac{k(N)}{2}+\ell N^{\delta}\right)!\left(\frac{k(N)}{2}-\ell N^{\delta}\right)!}{\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)!\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!}.$ (A.3)
We want to show that $\alpha_{\ell,N}\rightarrow 1$ as $N\rightarrow\infty$, so that the difference in (A) is asymptotically much smaller than the main term $\binom{k(N)}{k_{1,\ell}}$. We have
$\frac{\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)^{N^{\delta}}}{\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)^{N^{\delta}}}$ $\ \leq\ \alpha_{\ell,N}\ \leq\ \frac{\left(\frac{k(N)}{2}-\ell N^{\delta}\right)^{N^{\delta}}}{\left(\frac{k(N)}{2}+\ell N^{\delta}\right)^{N^{\delta}}}$
$\left(1-\frac{4(\ell+1)N^{\delta}}{k(N)+2(\ell+1)N^{\delta}}\right)^{N^{\delta}}$ $\ \leq\ \alpha_{\ell,N}\ \leq\left(1-\frac{4\ell N^{\delta}}{k(N)+2\ell N^{\delta}}\right)^{N^{\delta}}.$ (A.4)
Since $\ell\leq\sqrt{k(N)}/2$, then
$0\ \leq\ \frac{4\ell N^{\delta}}{k(N)+2\ell N^{\delta}}\ \leq\ \frac{4\ell N^{\delta}}{k(N)}\ \leq\ \frac{2N^{\delta}}{\sqrt{k(N)}}.$ (A.5)
Since $k(N)\geq N^{\epsilon}$, and $\delta\in(0,\epsilon/10)$, then
$0\ \leq\ \frac{2N^{2\delta}}{\sqrt{k(N)}}$ $\ \leq\ \frac{2N^{\epsilon/10}}{N^{\epsilon/2}}\ =\ 2N^{-2\epsilon/5}\ =\ O(N^{-2\epsilon/5}).$ (A.6)
Similarly, we also have
$\frac{4(\ell+1)N^{\delta}}{k(N)+2(\ell+1)N^{\delta}}\ =\ O(N^{-2\epsilon/5}).$ (A.7)
Hence, for sufficiently large $N$,
[…]
Returning to (A), given (A), we have
$\left(1-\frac{4(\ell+1)N^{\delta}}{k(N)+2(\ell+1)N^{\delta}}\right)^{N^{\delta}}\ \leq\ \alpha_{\ell,N}\ \leq\ \left(1-\frac{4\ell N^{\delta}}{k(N)+2\ell N^{\delta}}\right)^{N^{\delta}}.$ (A.9)
Using binomial expansion,
$\sum_{j=1}^{N^{\delta}}\binom{N^{\delta}}{j}(-1)^{j}\left(\frac{4(\ell+1)N^{\delta}}{k(N)+2(\ell+1)N^{\delta}}\right)^{j}$ $\ \leq\ \alpha_{\ell,N}-1\ \leq\ \sum_{j=1}^{N^{\delta}}\binom{N^{\delta}}{j}(-1)^{j}\left(\frac{4\ell N^{\delta}}{k(N)+2\ell N^{\delta}}\right)^{j}.$ (A.10)
We first bound the right sum in (A). Using the assumption that $\ell\leq\sqrt{k(N)}/2$, we have
$\left|\sum_{j=1}^{N^{\delta}}\binom{N^{\delta}}{j}(-1)^{j}\left(\frac{4\ell N^{\delta}}{k(N)+2\ell N^{\delta}}\right)^{j}\right|$ $\ \leq\ \sum_{j=1}^{N^{\delta}}N^{j\delta}\left(\frac{4\ell N^{\delta}}{k(N)}\right)^{j}$
$\ \leq\ \sum_{j=1}^{N^{\delta}}N^{j\delta}\left(\frac{2\sqrt{k(N)}N^{\delta}}{k(N)}\right)^{j}$
$\ =\ \sum_{j=1}^{N^{\delta}}\left(\frac{2N^{2\delta}}{\sqrt{k(N)}}\right)^{j}$
$\ =\ \frac{2N^{2\delta}}{\sqrt{k(N)}}\cdot\frac{1-\left(\frac{2N^{2\delta}}{\sqrt{k(N)}}\right)^{N^{\delta}}}{1-\frac{2N^{2\delta}}{\sqrt{k(N)}}},$ (A.11)
where on the last line we use the geometric series formula. Since $k(N)\geq N^{\epsilon}$, and $\delta\in(0,\epsilon/10)$, then
$0\ \leq\ \frac{2N^{2\delta}}{\sqrt{k(N)}}$ $\ \leq\ \frac{2N^{\epsilon/5}}{N^{\epsilon/2}}\ =\ 2N^{-3\epsilon/10}.$ (A.12)
Hence
$\left|\sum_{j=1}^{N^{\delta}}\binom{N^{\delta}}{j}(-1)^{j}\left(\frac{4\ell N^{\delta}}{k(N)+2\ell N^{\delta}}\right)^{j}\right|$ $\ \leq\ 2N^{-3\epsilon/10}\cdot\frac{1}{1-2N^{-3\epsilon/10}}$
$\ =\ O\left(N^{-3\epsilon/10}\right).$ (A.13)
Similarly, for the left sum in (A), we also have
$\left|\sum_{j=1}^{N^{\delta}}\binom{N^{\delta}}{j}(-1)^{j}\left(\frac{4(\ell+1)N^{\delta}}{k(N)+2(\ell+1)N^{\delta}}\right)^{j}\right|\ =\ O\left(N^{-3\epsilon/10}\right).$ (A.14)
Applying (A) and (A.14) to (A), we get
$\left|\alpha_{\ell,N}-1\right|\ =\ O\left(N^{-3\epsilon/10}\right).$ (A.15)
Thus, substituting the estimate (A.15) back to (A) gives us
$\left|\binom{k(N)}{k_{1,\ell}}-\binom{k(N)}{k_{1,\ell+1}}\right|\ \leq\ O\left(\binom{k(N)}{k_{1,\ell}}\cdot N^{-3\epsilon/10}\right).$ (A.16)
Appendix B Case for $k_{S}<d$
In this appendix, we want to show that $D_{1}(y)\ll N^{-1/2-\delta}$ when $k_{S}<d$. Recall that $S$ is a proper subset of $[d]$ and $k_{S}$ is the largest index such that $k\in[d]\setminus S$, and that $D_{1}(y)$ is defined in (3.4) to be
$D_{1}(y)$ $\ :=\ \sum_{\begin{subarray}{c}S\subset[d]\\ S\neq[d]\end{subarray}}\int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\Bigg{|}_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}$
$\quad\quad dz_{m-1}\cdots dz_{m+1-d}.$ (B.1)
Let $\ell_{1};\ell_{2}]$ denote $\{\ell_{1},\dots,\ell_{2}\}$ if $\ell_{1},\ell_{2}$ are integers such that $\ell_{1}\leq\ell_{2}$, and let it be $\emptyset$ otherwise. Then the integrand of (3.4) becomes
$\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\Bigg{|}_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}$
$\quad\ =\ \left(\prod_{i\in S\setminus[k_{S}+1;d]}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus(S\cup\{k_{S}\})}A(z_{m+i-d})\right)\cdot A(z_{m+k_{S}-d})$
$\quad\quad\quad\cdot\left(\prod_{i\in[k_{S}+1;d-1]}\varphi(z_{m+i-d})\right)\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right).$ (B.2)
Hence,
$\int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\Bigg{|}_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}$
$dz_{m-1}\cdots dz_{m+1-d}$
$\quad\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S\setminus[k_{S}+1;d]}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus(S\cup\{k_{S}\})}A(z_{m+i-d})\right)$
$\quad\quad\quad\cdot A(z_{m+k_{S}-d})\cdot\left(\prod_{i=k_{S}+1}^{d-1}\varphi(z_{m+i-d})\right)\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\cdots dz_{m+1-d}.$ (B.3)
To bound (B), we first give some estimate on the following
$\prod_{j=k_{S}+1}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i=k_{S}+1}^{d-1}\varphi(z_{m+i-d})\right)\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\cdots dz_{m+(k_{S}+1)-d}$
$\quad\ \ll\ \underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-k_{S}-1}\left(\prod_{i=k_{S}+1}^{d-1}\varphi(z_{m+i-d})\right)\cdot\varphi\left(\left(y-\sum_{i=1}^{k_{S}}z_{m+i-d}\right)-\sum_{i=k_{S}+1}^{d-1}z_{m+i-d}\right)dz_{m-1}$
$\quad\quad\quad\cdots dz_{m+(k_{S}+1)-d}$
$\quad\ \ll\ \frac{1}{\sqrt{2\pi(d-k_{S})}}e^{-\frac{\left(y-\sum_{i=1}^{k_{S}}z_{m+i-d}\right)^{2}}{2(d-k_{S})}},$ (B.4)
where in the last line we use the fact that the second last line is exactly the convolution of $d-k_{S}$ standard Gaussian density function evaluated at $y-\sum_{i=1}^{k_{S}}z_{m+i-d}$, which is exactly the probability density function of sum of $d-k_{S}$ independent standard Gaussian random variables and thus is itself also Gaussian with mean 0 and variance $d-k_{S}$, evaluated at $y-\sum_{i=1}^{k_{S}}z_{m+i-d}$. Hence, (B) is bounded above by
$\int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{k_{S}-1}\int_{z_{m}+(j-1)-d}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S\setminus[k_{S}+1;d]}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus(S\cup\{k_{S}\})}A(z_{m+i-d})\right)$
$\quad\quad\int_{z_{m}+(k_{S}-1)-d}^{\frac{y-\sum_{i=1}^{k_{S}-1}z_{m+i-d}}{d-k_{S}+1}}A(z_{m+k_{S}-d})\cdot\frac{1}{\sqrt{2\pi(d-k_{S})}}e^{-\frac{\left(y-\sum_{i=1}^{k_{S}}z_{m+i-d}\right)^{2}}{2(d-k_{S})}}dz_{m+k_{S}-d}\cdots dz_{m+1-d}$
$\quad\ \ll\ N^{-\frac{1}{2}-\delta}\underbrace{\int_{-C^{\prime}\sqrt{N}}^{C^{\prime}\sqrt{N}}\cdots\int_{-C^{\prime}\sqrt{N}}^{C^{\prime}\sqrt{N}}}_{k_{S}-1}\left(\prod_{i\in S\setminus[k_{S}+1;d]}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus(S\cup\{k_{S}\})}A(z_{m+i-d})\right)$
$\quad\quad\quad dz_{m+(k_{S}-1)-d}\cdots dz_{m+1-d}$
$\quad\ \ll\ N^{-\frac{1}{2}-\delta}\prod_{i\in S\setminus[k_{S}+1;d]}\left(\int_{-C^{\prime}\sqrt{N}}^{C^{\prime}\sqrt{N}}\varphi(z_{m+i-d})dz_{m+i-d}\right)\cdot\prod_{i\in[d]\setminus(S\cup\{k_{S}\})}\left(\int_{-C^{\prime}\sqrt{N}}^{C^{\prime}\sqrt{N}}A(z_{m+i-d})dz_{m+i-d}\right)$
$\quad\ \ll\ N^{-1/2-\left(1+\left|[d]\setminus(S\cup\{k_{S}\})\right|\right)\delta}\ \ll\ N^{-1/2-\delta},$ (B.5)
where in the second line, we use the change of variable $z_{m+k_{S}-d}\mapsto z_{m+k_{S}-d}-(\sum_{i=1}^{k_{S}-1}z_{m+i-d}+y)$ and the fact that
[…]
Since the number of proper subsets $S$ of $[d]$ is finite, then when $k_{S}<d$, we have that $D_{1}(y)\ll N^{-1/2-\delta}$ by definition of $D_{1}(y)$ in (B).
Appendix C Case for $d>2$
In this section, we want to obtain the following estimate on $d^{2}\mathcal{M}_{Y^{(N)}}(y)/dy^{2}$ when $d>2$
$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)$
$\quad\ \ll\ \underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)dz_{m-2}\cdots dz_{m+1-d}$
$\quad\quad+\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)dz_{m-2}\cdots dz_{m+1-d}.$ (C.1)
First, recall from (3.6) that
$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)$
$\quad\ \ll\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-2}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\Phi(z_{m+1-d})\right)^{m-d}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\right)^{2}$
$\quad\quad+\Bigg{(}\int_{z_{m-2}}^{\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}}(\Phi(z_{m+1-d}))^{m-d}\left(\prod_{i=1}^{d-1}\varphi(z_{m+i-d})\right)(-1)\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
$\quad\quad\quad\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\Bigg{)}dz_{m-2}\cdots dz_{m+1-d}.$ (C.2)
Since $y$ is an upper bound on $\sum_{i=1}^{d}z_{m+i-d}$ and $z_{m-2}\leq z_{m-1},z_{m}$, then we have
$z_{m-2}\ \leq\ \frac{z_{m-1}+z_{m}}{2}\ \leq\ \frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}.$ (C.3)
When $(y-\sum_{i=1}^{d-2}z_{m+i-d})/2\leq 0$, $\varphi(z_{m-1})$ is at most $\varphi((y-\sum_{i=1}^{d-2}z_{m+i-d})/2)$. Hence,
$\int_{z_{m-2}}^{\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}}(\Phi(z_{m+1-d}))^{m-d}\left(\prod_{i=1}^{d-1}\varphi(z_{m+i-d})\right)(-1)\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
$\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}$
$\quad\ \ll\ \left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\int_{-\infty}^{\infty}\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
$\quad\quad\quad\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}$
$\quad\ll\ \left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\int_{-\infty}^{\infty}|z_{m-1}|\varphi(z_{m-1})dz_{m-1},$ (C.4)
where in the last line we use the change of variable $z_{m-1}\mapsto-z_{m-1}+y-\sum_{i=1}^{d-2}z_{m+i-d}$. The integral $\int_{-\infty}^{\infty}|x|\varphi(x)dx$ is the expected value of the absolute value of a standard normal random variable, which is finite. Hence, when $(y-\sum_{i=1}^{d-2}z_{m+i-d})/2\leq 0$,
$\int_{z_{m-2}}^{\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}}(\Phi(z_{m+1-d}))^{m-d}\left(\prod_{i=1}^{d-1}\varphi(z_{m+i-d})\right)(-1)\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
$\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}$
$\quad\ \ll\ \left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right).$ (C.5)
On the other hand, when $(y-\sum_{i=1}^{d-2}z_{m+i-d})/2\geq 0$,
$y-\sum_{i=1}^{d-1}z_{m+i-d}\ \geq\ y-\sum_{i=1}^{d-2}z_{m+i-d}-\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\ =\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\ \geq\ 0.$ (C.6)
Hence, we find
$\int_{z_{m-2}}^{\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}}(\Phi(z_{m+1-d}))^{m-d}\left(\prod_{i=1}^{d-1}\varphi(z_{m+i-d})\right)(-1)\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
$\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}$
$\quad\ \ll\ \left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\int_{z_{m-2}}^{\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}}\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}$
$\quad\ =\ \left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)\Bigg{|}_{z_{m-1}=z_{m-2}}^{z_{m-1}=\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}}$
$\quad\ =\ \left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)-\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)\right).$ (C.7)
Thus, if $d>2$, regardless of whether $(y-\sum_{i=1}^{d-2}z_{m+i-d})/2\leq 0$ or $(y-\sum_{i=1}^{d-2}z_{m+i-d})/2\geq 0$, we have by combining (C) and (C) that
$\int_{z_{m-2}}^{\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}}(\Phi(z_{m+1-d}))^{m-d}\left(\prod_{i=1}^{d-1}\varphi(z_{m+i-d})\right)(-1)\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
$\quad\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}$
$\quad\ \ll\ \left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)+\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)\right).$ (C.8)
Hence, (3.6) becomes
$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)$
$\quad\ \ll\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-2}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\right)^{2}$
$\quad\quad\quad+\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)+\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)\right)$
$\quad\quad\quad dz_{m-2}\cdots dz_{m+1-d}$
$\quad\ \ll\ \underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)dz_{m-2}\cdots dz_{m+1-d}$
$\quad\quad+\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)dz_{m-2}\cdots dz_{m+1-d},$ (C.9)
which is exactly the estimate we need.

### v3

@@
 Benford Behavior Resulting From Stick and Box Fragmentation Processes
-Benford behavior resulting from stick and box fragmentation processes
-Bruce Fang fangbaojun2002@gmail.com Department of Mathematics and Statistics, Williams College, Williamstown, MA 01267  and  Steven J.
-Miller sjm1@williams.edu, Steven.Miller.MC.96@aya.yale.edu Department of Mathematics and Statistics, Williams College, Williamstown, MA 01267
-(Date: August 19, 2025)
-Benford’s law is the statement that in many real-world data set, the probability of having digit $d$ in base $B$, where $1\leq d\leq B$, as the first digit is $\log_{B}\left((d+1)/d\right)$.
-We sometimes refer to this as weak Benford behavior, and we say that a data set exhibits strong Benford behavior in base $B$ if the probability of having significand at most $s$, where $s\in[1,B)$, is $\log_{B}(s)$.
-We examine Benford behaviors in two different probabilistic model: stick and box fragmentation.
-Building on the joint work of [B–] on the 1-dimensional stick fragmentation model, we employ combinatorial identities on multinomial coefficients to reduce the high dimensional stick fragmentation model to the 1-dimensional model and provide a necessary and sufficient condition for the lengths of the stick fragments to converge to strong Benford behavior.
-Then we answer a conjecture posed by [BDMMM] on the high dimensional box fragmentation model.
-Using tools from order statistics, we prove that under some mild conditions, faces of any arbitrary dimension of the box have total volume converging to strong Benford’s behavior.
-1. Introduction
-At the beginning of the $20$th century, astronomer and mathematician Simon Newcomb [Ne] noticed an unusual pattern in the logarithmic tables he used at work.
+Abstract
+Benford’s law is the statement that in many real-world data set, the probability of having digit $d$ in base $B$ as the first digit is $\log_{B}\left((d+1)/d\right)$ for all $1\leq d\leq B$.
+We sometimes refer to this as weak Benford behavior, and we say that a data set satisfies strong Benford behavior in base $B$ if the probability of having significand at most $s$ is $\log_{B}(s)$ for all $s\in[1,B)$.
+We examine Benford behaviors in two different probabilistic model: stick and box fragmentation models.
+Building on the joint work of Becker et al.
+[1] on the single proportion stick fragmentation model, we employ combinatorial identities on multinomial coefficients to reduce the multi-proportion stick fragmentation model to the single proportion model.
+We then provide a necessary and sufficient condition for the lengths of the stick fragments to converge to strong Benford behavior along with a quantification of the discrepancy from uniform distribution on $[0,1]$ in terms of irrationality exponent.
+Then we answer a conjecture posed by Betti et al.
+[6] on the high dimensional box fragmentation model.
+Using tools from Fourier analysis and order statistics, we prove that under some conditions, faces of any arbitrary dimension of the box have total volume converging to strong Benford’s behavior.
+keywords:
+Benford’s law , high-dimensional fragmentation , equidistribution mod 1 , irrationality exponent , multinomial identities , order statistics
+2008 MSC:
+60A10 , 11K06 (primary) , 60E10 (secondary)
+journal: Journal of Number Theory
+[…]
+[inst1]organization=Department of Mathemtics, addressline=Williams College, city=Williamstown, postcode=01267, state=MA, country=USA
+1 Introduction
+At the end of the 19th century, astronomer and mathematician Simon Newcomb [18] noticed an unusual pattern in the logarithmic tables he used at work.
 The early pages of the books were far more worn than the later ones, suggesting that numbers starting with smaller digits were consulted more often.
-From this, Newcomb inferred that people had a tendency to encounter numbers beginning with $1$ more frequently.
-Specifically, he found that $1$ appeared as the leading digit about $30\%$ of the time, $2$ about $17\%$, with the frequency decreasing for larger digits.
-Although he formulated a mathematical explanation for this curious phenomenon, his discovery initially went largely overlooked.
-It took another 57 years after Newcomb’s discovery for physicist Frank Benford to make the exact same observation as Newcomb: the first pages of logarithmic tables were used far more than others.
+Specifically, Newcomb found that $1$ appeared as the leading digit for about $30\%$ of the time, $2$ for about $17\%$, with the frequency decreasing for larger digits.
+Although he formulated a mathematical law for this curious phenomenon, his discovery initially went largely overlooked.
+It took over 50 years after Newcomb’s discovery for physicist Frank Benford to make the same observation as Newcomb.
 He formulated this law as follows.
-Definition 1.1.
-[Ben, Page 554] Data exhibits (weak) Benford behavior base $B$ if the frequency $F_{d}$ of leading digit $d$ is
-$F_{d}\ =\ \log_{B}\left(\frac{d+1}{d}\right).$ (1.1)
-Nowadays, Benford’s Law is used in detecting many different forms of fraud, and its prevalence in the world fascinates not only mathematicians, but many other scientists as well (to learn more about Benford’s law and its many applications, we recommend [BeHi, Nig, Mil1, NiMi] to name a few).
-In 1986, Lemons [Lem] proposed using Benford’s law to analyze the partitioning of a conserved quantity.
-Since then, driven by the potential application to nuclear fragmentation, mathematicians and physicists have taken an interest in the Benford behavior of various fragmentation processes.
-Among these processes of interest is stick fragmentation.
-In a 1-dimensional stick fragmentation model, one begins with a stick of length $L$.
+Although the law was originally only stated for base 10, we state the generalized form for any base $B\geq 2$.
+Definition 1.
+[5, Page 554] A data set is said to satisfy Benford’s law for the leading digit in base $B$ if the frequency $F_{d}$ of leading digit $d$ is
+$F_{d}\ =\ \log_{B}\left(\frac{d+1}{d}\right).$ (1)
+Nowadays, Benford’s Law is used in detecting many different forms of fraud, and its prevalence in the world fascinates not only mathematicians, but many other scientists as well (to learn more about Benford’s law and its many applications, we recommend [3, 17, 19] to name a few).
+In 1986, Lemons [16] proposed applying Benford’s law to the analysis of the partitioning of a conserved quantity.
+Motivated in part by potential applications to nuclear fragmentation, both mathematicians and physicists have since investigated the Benford behavior of a range of fragmentation models.
+One such model is stick fragmentation.
+In a 1-dimensional stick fragmentation model, we begin with a stick of length $L$.
 Draw $p_{1}$ from a probability distribution on $(0,1)$.
-This fragments the stick into two sub-sticks of lengths $p_{1}L$ and $(1-p_{1})L$.
-For each sub-stick, draw two probabilities $p_{2}$ and $p_{3}$ from the same distribution.
-We perform this fragmentation process for a total of $N$ stages, where at each stage we fragment all the sub-sticks created during the previous stage according to probabilities.
-Hence, by the end of $N$ stages, there are $2^{N}$ sticks.
-Of particular interest is whether or not this fragmentation process converges to Benford’s behavior.
-1.1. Previous Work on Fragmentation
-An important definition when studying a more precise statistical version of Benford behavior is the notion of the significand of a real number, i.e., its leading digits in scientific notation.
-Definition 1.2.
-Given a positive real number $x$, we say that its significand base $B>1$, denoted $S_{B}(x)$, is the unique real number $S_{B}(x)\in[1,B)$ such that $x=S_{B}(x)\cdot B^{k}$, where $k$ is an integer.
-As is common practice with these techniques involving proofs of Benford behavior, we define a more general version of Benford behavior that allows for processes that do not necessarily exhibit Benford behavior at first but converge to it in the limit.
-Definition 1.3.
-We say that a sequence of random variables $X^{(n)}$ converges to strong Benford behavior base $B$ if and only if
-$\mathbb{P}(S_{B}(X^{(n)})\leq s)\to\log_{B}(s)$ (1.2)
-for all $s\in[1,B]$. Notice by compactness that this implies uniform convergence of (1.2).
-An equivalent formulation to the above is the Uniform Distribution Characterization, which is especially suited for investigation of products of random variables; see [Dia] for a proof.
-Proposition 1.4 (Uniform Distribution Characterization).
-[Dia] A sequence of random variables $X^{(n)}$ converges to strong Benford behavior in base $B$ if and only if
-$\mathbb{P}\left(\log_{B}(X^{(n)})\textup{ mod }1\leq t\right)\rightarrow t,$ (1.3)
-for all $t\in[0,1]$. If (1.4) is satisfied, then we say that $\log_{B}(X^{(n)})$ converges to being equidistributed mod 1.
-We may now state the previous results on stick fragmentation. Becker et al. [B–] proved a theorem regarding fixed proportion 1-dimensional stick fragmentation (compare with their Theorem 1.11).
-Theorem 1.5.
-[B–] Choose any $p\in(0,1)$.
-In Stage 1, cut a given stick at proportion $p$ to create two pieces of length $p$ and 1p1-p.
-In Stage 2, cut each resulting piece into two pieces at the same proportion $p$.
-Perform this process for a total of $N$ stages, generating $2^{N}$ sticks with $N+1$ distinct lengths (assuming $p\neq 1/2$) given by
+This divides the stick into two sub-sticks of lengths $p_{1}L$ and $(1-p_{1})L$.
+For each resulting sub-stick, draw two probabilities $p_{2}$ and $p_{3}$ from the same distribution and fragment them accordingly.
+This procedure is repeated for a total of $N$ stages, with every sub-stick produced at a given stage being fragmented at the next stage using newly drawn probabilities.
+After $N$ stages, the process yields $2^{N}$ sticks.
+A central question is whether the distribution of stick lengths produced by this process converges to Benford’s behavior.
+1.1 Previous Work on Fragmentation
+An key definition in formulating a more general version of Benford behavior is the notion of the significand of a real number, i.e., its leading digits in scientific notation.
+Definition 2.
+For any $x>0$, we can uniquely write
+$x\ =\ S_{B}(x)\cdot B^{k_{B}(x)},$ (2)
+where $S_{B}(x)\in[1,B)$ and $k_{B}(x)\in\mathbb{Z}$. Equivalently, $k_{B}(x)=\lfloor\log_{B}(x)\rfloor$ and $S_{B}(x)=x/B^{k_{B}(x)}$. We call $S_{B}(x)$ the significand of $x$.
+We now define a more general version of Benford behavior that captures processes that only display Benford behaviors in the limit.
+Definition 3.
+A sequence of random variables $\{X^{(N)}\}_{N=1}^{\infty}$ is said to converge to strong Benford behavior in base $B$ if
+$\lim_{N\rightarrow\infty}\mathbb{P}(S_{B}(X^{(N)})\leq s)\ =\ \log_{B}(s)$ (3)
+for all $s\in[1,B)$..
+An equivalent formulation to the above is the Uniform Distribution Characterization, which is especially suited for investigation of products of random variables; see [8] for a proof.
+Proposition 1 (Uniform Distribution Characterization).
+[8] A sequence of random variables $\{X^{(N)}\}_{N=1}^{\infty}$ converges to strong Benford behavior in base $B$ if and only if
+$\lim_{N\rightarrow\infty}\mathbb{P}\left(\log_{B}(X^{(N)})\textup{ mod }1\leq t\right)\ =\ t,$ (4)
+for all $t\in[0,1)$. If (4) is satisfied, then we say that $\{\log_{B}(X^{(N)})\}_{N=1}^{\infty}$ converges to being equidistributed mod 1.
+We are now ready to review some previous results on stick fragmentation models and introduce the two models of our paper.
+The first model is an extension of the model introduced by Becker et al.
+called the fixed single proportion stick fragmentation model [1].
+Their main result on this model is a necessary and sufficient condition for the distribution of stick lengths to converge to strong Benford behavior as well as a quantification of the error term in terms of irrationality exponent, which we define now.
+Definition 4.
+[15] Suppose that $x$ is a real number.
+The irrationality exponent $\mu$ of $x$ is the supremum of the set of $\mu$ such that $0<|x-p/q|<1/q^{\mu}$ is satisfied by an infinite number of coprime integer pairs $(p,q)$ with $q>0$.
+If such a set does not exist, then we say $x$ has irrationality exponent $\infty$.
+Definition 5.
+[15] For a finite sequence $\{x_{i}\}_{i=1}^{N}$, denote by $A([a,b),N)$ the number of $x_{i}$’s such that $x_{i}\textup{ mod 1}\in[a,b)$ for $0\leq a<b\leq 1$. Then we call the number
+$D_{N}\ :=\ \sup_{0\leq a<b\leq 1}\left|\frac{A([a,b),N)}{N}-(b-a)\right|$ (5)
+the discrepancy of the sequence. It measures how close a sequence is from being uniform mod 1.
+Now we are ready to state a main theorem of [1].
+Theorem 1.
+[1] Fix a proportion $p\in(0,1)$ and a stick of length $L$.
+At Stage 1, cut the stick at proportion $p$ to create two sub-sticks of lengths $Lp$ and L(1p)L(1-p).
+At Stage 2, cut each of these two sub-sticks into two smaller sub-sticks at the same proportion $p$.
+Repeat this procedure for a total of $N$ stages, generating $2^{N}$ sticks with $N+1$ distinct lengths (assuming $p\neq 1/2$) given by
 $x_{1}$ $\ =\ Lp^{N}$
@@
 $x_{N}$ $\ =\ Lp(1-p)^{N-1}$
-$x_{N+1}$ $\ =\ L(1-p)^{N},$ (1.4)
+$x_{N+1}$ $\ =\ L(1-p)^{N},$ (6)
 where the frequency of $x_{n}$ is $\binom{N}{n}/2^{N}$.
+Consider the sequence of random lengths $\{X^{(N)}\}_{N=1}^{\infty}$, where $X^{(N)}$ is the length of a stick picked uniformly at random from the $2^{N}$ sticks obtained at stage $N$.
 Choose $y$ so that $B^{y}=(1-p)/p$, which is the ratio of adjacent lengths (i.e., $x_{i+1}/x_{i}$).
-The decomposition process results in stick lengths that converge to strong Benford’s behavior base $B$ if and only if $y\not\in\mathbb{Q}$.
-If $y$ has finite irrationality exponent, the convergence rate can be quantified in terms of that exponent, and there is a power savings.
-Theorem 1.5 suggests the possibility of examining another fixed proportion 1-dimensional stick fragmentation model, where we cut a piece into an arbitrary $m\geq 2$ number of pieces according m1m-1 fixed proportions at each stage.
-Using new techniques involving multinomial coefficients to reduce the problem into the original fixed proportion 1-dimensional stick fragmentation model in [B–], this paper establishes the following result (see Section 2).
-Theorem 1.6.
-For any integer $m>2$, choose $p_{1},p_{2},\dots,p_{m-1}\in(0,1)$ such that $p_{1}+p_{2}+\dots+p_{m-1}<1$.
+Then $\{X^{(N)}\}_{N=1}^{\infty}$ converges to strong Benford behavior base in $B$ if and only if $y\not\in\mathbb{Q}$.
+If $y$ has a finite irrationality exponent, then for sufficiently large $N$, the discrepancy of the sequence of the $2^{N}$ stick lengths (not distinct) can be quantified in terms of that exponent, and there is a power savings.
+Theorem 1 suggests the possibility of generalizing the fixed single proportion stick fragmentation model to a fixed multi-proportion stick fragmentation model, where at each stage, instead of cutting a stick into 2 sub-sticks according to a fixed proportion, we cut a stick into an arbitrary but fixed $m$ number of sub-sticks according to m1m-1 fixed proportions.
+Again, we are interested in the conditions under which the distribution of stick lengths produced in this model converges to strong Benford’s behavior.
+A main result of this paper provides a definitive answer to this question, as follows:
+Theorem 2.
+For a fixed integer $m>2$, choose $p_{1},p_{2},\dots,p_{m-1}\in(0,1)$ such that $p_{1}+p_{2}+\dots+p_{m-1}<1$.
 Set $p_{m}:=1-(p_{1}+p_{2}+\dots+p_{m-1})$.
-At each stage, we cut all sticks according to proportions $p_{1},p_{2},\dots,p_{m-1}$ to create $m$ pieces.
+Now, consider a stick of length $L$.
+At each stage, we cut every remaining stick into $m$ sub-sticks according to proportions $p_{1},p_{2},\dots,p_{m-1}$.
 After stage $N$, we have $m^{N}$ sticks in total, of lengths
-$A_{k_{1},k_{2},\dots,k_{m}}\ =\ p_{1}^{k_{1}}p_{2}^{k_{2}}\cdots p_{m}^{k_{m}},$ (1.5)
-for $0\leq k_{1},k_{2},\dots,k_{m}\leq N$ such that $k_{1}+k_{2}+\cdots+k_{m}=N$.
-Let $y_{i}=\log_{B}(p_{i}/p_{i+1})$ for $1\leq i\leq m-1$.
-Then the decomposition process results in stick lengths that converge to strong Benford’s behavior base $B$ if and only if $y_{i}\not\in\mathbb{Q}$ for some $1\leq i\leq m-1$.
-Another possibility is to consider fragmentation processes in higher dimensions.
-This was considered by Durmić and Miller in [DM] and Benford behavior was established in the case of volume of an arbitrary $m$-dimensional box under mild assumptions.
-This was generalized by Betti et al.
-in [BDMMM] to any $d$-dimensional volume of $d$-dimensional faces of an arbitrary $m$-dimensional box, where $d\leq m$.
+$A^{(N)}_{k_{1},k_{2},\dots,k_{m}}\ =\ Lp_{1}^{k_{1}}p_{2}^{k_{2}}\cdots p_{m}^{k_{m}},$ (7)
+for all $0\leq k_{1},k_{2},\dots,k_{m}\leq N$ such that $k_{1}+k_{2}+\cdots+k_{m}=N$.
+Consider the sequence of random lengths $\{X^{(N)}\}_{N=1}^{\infty}$, where $X^{(N)}$ is the length of a stick picked uniformly at random from the $m^{N}$ sticks obtained at stage $N$.
+Let $y_{i}=\log_{B}(p_{i}/p_{i+1})$ for all $1\leq i\leq m-1$.
+Then $\{X^{(N)}\}_{N=1}^{\infty}$ converges to strong Benford behavior base in $B$ if and only if $y_{i}\not\in\mathbb{Q}$ for some $1\leq i\leq m-1$.
+Let $\kappa_{0}$ be the least irrationality exponent among all the irrational $y_{i}$’s.
+Then for sufficiently large $N$, the discrepancy of the sequence of the $m^{N}$ stick lengths (not distinct) is $O(N^{\delta(-1/(\kappa_{0}-1)+\epsilon^{\prime})})$ for some $\delta>0$ and for every $\epsilon^{\prime}>0$.
+In [1], Becker et al.
+also looked into other variants of the stick fragmentation model, in particular, the Unrestricted 1-Dimension Decomposition Model, where the proportion cuts are independently drawn from the same unspecified probability density $f$.
+They were able to establish a sufficient condition in terms of the Mellin transforms of the probability density function $f$ for strong Benford behavior to emerge from the stick length distribution.
+This direction was generalized by Durmić and Miller to a higher dimensional model called the box fragmentation model in [10], and strong Benford behavior was established for the volume of a fragmented box of arbitrary dimension $m$ under a similar condition on the Mellin transforms of $f$.
+In [6], Betti et al.
+further generalized the work of Durmić and Miller on box fragmentation model and considered the volume of the lower $d$-dimensional faces of a fragmented box of arbitrary dimension $m$, where $d\leq m$.
 To state their results, we start with some definitions.
-Definition 1.7.
-We say that a set $\mathfrak{B}\subset\mathbb{R}^{m}$ is an $m$-dimensional box if it is a set of the form $a_{1},b_{1}]\times\cdots\times[a_{m},b_{m}]\subset\mathbb{R}^{m}$, where $a_{i}<b_{i}$ are finite numbers.
-Definition 1.8.
+Definition 6.
+A set $\mathfrak{B}\subset\mathbb{R}^{m}$ is said to be an $m$-dimensional box if it is a set of the form $a_{1},b_{1}]\times\cdots\times[a_{m},b_{m}]\subset\mathbb{R}^{m}$, where $a_{i}<b_{i}$ are finite numbers.
+Definition 7.
 A linear-fragmentation process is a sequence of random variables $\mathfrak{B}_{0},\mathfrak{B}_{1},$ $\mathfrak{B}_{2},\dots$ such that the following holds.
-  1. (1)
+  1. 1.
     The random variables $\mathfrak{B}_{i}$ are $m$-dimensional boxes.
-  2. (2)
+  2. 2.
     The random variables $\mathfrak{B}_{i}$ form a descending chain $\mathfrak{B}_{0}\supset\mathfrak{B}_{1}\supset\mathfrak{B}_{2}\supset\dots$.
-  3. (3)
+  3. 3.
     The distribution $\mathfrak{B}_{n+1}$ conditioned on $\mathfrak{B}_{n}$ is some fixed distribution of independent proportion cuts $P_{1},\dots,P_{m}$ along each Cartesian axis.
 These $P_{i}$ are fixed over all $n\geq 0$.
-  4. (4)
+  4. 4.
     The proportion cuts $P_{i}$ are continuous random variables with finite mean, variance, and third moment.
-  5. (5)
+  5. 5.
     $\mathbb{E}[\log_{B}P_{i}]=\mu_{P}\in\mathbb{R}$ and $\textup{Var}[\log_{B}P_{i}]=\sigma_{P}^{2}>0$ for all $1\leq i\leq m$.
-Definition 1.9.
+Definition 8.
 Given an $m$-dimensional box $\mathfrak{B}$ and a positive integer $d\leq m$, we say the $d$-volume of $\mathfrak{B}=\prod_{i}[a_{i},b_{i}]$ is the sum of the $d$-dimensional volumes of the $d$-dimensional faces of $\mathfrak{B}$:
-$\textup{Vol}_{d}(\mathfrak{B})\ :=\ 2^{m-d}\sum_{|I|=d}\prod_{i\in I}(b_{i}-a_{i}),$ (1.6)
+$\textup{Vol}_{d}(\mathfrak{B})\ :=\ 2^{m-d}\sum_{|I|=d}\prod_{i\in I}(b_{i}-a_{i}),$ (8)
 where we are summing over all subsets $I\subset\{1,\dots,m\}$ with cardinality $d$.
-[BDMMM] established a sufficient condition for strong Benford behavior to emerge (see their Theorem 1.9), which involves the maximum $d$-dimensional volume of a $d$-dimensional face.
-Theorem 1.10 (Maximum Criterion).
-[BDMMM] Let $\mathfrak{B}:=\mathfrak{B}_{0}$ be a fixed $m$-dimensional box and $\mathfrak{B}_{0}\supset\mathfrak{B}_{1}\supset\cdots$ be a linear-fragmentation process whose proportion cuts $P_{i}$ have probability density functions $f_{i}:(0,1)\rightarrow(0,\infty)$.
+[6] established a sufficient condition for strong Benford behavior to emerge (see their Theorem 1.9), which involves the volume of the largest $d$-dimensional face.
+Theorem 3 (Maximum Criterion).
+[6] Let $\mathfrak{B}:=\mathfrak{B}_{0}$ be a fixed $m$-dimensional box and $\mathfrak{B}_{0}\supset\mathfrak{B}_{1}\supset\cdots$ be a linear-fragmentation process with proportion cuts $P_{i}$ having probability density functions $f_{i}:(0,1)\rightarrow(0,\infty)$.
 Let
-$V_{d}^{(N)}\ :=\ \textup{Vol}_{d}(\mathfrak{B}_{N})$ (1.7)
+$V_{d}^{(N)}\ :=\ \textup{Vol}_{d}(\mathfrak{B}_{N})$ (9)
 be the sequence of volumes obtained from this process and $\mathfrak{m}_{d}^{(N)}$ denote the maximum product of $d$ sides of $\mathfrak{B}_{N}$.
-If $\mathfrak{m}_{d}^{(N)}$ converges to strong Benford behavior base $B$ as $N\rightarrow\infty$, then so too does $V_{d}^{(N)}$ converge to strong Benford behavior base $B$ as $N\rightarrow\infty$.
-In the same paper, the Maximum Criterion was verified for $d=m$ for any linear fragmentation process, and for $d=1$ in a special case.
-Theorem 1.11.
-Let $P_{i}^{(j)}$ be i.i.d. log-uniform distributions, i.e., $\log_{B}P_{i}^{(j)}$’s have uniform distributions. In the case of $d=1$, i.e., perimeter, the maximum side length of each box
-$\mathfrak{m}_{1}^{(N)}\ =\ \textup{max}_{1\leq i\leq m}P_{i}^{(1)}\cdots P_{i}^{(N)}$ (1.8)
-converges to strong Benford behavior base $B$ as $N\rightarrow\infty$.
-Hence, it was conjectured in the same paper (see Conjecture 4.1) that for any linear fragmentation process, $m_{d}^{(N)}$ converges to strong Benford behavior.
-Conjecture 1.12.
-[BDMMM] Every linear fragmentation process satisfies the Maximum Criterion in all dimensions $1\leq d\leq m$.
-We prove the conjecture for any $m\geq 2$ and $d\leq m$ under some mild conditions on the linear fragmentation process (see Section 3).
-Theorem 1.13.
-Suppose that $\mathfrak{B}_{0}\supset\mathfrak{B}_{1}\supset\mathfrak{B}_{2}\supset\cdots$ is a linear fragmentation process on an $m$-dimensional box for an arbitrary $m\geq 2$, such that the proportion cut $P_{i}$ at every stage satisfies the following.
-  1. (1)
-    We have $\mathbb{E}[\log_{B}(P_{i})]=0$ and $\textup{Var}[\log_{B}(P_{i})]=1$.
-  2. (2)
-    We have $\log(P_{i})$ is supported on [C,C][-C,C], where $C>0$ is a constant.
-  3. (3)
-    Suppose that $P^{(k)}_{i}$ is a sequence of i.i.d. random variables $\sim P_{i}$. Define the random variable
-    $Z_{i}^{(N)}\ :=\ \frac{\log_{B}(P_{i}^{(1)}\cdots P_{i}^{(N)})}{\sqrt{N}}.$ (1.9)
-    Then $Z_{i}^{(N)}$ has probability density function $f_{Z_{i}^{(N)}}(z_{i})=\varphi(z_{i})+A(z_{i})$ and cumulative density function $F_{Z_{i}^{(N)}}(z_{i})=\Phi(z_{i})+B(z_{i})$, where $\varphi(x)$ and $\Phi(x)$ are the probability density function and cumulative density function of the standard normal $N(0,1)$, $A(x)=O(N^{-1/2-\delta})$ and $B(x)=O(N^{-\delta})$ for some $\delta>0$.
-Then this linear fragmentation process satisfies the condition of the Maximum Criterion for any $d\leq m$, i.e., $\mathfrak{m}_{d}^{(N)}$ converges to strong Benford behavior base $B$ as $N\rightarrow\infty$ for any $d\leq m$, thus by Theorem 1.10, $V_{d}^{(N)}$ converges to strong Benford behavior base $B$ as $N\rightarrow\infty$ for any $d\leq m$.
-2. Proof of Theorem 1.6
-2.1. Preliminary
-In [FILM], Fang, Irons, Lippelman and Miller proved that if $y_{i}\in\mathbb{Q}$ for all $1\leq i\leq m-1$, then the distribution of the stick lengths does not follow Benford behavior as $N\rightarrow\infty$.
-To be exact, they were able to show that the distribution of the logarithms of the stick lengths converges to a discrete distribution, contrary to a uniform distribution on $[0,1]$ which is an equivalent condition for convergence to strong Benford behavior by Uniform Distribution Characterization 1.4.
-Hence, having $y_{i}\not\in\mathbb{Q}$ for some $1\leq i\leq m-1$ is a necessary condition for the distribution of the stick lengths to follow Benford behavior.
-For the remainder of this section, we show that having $y_{i}\not\in\mathbb{Q}$ for some $1\leq i\leq m-1$ is also a sufficient condition for the distribution of the stick lengths to converge to Benford’s behavior.
-Since we can always reorder the $p_{i}^{k_{i}}$’s in the stick length $A_{k_{1},k_{2},\dots,k_{m}}:=p_{1}^{k_{1}}p_{2}^{k_{2}}\cdots p_{m}^{k_{m}}$ and thus the order of $y_{i}$’s in the factorization, then it suffices to show that the first exponent $y_{1}\not\in\mathbb{Q}$ is a sufficient condition for the distribution of the stick lengths to converge to Benford’s behavior.
-We know that after stage $N$, the number of sticks of length $A_{k_{1},k_{2},\dots,k_{m}}$ is equal to the multinomial coefficient
-$\binom{k_{1}+k_{2}+\cdots+k_{m}}{k_{1},k_{2},\dots,k_{m}}\ =\ \binom{N}{k_{1},k_{2},\dots,k_{m}}$. (2.1)
-From the proof of Theorem , we know that the stick length $A_{k_{1},k_{2},\dots,k_{m}}$ has the factorization
-$A_{k_{1},k_{2},\dots,k_{m}}\ =\ p_{1}^{k_{1}}p_{2}^{k_{2}}\cdots p_{m}^{k_{m}}\ =\ \left(\frac{p_{1}}{p_{2}}\right)^{k_{1}}\left(\frac{p_{2}}{p_{3}}\right)^{k_{1}+k_{2}}\cdots\left(\frac{p_{m-1}}{p_{m}}\right)^{\sum_{j=1}^{m-1}k_{j}}(p_{m})^{N},$ (2.2)
-which motivates us to define the exponents $y_{i}=\log(p_{i}/p_{i+1})$ for $1\leq i\leq m-1$. Now, fix an interval $(a,b)\subset(0,1)$. Let
-$\chi_{(a,b)}(x)\ :=\ \mathbbm{1}\Big{(}\log(x)\textup{ mod 1}\in(a,b)\Big{)}$ (2.3)
-be the indicator function for $\log(x)\textup{ mod 1}\in(a,b)$. Then after stage $N$, the probability that a stick length mod 1 is in $(a,b)$ equals
-$F_{N}(a,b)\ :=\ \frac{1}{m^{N}}\sum_{\begin{subarray}{c}0\leq k_{1},k_{2}\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}\chi_{(a,b)}\left(\prod_{i=1}^{m}p_{i}^{k_{i}}\right).$ (2.4)
+If $\mathfrak{m}_{d}^{(N)}$ converges to strong Benford behavior in base $B$ as $N\rightarrow\infty$, then $V_{d}^{(N)}$ also converges to strong Benford behavior in base $B$ as $N\rightarrow\infty$.
+In the same paper, the Maximum Criterion was verified for $d=1$ in a special case.
+Theorem 4.
+Let $P_{i}^{(j)}$ be i.i.d. with $\log_{B}P_{i}^{(j)}$’s having uniform distributions. When $d=1$, the maximum side length of each box
+$\mathfrak{m}_{1}^{(N)}\ =\ \textup{max}_{1\leq i\leq m}P_{i}^{(1)}\cdots P_{i}^{(N)}$ (10)
+converges to strong Benford behavior in base $B$ as $N\rightarrow\infty$.
+It was conjectured in the same paper (see Conjecture 4.1) that for any linear fragmentation process, $m_{d}^{(N)}$ converges to strong Benford behavior.
+Conjecture 1.
+Every linear fragmentation process satisfies the Maximum Criterion in all dimensions $1\leq d\leq m$.
+We prove the conjecture for any $m\geq 2$ and $d\leq m$ under some conditions on the linear fragmentation process, which we shall explain more in Section 3.
+Theorem 5.
+Suppose that $\mathfrak{B}_{0}\supset\mathfrak{B}_{1}\supset\mathfrak{B}_{2}\supset\cdots$ is a linear fragmentation process on an $m$-dimensional box for an arbitrary $m\geq 2$, such that the proportion cut $P_{i}^{(j)}$ at every stage satisfies the following.
+  1. 1.
+    Condition 1: $P_{i}^{(j)}$’s are i.i.d.
+for all $1\leq i\leq m$ and $1\leq j\leq N$.
+That is, $P_{i}^{(j)}$’s are all independent and there exists a random variable $P$ such that $P_{i}^{(j)}\sim P$ for all $i,j$;
+  2. 2.
+    Condition 2: Let the mean and the variance of $\log_{B}(P)$ be $\mu_{P}$ and $\sigma_{P}$.
+Then $\log_{B}(P)$ has finite centered moments, i.e., $\mathbb{E}[(\log_{B}(P)-\mu_{P})^{k}]<\infty$ for all $k\geq 1$ and the characteristic function of $(P-\mu_{P})/\sigma_{P}$ admits Taylor expansion.
+Moreover, the variance is nonzero;
+  3. 3.
+    Condition 3: Let $f(x)$ be the probability density function of $\log_{B}(P)$, then we have $\textup{ess sup}_{x}f(x)<\infty$;
+  4. 4.
+    Condition 4: $f(x)$ is differentiable with $\textup{sup}_{x}|f^{\prime}(x)|<\infty$.
+Then this linear fragmentation process satisfies the condition of the Maximum Criterion for any $d\leq m$, i.e., $\mathfrak{m}_{d}^{(N)}$ converges to strong Benford behavior in base $B$ as $N\rightarrow\infty$ for any $d\leq m$, thus by Theorem 3, $V_{d}^{(N)}$ converges to strong Benford behavior in base $B$ as $N\rightarrow\infty$ for any $d\leq m$.
+2 Proof of Theorem 2
+2.1 Preliminary
+In [13], Fang, Irons, Lippelman and Miller proved that if $y_{i}\in\mathbb{Q}$ for all $1\leq i\leq m-1$, the distribution of the logarithms mod 1 of the stick lengths converges to a discrete distribution instead of a uniform distribution on $[0,1]$.
+Hence, having $y_{i}\not\in\mathbb{Q}$ for some $1\leq i\leq m-1$ is a necessary condition for the distribution of the stick lengths to follow strong Benford behavior.
+For the remainder of this section, we show that having $y_{i}\not\in\mathbb{Q}$ for some $1\leq i\leq m-1$ is also a sufficient condition for the distribution of the stick lengths to converge to strong Benford’s behavior.
+After reordering the $p_{i}^{k_{i}}$’s and thus the $y_{i}$’s in the factorization of the stick length $A^{(N)}_{k_{1},k_{2},\dots,k_{m}}:=p_{1}^{k_{1}}p_{2}^{k_{2}}\cdots p_{m}^{k_{m}}$, it suffices to show that having the first exponent $y_{1}\not\in\mathbb{Q}$ is a sufficient condition for the distribution of the stick lengths to converge to strong Benford’s behavior.
+We know that after stage $N$, there are
+$\binom{k_{1}+k_{2}+\cdots+k_{m}}{k_{1},k_{2},\dots,k_{m}}\ =\ \binom{N}{k_{1},k_{2},\dots,k_{m}}$. (11)
+number of sticks of length $A^{(N)}_{k_{1},k_{2},\dots,k_{m}}$. From the proof of Theorem 1 of [13], we know that the stick length $A^{(N)}_{k_{1},k_{2},\dots,k_{m}}$ has the factorization
+$A^{(N)}_{k_{1},k_{2},\dots,k_{m}}\ =\ p_{1}^{k_{1}}p_{2}^{k_{2}}\cdots p_{m}^{k_{m}}\ =\ \left(\frac{p_{1}}{p_{2}}\right)^{k_{1}}\left(\frac{p_{2}}{p_{3}}\right)^{k_{1}+k_{2}}\cdots\left(\frac{p_{m-1}}{p_{m}}\right)^{\sum_{j=1}^{m-1}k_{j}}(p_{m})^{N},$ (12)
+which motivates us to define the exponents $y_{i}=\log_{B}(p_{i}/p_{i+1})$ for $1\leq i\leq m-1$. Now, fix an interval $(a,b)\subset(0,1)$. Let
+$\chi_{(a,b)}(x)\ :=\ \mathbbm{1}\Big(\log_{B}(x)\textup{ mod 1}\in(a,b)\Big)$ (13)
+be the indicator function for $\log_{B}(x)\textup{ mod 1}\in(a,b)$. Then after stage $N$, the probability that the logarithm mod 1 of a stick length is in $(a,b)$ equals
+$F_{N}(a,b)\ :=\ \frac{1}{m^{N}}\sum_{\begin{subarray}{c}0\leq k_{1},k_{2}\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}\chi_{(a,b)}\left(\prod_{i=1}^{m}p_{i}^{k_{i}}\right).$ (14)
 Our goal is to prove that as $N\rightarrow\infty$,
-$F(a,b)\ :=\ \lim_{N\rightarrow\infty}F_{N}(a,b)\ =\ b-a,$ (2.5)
-which is a sufficient condition for Benford behavior.
-2.2. Combinatorial identities
-The key steps to our proof are two combinatorial identities related to multinomial coefficients, which we present and prove below. We first review the multinomial coefficients.
-Definition 2.1.
+$F(a,b)\ :=\ \lim_{N\rightarrow\infty}F_{N}(a,b)\ =\ b-a,$ (15)
+which by the Uniform Distribution Characterization implies strong Benford behavior.
+2.2 Combinatorial identities
+The key steps to our proof are two combinatorial identities related to multinomial coefficients, which we present in this subsection. We first review the definition of multinomial coefficients.
+Definition 9.
 Suppose that $n,k_{1},k_{2},\dots,k_{m}\geq 0$ and $k_{1}+\cdots+k_{m}=n$, then we define the multinomial coefficients by
-$\binom{n}{k_{1},k_{2},\dots,k_{m}}\ :=\ \frac{n!}{k_{1}!k_{2}!\cdots k_{m}!}.$ (2.6)
-The following identity on the factorization of multinomial coefficients allows us to express multinomial coefficients in terms of product of binomial coefficients.
-As we shall see, this forms the key ingredient that reduces our multi-proportion problem into the original single-proportion problem in [B–].
-Lemma 2.2.
-[Stan] For $m\geq 1$,
-$\binom{k_{1}+k_{2}+\cdots+k_{m}}{k_{1},k_{2},\dots,k_{m}}\ =\ \prod_{i=1}^{m}\binom{k_{1}+\cdots+k_{i}}{k_{i}}$. (2.7)
-The proof for the identity above is a straightforward induction through the expansion of the multinomial and binomial coefficients.
-Below is another identity that is especially useful later for evaluating sums of multinomial coefficients.
-Lemma 2.3.
-[Stan] For $N\in\mathbb{Z}_{\geq 0}$ and $m\in\mathbb{Z}_{>0}$,
+$\binom{n}{k_{1},k_{2},\dots,k_{m}}\ :=\ \frac{n!}{k_{1}!k_{2}!\cdots k_{m}!}.$ (16)
+The following identity allows us to express multinomial coefficients as the product of binomial coefficients.
+As we shall see, this forms a key ingredient to our proof, as it helps reduce our multi-proportion problem to the original single-proportion problem in [1].
+Lemma 1.
+[20] For $m\geq 1$,
+$\binom{k_{1}+k_{2}+\cdots+k_{m}}{k_{1},k_{2},\dots,k_{m}}\ =\ \prod_{i=1}^{m}\binom{k_{1}+\cdots+k_{i}}{k_{i}}$. (17)
+The proof for Lemma 1 follows a straightforward inductive argument based on the expansion of the multinomial and binomial coefficients.
+Below is another identity that comes in handy when evaluating sums of multinomial coefficients.
+Lemma 2.
+[20] For $N\in\mathbb{Z}_{\geq 0}$ and $m\in\mathbb{Z}_{>0}$,
 $m^{N}$ $\ =\ \sum_{\begin{subarray}{c}0\leq k_{1},k_{2},\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}\ =\ \sum_{\begin{subarray}{c}0\leq k_{1},\dots,k_{m-1}\leq N\\ k_{1}+k_{2}+\cdots+k_{m-1}=N\end{subarray}}2^{k_{1}}\binom{N}{k_{1},k_{2},\cdots,k_{m-1}}$.
-(2.8)
-The most straightforward proof of the identity above relies on two different ways of representing and expanding $m^{N}$: one as $(\underbrace{1+\cdots+1}_{\textrm{$N$}})^{N}$ and the other as $(2+\underbrace{1+\cdots+1}_{\textrm{$N-2$}})^{N}$.
-See [Stan] for more.
-2.3. Truncation
-We prove that for any $\epsilon\in(0,1)$, the contribution to $F_{N}(a,b)$ of multinomial coefficients with $k_{1}+k_{2}<N^{\epsilon}$ is negligible.
-This allows us to only consider $k_{1}+k_{2}$ that are sufficiently large for these stick lengths $A_{k_{1},k_{2},\dots,k_{m}}$.
-Proposition 2.4.
-For any $\epsilon\in(0,1)$,
-$\sum_{\begin{subarray}{c}0\leq k_{1},k_{2}\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}\ \leq\ (m-2)^{N}N^{N^{\epsilon}+2}.$ (2.9)
+(18)
+The proof for Lemma 2 relies on two different representation and expansion of $m^{N}$: one as $(\underbrace{1+\cdots+1}_{\textrm{$N$}})^{N}$ and the other as $(2+\underbrace{1+\cdots+1}_{\textrm{$N-2$}})^{N}$.
+See [20] for more.
+2.3 Truncation
+First, we would like to make some truncation to the sum in (14).
+We show that for any $\epsilon\in(0,1)$, the contribution to $F_{N}(a,b)$ of multinomial coefficients with $k_{1}+k_{2}<N^{\epsilon}$ is negligible.
+This allows us to only consider stick lengths $A_{k_{1},k_{2},\dots,k_{m}}$ where $k_{1}+k_{2}$ are sufficiently large.
+Proposition 2.
+For any $0<\epsilon<1$,
+$\sum_{\begin{subarray}{c}0\leq k_{1},k_{2}\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}\ \leq\ (m-2)^{N}N^{N^{\epsilon}+2}.$ (19)
 Proof.
@@
 $\binom{N}{k_{1},k_{2},\dots,k_{m}}$ $\ =\ \frac{N(N-1)\cdots(N-(k_{1}+k_{2})+1)}{k_{1}!k_{2}!}\cdot\frac{(N-(k_{1}+k_{2}))!}{k_{3}!k_{4}!\cdots k_{m}!}$
-$\ \leq\ N^{k_{1}+k_{2}}\cdot\binom{N-(k_{1}+k_{2})}{k_{3},k_{4},\dots,k_{m}}$. (2.10)
+$\ \leq\ N^{k_{1}+k_{2}}\cdot\binom{N-(k_{1}+k_{2})}{k_{3},k_{4},\dots,k_{m}}$. (20)
 Hence,
-$\sum_{\begin{subarray}{c}0\leq k_{1},k_{2},\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}$ $\ =\ \sum_{\begin{subarray}{c}0\leq k_{1},k_{2}<N^{\epsilon}\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}N^{k_{1}+k_{2}}\sum_{\begin{subarray}{c}0\leq k_{3},k_{4},\dots,k_{m}\leq N-(k_{1}+k_{2})\\ k_{3}+k_{4}+\cdots+k_{m}=N-(k_{1}+k_{2})\end{subarray}}\binom{N-(k_{1}+k_{2})}{k_{3},k_{4},\dots,k_{m}}$
+$\sum_{\begin{subarray}{c}0\leq k_{1},k_{2},\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}$
+$\ =\ \sum_{\begin{subarray}{c}0\leq k_{1},k_{2}<N^{\epsilon}\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}N^{k_{1}+k_{2}}\sum_{\begin{subarray}{c}0\leq k_{3},k_{4},\dots,k_{m}\leq N-(k_{1}+k_{2})\\ k_{3}+k_{4}+\cdots+k_{m}=N-(k_{1}+k_{2})\end{subarray}}\binom{N-(k_{1}+k_{2})}{k_{3},k_{4},\dots,k_{m}}$
 $\ \leq\ (m-2)^{N}\sum_{\begin{subarray}{c}0\leq k_{1},k_{2}<N^{\epsilon}\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}N^{k_{1}+k_{2}}$
-$\ \leq\ (m-2)^{N}(N^{\epsilon})^{2}N^{N^{\epsilon}}\ \leq\ (m-2)^{N}N^{N^{\epsilon}+2}.$ (2.11)
+$\ \leq\ (m-2)^{N}(N^{\epsilon})^{2}N^{N^{\epsilon}}\ \leq\ (m-2)^{N}N^{N^{\epsilon}+2}.$ (21)
 We know the probability that a stick length $p_{1}^{k_{1}}p_{2}^{k_{2}}\cdots p_{m}^{k_{m}}$ satisfies $k_{1}+k_{2}<N^{\epsilon}$ is
-$\frac{1}{m^{N}}\sum_{\begin{subarray}{c}0\leq k_{1},k_{2}\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}$. (2.12)
-According to Proposition 2.4, we have that (2.12) is bounded above by $((m-2)/m)^{N}N^{N^{\epsilon}+2}$. Hence, the logarithm of the probability is bounded above by
-$\log\left(\left(\frac{m-2}{m}\right)^{N}N^{N^{\epsilon}+2}\right)\ =\ N\log\left(\frac{m-2}{m}\right)+(N^{\epsilon}+2)\log(N).$ (2.13)
-It is clear that (2.13) goes to -$\infty$ as $N\rightarrow\infty$.
+$\frac{1}{m^{N}}\sum_{\begin{subarray}{c}0\leq k_{1},k_{2}\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}$. (22)
+According to Proposition 2, we have that (22) is bounded above by $((m-2)/m)^{N}N^{N^{\epsilon}+2}$. Hence, the logarithm of the probability is bounded above by
+$\log\left(\left(\frac{m-2}{m}\right)^{N}N^{N^{\epsilon}+2}\right)\ =\ N\log\left(\frac{m-2}{m}\right)+(N^{\epsilon}+2)\log(N).$ (23)
+It is clear that the above goes to -$\infty$ as $N\rightarrow\infty$.
 Thus the probability goes to 0 as $N\rightarrow\infty$.
@@
 Now, fix $k(N)\geq N^{\epsilon}$ and let $0\leq k_{1},k_{2}\leq k(N)$ with $k_{1}+k_{2}=k(N)$ and $0\leq k_{3},k_{4},\dots,k_{m}\leq N-k(N)$ with $k_{3}+k_{4}+\dots+k_{m}=N-k(N)$.
-By [B–], we know that the frequency $k_{1}$ of stick length $A_{k_{1},k_{2},\dots,k_{m}}=p_{1}^{k_{1}}p_{2}^{k_{2}}\dots p_{m}^{k_{m}}$ follows a binomial distribution with mean $k(N)/2$ and standard deviation $\sqrt{k(N)}/2$.
+By [1], we know that the frequency $k_{1}$ follows a binomial distribution with mean $k(N)/2$ and standard deviation $\sqrt{k(N)}/2$.
 Pick some $\delta\in(0,\epsilon/10)$.
 We see that it suffices to consider cases where $|k_{1}-k(N)/2|<(\lceil N^{\delta}\rceil\sqrt{k(N)})/2$, because the probability that $k_{1}$ is outside this range is asymptotically small by Chebyshev’s inequality:
-$\mathbb{P}\left(\left|k_{1}-\frac{k(N)}{2}\right|\ \geq\ \frac{\lceil N^{\delta}\rceil\sqrt{k(N)}}{2}\right)\ \leq\ \frac{1}{\lceil N^{\delta}\rceil^{2}}.$ (2.14)
-2.4. Near uniform probability within small intervals
+$\mathbb{P}\left(\left|k_{1}-\frac{k(N)}{2}\right|\ \geq\ \frac{\lceil N^{\delta}\rceil\sqrt{k(N)}}{2}\right)\ \leq\ \frac{1}{\lceil N^{\delta}\rceil^{2}}.$ (24)
+2.4 Near uniform probability within small intervals
 […]
-$k_{1,\ell}\ :=\ f_{\ell}\left(\frac{k(N)}{2}\right)+\ell\lceil N^{\delta}\rceil,\quad k_{1,\ell,i}\ :=\ f_{\ell}\left(\frac{k(N)}{2}\right)+\ell\lceil N^{\delta}\rceil+i,\quad 0\leq i\leq N^{\delta}-1,$ (2.15)
+$k_{1,\ell}\ :=\ f_{\ell}\left(\frac{k(N)}{2}\right)+\ell\lceil N^{\delta}\rceil,$
+$k_{1,\ell,i}\ :=\ f_{\ell}\left(\frac{k(N)}{2}\right)+\ell\lceil N^{\delta}\rceil+i,\quad 0\leq i\leq N^{\delta}-1,$ (25)
 […]
 […]
-We follow Section (5.52) of [B–] to obtain a bound for the difference.
-Proposition 2.5.
+We follow Section (5.52) of [1] to obtain a bound for the difference.
+Proposition 3.
 For $\ell\leq\sqrt{k(N)}/2$,
-$\left|\binom{k(N)}{k_{1,\ell}}-\binom{k(N)}{k_{1,\ell+1}}\right|\ \leq\ O\left(\binom{k(N)}{k_{1,\ell}}\cdot N^{-\frac{3\epsilon}{10}}\right).$ (2.17)
-Since the proof follows similar argument to Section (5.52) of [B–], we leave the proof details to Appendix A.
-2.5. Equidistribution within small intervals
-For fixed $k(N):=k_{1}+k_{2},k_{3},k_{4}\dots,k_{m}$, and $\ell$, the stick length
-$A_{k_{1,\ell,i},k_{2},\dots,k_{m}}\ =\ \left(\frac{p_{1}}{p_{2}}\right)^{k_{1,\ell,i}}\left(\frac{p_{2}}{p_{3}}\right)^{k(N)}\left(\frac{p_{3}}{p_{4}}\right)^{k(N)+k_{3}}\cdots\left(\frac{p_{m-1}}{p_{m}}\right)^{k(N)+\sum_{j=3}^{m-1}k_{j}}(p_{m})^{N}$ (2.18)
+$\left|\binom{k(N)}{k_{1,\ell}}-\binom{k(N)}{k_{1,\ell+1}}\right|\ \leq\ O\left(\binom{k(N)}{k_{1,\ell}}\cdot N^{-\frac{3\epsilon}{10}}\right).$ (27)
+Since the proof follows a similar argument to Section (5.52) of [1], we leave the proof details to A.
+2.5 Equidistribution within small intervals
+In this subsection, we want to show that for fixed $k(N),k_{3},\dots,k_{m}$, the logarithm of the stick length $\log_{B}(A^{(N)}_{k_{1,\ell,i},k_{2},\dots,k_{m}})$ for $0\leq i\leq\lceil N^{\delta}\rceil-1$ converges to being equidistributed mod 1, in the sense that the logarithm of the stick length is thought of as a random variable that takes value in $\log_{B}(A^{(N)}_{k_{1,\ell,i},k_{2},\dots,k_{m}})$ for $0\leq i\leq\lceil N^{\delta}\rceil-1$ with equal probabilities.
+We first state the following theorem which provides an easy criterion for checking equidistribution mod 1.
+Corollary 1.
+[15] For any $a\in\mathbb{R}$, the arithmetic progression $\{a+nd\}_{n=1}^{\infty}$ is equidistributed mod 1 if and only if $d\not\in\mathbb{Q}$.
+Now, note that for fixed $k(N)$ and $\ell$,
+$\log_{B}(A^{(N)}_{k_{1,\ell,i},k_{2},\dots,k_{m}})$
+$\ =\ k_{1,\ell,i}\log_{B}\left(\frac{p_{1}}{p_{2}}\right)+(k_{1}+k_{2})\log_{B}\left(\frac{p_{2}}{p_{3}}\right)+\cdots+\left(\sum_{j=1}^{m-1}k_{j}\right)\log_{B}\left(\frac{p_{m-1}}{p_{m}}\right)+N\log(p_{m})$
+$\ =\ k_{1,\ell,i}\log_{B}\left(\frac{p_{1}}{p_{2}}\right)+k(N)\log_{B}\left(\frac{p_{2}}{p_{3}}\right)+\cdots+\left(\sum_{j=1}^{m-1}k_{j}\right)\log_{B}\left(\frac{p_{m-1}}{p_{m}}\right)+N\log(p_{m})$ (28)
 […]
-Definition 2.6.
-[KN] Suppose that $x$ is a real number.
+Definition 10.
+[15] Suppose that $x$ is a real number.
 The irrationality exponent $\mu$ of $x$ is the supremum of the set of $\mu$ such that $0<|x-p/q|<1/q^{\mu}$ is satisfied by an infinite number of coprime integer pairs $(p,q)$ with $q>0$.
-Theorem 2.7.
-[KN, Theorem 3.2, Example 3.2] Let $\kappa$ be the irrationality exponent of $\log_{B}(p_{1}/p_{2})$. If $\kappa<\infty$, then
-$|J_{\ell}(a,b)|\ =\ (b-a)N^{\delta}+O\left(N^{\delta\left(1-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\right),$ (2.19)
-for some $\epsilon^{\prime}>0$ such that $\frac{1}{\kappa}+\epsilon^{\prime}<0$ and there is a power saving with the error term. If $\kappa=\infty$, then
-$|J_{\ell}(a,b)|\ =\ (b-a)N^{\delta}+o(N^{\delta}).$ (2.20)
-In either case, the error term is asymptotically much smaller than the main term $(b-a)N^{\delta}$.
-Thus, if $\log_{B}(p_{1}/p_{2})\not\in\mathbb{Q}$, then $\log_{B}(A_{k_{1,\ell,i},k_{2},\dots,k_{m}})$ converges to being equidistributed mod 1.
-2.6. Evaluation of sums within and across intervals
+If such a set does not exist, then we say $x$ has irrationality exponent $\infty$.
+Since $\mathbb{Q}$ is dense in $\mathbb{R}$, then every real number can be approximated by rational numbers.
+However, what differentiates them is how well they can be approximated.
+Irrationality exponent measures exactly how well a real number can be approximated by rational numbers.
+The bigger the irrationality exponent, the finer the rational approximation can be.
+The following are some well-known facts about irrationality exponent, which comes handy later in this section.
+Proposition 4.
+[15] The irrationality exponent of any rational number is 1. As a consequence of Dirichlet’s approximation theorem, which states that for any irrational number $x$,
+$\left|x-\frac{p}{q}\right|\ <\ \frac{1}{q^{2}}$ (29)
+for infinitely $(p,q)$ where $p$ and $q$ are coprime, the irrationality exponent of any irrational number is at least 2.
+Further more, any algebraic irrational (irrational numbers that are zeros of polynomials over $\mathbb{Q}$) has irrationality exponent exactly 2.
+We now state a theorem that quantifies the discrepancy of the logarithm of the stick length in terms of irrationality exponent.
+Theorem 6.
+[15, Theorem 3.2] Let $\kappa$ be the irrationality exponent of $\log_{B}(p_{1}/p_{2})$ and $J_{\ell}(a,b)\subset\{0,1,\dots,N^{\delta}-1\}$ be the set of indices $i$ such that $\log_{B}(A^{(N)}_{k_{1,\ell,i},k_{2},\dots,k_{m}})$ mod 1 is in $(a,b)$.
+If $\kappa<\infty$, then
+$|J_{\ell}(a,b)|\ =\ (b-a)N^{\delta}+O\left(N^{\delta\left(1-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}\right),$ (30)
+for every $\epsilon^{\prime}>0$ and there is a power saving with the error term. The error term is optimal in the sense that
+$\left|\frac{|J_{\ell}(a,b)|}{N^{\delta}}-(b-a)\right|\ =\ \Omega\left(N^{\delta(-\frac{1}{\kappa-1})-\epsilon^{\prime}}\right)$ (31)
+for every $\epsilon^{\prime}>0$. If $\kappa=\infty$, then
+$|J_{\ell}(a,b)|\ =\ (b-a)N^{\delta}+o(N^{\delta}).$ (32)
+Note that [15] has a different convention for irrationality exponent and if the irrationality exponent of $x$ is $\mu$ in our definition, then it is $\mu-1$ in their definition.
+This is the reason for the change made to the presentation of Theorem [15], where there is a $1/(\kappa-1)$ instead of a $1/\kappa$ in the exponent.
+2.6 Evaluation of sums within and across intervals
 […]
-$\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$. (2.21)
-By Proposition 2.5 and Section 2.5, we have
+$\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$. (33)
+By Proposition 3 and Section 2.5, we have
 $\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$ $\ =\ \sum_{i\in J_{\ell}(a,b)}\Bigg\{\binom{k(N)}{k_{1,\ell}}+O\left(\binom{k(N)}{k_{1,\ell}}N^{-\frac{3\epsilon}{10}}\right)\Bigg\}$
 $\ =\ \Bigg\{\binom{k(N)}{k_{1,\ell}}\sum_{i\in J_{\ell}(a,b)}1\Bigg\}+O\left(\binom{k(N)}{k_{1,\ell}}N^{-\frac{3\epsilon}{10}}\sum_{i\in J_{\ell}(a,b)}1\right)$
-$\ =\ (b-a)N^{\delta}\binom{k(N)}{k_{1,\ell}}+O\left(\binom{k(N)}{k_{1,\ell}}N^{\delta\left(1-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\right).$ (2.22)
+$\ =\ (b-a)N^{\delta}\binom{k(N)}{k_{1,\ell}}+O\left(\binom{k(N)}{k_{1,\ell}}N^{\delta\left(1-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}\right).$ (34)
 We justify the error term in the last line of (2.6).
-Since $\kappa$ is the irrationality exponent of an irrational number, then $\kappa>1$ (see [BBS]).
-Also, $\epsilon^{\prime}>0$, so $1<-1/\kappa+\epsilon^{\prime}$.
-Moreover, $1/\kappa+\epsilon^{\prime}<0$ by assumption.
+Since $\kappa$ is the irrationality exponent of an irrational number, then $\kappa\geq 2$ (see [2]).
+Also, $\epsilon^{\prime}>0$, so $1<-1/(\kappa-1)+\epsilon^{\prime}$.
+Moreover, for some proper choice of $\epsilon^{\prime}$, we have $1/(\kappa-1)+\epsilon^{\prime}<0$.
 Combining these with the fact that $\delta\in(0,\epsilon/10)$ gives the dominant error term in the last line of (2.6).
-Now, we count the number of stick lengths $A_{k_{1,\ell,i},k_{2},\dots,k_{m}}$ such that $\log_{B}(A_{k_{1,\ell,i},k_{2},\dots,k_{m}})$ mod 1 is in $(a,b)$ over all the intervals.
+Now, we count the number of stick lengths $A^{(N)}_{k_{1,\ell,i},k_{2},\dots,k_{m}}$ such that $\log_{B}(A^{(N)}_{k_{1,\ell,i},k_{2},\dots,k_{m}})$ mod 1 is in $(a,b)$ over all the intervals.
 By the truncation in Section 2.3 the main term comes from the sum over $\ell$ from $\sqrt{k(N)}/2$ to $\sqrt{k(N)}/2$.
@@
 $\sum_{\ell=-k(N)/(2N^{\delta})}^{k(N)/(2N^{\delta})}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}\ =\ \sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}+\sum_{|\ell|>\sqrt{k(N)}/2}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$.
-(2.23)
-Using (2.14), we can provide an upper bound on the second sum on the RHS of (2.23):
-$\sum_{|\ell|>\sqrt{k(N)}/2}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$ $\ \leq\ \mathbb{P}\left(\left|k_{1}-\frac{k(N)}{2}\right|\ \geq\ \frac{N^{\delta}\sqrt{k(N)}}{2}\right)\cdot\sum_{0\leq k_{1}\leq N}\binom{k(N)}{k_{1}}$
-$\ =\ \mathbb{P}\left(\left|k_{1}-\frac{k(N)}{2}\right|\ \geq\ \frac{N^{\delta}\sqrt{k(N)}}{2}\right)\cdot 2^{k(N)}\ =\ O\left(\frac{2^{k(N)}}{N^{2\delta}}\right).$ (2.24)
-On the other hand, we apply (2.6) to the first sum on the RHS of (2.23):
-$\sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}\ =\ \sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}(b-a)N^{\delta}\binom{k(N)}{k_{1,\ell}}+O\left(\sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}\binom{k(N)}{k_{1,\ell}}N^{\delta\left(1-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\right).$ (2.25)
-By (2.6) and (5.17) and (5.18) of [B–], by choosing $q=N^{\delta}$, we have
-$\sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$ $\ =\ (b-a)2^{k(N)}+O\left(2^{k(N)}N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\right).$ (2.26)
-Substituting (2.6) and (2.26) into (2.23) yields
-$\sum_{\ell=-k(N)/(2N^{\delta})}^{k(N)/(2N^{\delta})}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}\ =\ (b-a)2^{k(N)}+O\left(2^{k(N)}N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\right).$ (2.27)
-Finally, we sum over all $k(N),k_{3},k_{4},\dots,k_{m}$, which gives us the total number of lengths $A_{k_{1},k_{2},\dots,k_{m}}$ such that $\log_{B}(A_{k_{1},k_{2},\dots,k_{m}})$ mod 1 is in $(a,b)$.
+(35)
+Using (24), we can provide an upper bound on the second sum on the RHS of (35):
+$\sum_{|\ell|>\sqrt{k(N)}/2}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$ $\ \leq\ \mathbb{P}\left(\left|k_{1}-\frac{k(N)}{2}\right|\ \geq\ \frac{N^{\delta}\sqrt{k(N)}}{2}\right)\cdot\sum_{0\leq k_{1}\leq k(N)}\binom{k(N)}{k_{1}}$
+$\ =\ \mathbb{P}\left(\left|k_{1}-\frac{k(N)}{2}\right|\ \geq\ \frac{N^{\delta}\sqrt{k(N)}}{2}\right)\cdot 2^{k(N)}\ =\ O\left(\frac{2^{k(N)}}{N^{2\delta}}\right).$ (36)
+On the other hand, we apply (2.6) to the first sum on the RHS of (35):
+$\sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$ $\ =\ \sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}(b-a)N^{\delta}\binom{k(N)}{k_{1,\ell}}$
+$\quad\quad+O\left(\sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}\binom{k(N)}{k_{1,\ell}}N^{\delta\left(1-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}\right).$ (37)
+We know that
+$\sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}\binom{k(N)}{k_{1,\ell}}$ $\ =\ \sum_{\ell=-k(N)/(2N^{\delta})}^{k(N)/(2N^{\delta})}\binom{k(N)}{k_{1,\ell}}+O\left(\sum_{|\ell|>\sqrt{k(N)}/2}\binom{k(N)}{k_{1,\ell}}\right)$
+$\ =\ \sum_{\ell=-k(N)/(2N^{\delta})}^{k(N)/(2N^{\delta})}\binom{k(N)}{k_{1,\ell}}+O\left(\frac{2^{k(N)}}{N^{2\delta}}\right).$ (38)
+By (5.17) of [1], by choosing $q=N^{\delta}$, we have
+$\sum_{\ell=-k(N)/(2N^{\delta})}^{k(N)/(2N^{\delta})}\binom{k(N)}{k_{1,\ell}}\ =\ \frac{2^{k(N)}}{N^{\delta}}+O\left(2^{k(N)}e^{-3k(N)^{1-2\delta}}\right)$ (39)
+Substitute (39) into (2.6) and then substitute the resulting formula into (2.6), we have
+$\sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$ $\ =\ (b-a)2^{k(N)}+O\left(2^{k(N)}N^{-\delta}+2^{k(N)}N^{\delta\left(-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}\right)$
+$\ =\ (b-a)2^{k(N)}+O\left(2^{k(N)}N^{\delta\left(-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}\right),$ (40)
+where the last line follows from $1<-1/(\kappa-1)+\epsilon^{\prime}<0$. Substituting (2.6) and (2.6) into (35) yields
+$\sum_{\ell=-k(N)/(2N^{\delta})}^{k(N)/(2N^{\delta})}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}\ =\ (b-a)2^{k(N)}+O\left(2^{k(N)}N^{\delta\left(-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}\right).$ (41)
+Finally, we sum over all $k(N),k_{3},k_{4},\dots,k_{m}$, which gives us the total number of lengths $A^{(N)}_{k_{1},k_{2},\dots,k_{m}}$ such that $\log_{B}(A^{(N)}_{k_{1},k_{2},\dots,k_{m}})$ mod 1 is in $(a,b)$.
 We break the sum based on whether $k(N)\geq N^{\epsilon}$ or not.
-For $k(N)\geq N^{\epsilon}$, we apply (2.27):
+For $k(N)\geq N^{\epsilon}$, we apply (41):
 $\sum_{\begin{subarray}{c}0\leq k(N),k_{3},k_{4},\dots,k_{m}\leq N\\ k(N)+k_{3}+k_{4}+\cdots+k_{m}=N\\ k(N)\geq N^{\epsilon}\end{subarray}}\binom{N}{k(N),k_{3},k_{4},\dots,k_{m}}\sum_{\ell=-k(N)/(2N^{\delta})}^{k(N)/(2N^{\delta})}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$
 $\ =\ (b-a)\sum_{\begin{subarray}{c}0\leq k(N),k_{3},k_{4},\dots,k_{m}\leq N\\ k(N)+k_{3}+k_{4}+\cdots+k_{m}=N\\ k(N)\geq N^{\epsilon}\end{subarray}}2^{k(N)}\binom{N}{k(N),k_{3},k_{4},\dots,k_{m}}$
-$\quad+O\left(N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\sum_{\begin{subarray}{c}0\leq k(N),k_{3},k_{4},\dots,k_{m}\leq N\\ k(N)+k_{3}+k_{4}+\cdots+k_{m}=N\\ k(N)\geq N^{\epsilon}\end{subarray}}2^{k(N)}\binom{N}{k(N),k_{3},k_{4},\dots,k_{m}}\right)$.
-(2.28)
-Note that by Lemma 2.3
+$\quad\quad+O\left(N^{\delta\left(-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}\sum_{\begin{subarray}{c}0\leq k(N),k_{3},k_{4},\dots,k_{m}\leq N\\ k(N)+k_{3}+k_{4}+\cdots+k_{m}=N\\ k(N)\geq N^{\epsilon}\end{subarray}}2^{k(N)}\binom{N}{k(N),k_{3},k_{4},\dots,k_{m}}\right)$.
+(42)
+Note that by Lemma 2
 $\sum_{\begin{subarray}{c}0\leq k(N),k_{3},k_{4},\dots,k_{m}\leq N\\ k(N)+k_{3}+k_{4}+\cdots+k_{m}=N\\ k(N)\geq N^{\epsilon}\end{subarray}}2^{k(N)}\binom{N}{k(N),k_{3},k_{4},\dots,k_{m}}$
 $\ =\ m^{N}-\sum_{\begin{subarray}{c}0\leq k_{1},k_{2},\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}\sum_{k_{1}=0}^{k_{1}+k_{2}}\binom{N}{k_{1},k_{2},\dots,k_{m}}$
-$\ =\ m^{N}+O((m-2)^{N}N^{N^{\epsilon}+2}),$ (2.29)
-where in the last line, we apply Proposition 2.4. Substituting (2.6) into (2.6) gives
+$\ =\ m^{N}+O((m-2)^{N}N^{N^{\epsilon}+2}),$ (43)
+where in the last line, we apply Proposition 2. Substituting (2.6) into (2.6) gives
 $\sum_{\begin{subarray}{c}0\leq k(N),k_{3},k_{4},\dots,k_{m}\leq N\\ k(N)+k_{3}+k_{4}+\cdots+k_{m}=N\\ k(N)\geq N^{\epsilon}\end{subarray}}\binom{N}{k(N),k_{3},k_{4},\dots,k_{m}}\sum_{\ell=-k(N)/(2N^{\delta})}^{k(N)/(2N^{\delta})}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$
-$\ =\ (b-a)m^{N}+O\left((m-2)^{N}N^{N^{\epsilon}+2}+m^{N}N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}+(m-2)^{N}N^{N^{\epsilon}+2}N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\right)$
-$\ =\ (b-a)m^{N}+O\left(m^{N}N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\right).$ (2.30)
-For $k(N)<N^{\epsilon}$, by Lemma 2.7 and Proposition 2.4 we have
+$\ =\ (b-a)m^{N}+O\left((m-2)^{N}N^{N^{\epsilon}+2}+m^{N}N^{\delta\left(-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}+(m-2)^{N}N^{N^{\epsilon}+2}N^{\delta\left(-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}\right)$
+$\ =\ (b-a)m^{N}+O\left(m^{N}N^{\delta\left(-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}\right).$ (44)
+For $k(N)<N^{\epsilon}$, by Lemma 1 and Proposition 2 we have
 $\sum_{\begin{subarray}{c}0\leq k(N),k_{3},k_{4},\dots,k_{m}\leq N\\ k(N)+k_{3}+k_{4}+\cdots+k_{m}=N\\ k(N)<N^{\epsilon}\end{subarray}}\binom{N}{k(N),k_{3},k_{4},\dots,k_{m}}\sum_{\ell=-k(N)/(2N^{\delta})}^{k(N)/(2N^{\delta})}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$
 $\ \leq\ \sum_{\begin{subarray}{c}0\leq k(N),k_{3},k_{4},\dots,k_{m}\leq N\\ k(N)+k_{3}+k_{4}+\cdots+k_{m}=N\\ k(N)<N^{\epsilon}\end{subarray}}\binom{N}{k(N),k_{3},k_{4},\dots,k_{m}}\sum_{k_{1}=0}^{k(N)}\binom{k(N)}{k_{1}}$
-$\ =\ \sum_{\begin{subarray}{c}0\leq k_{1},k_{2},\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}\ =\ O\left((m-2)^{N}N^{N^{\epsilon}+2}\right).$ (2.31)
+$\ =\ \sum_{\begin{subarray}{c}0\leq k_{1},k_{2},\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}\ =\ O\left((m-2)^{N}N^{N^{\epsilon}+2}\right).$ (45)
 Combining (2.6) and (2.6), we have
 $\sum_{\begin{subarray}{c}0\leq k(N),k_{3},k_{4},\dots,k_{m}\leq N\\ k(N)+k_{3}+k_{4}+\cdots+k_{m}=N\end{subarray}}\binom{N}{k(N),k_{3},k_{4},\dots,k_{m}}\sum_{\ell=-k(N)/(2N^{\delta})}^{k(N)/(2N^{\delta})}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$
-$\ =\ (b-a)m^{N}+O\left(N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}m^{N}+N^{N^{\epsilon}+2}(m-2)^{N}\right)$
-$\ =\ (b-a)m^{N}+O\left(N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}m^{N}\right),$ (2.32)
+$\ =\ (b-a)m^{N}+O\left(N^{\delta\left(-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}m^{N}+N^{N^{\epsilon}+2}(m-2)^{N}\right)$
+$\ =\ (b-a)m^{N}+O\left(N^{\delta\left(-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}m^{N}\right),$ (46)
 After dividing by $m^{N}$, the total number of sticks after stage $N$, we arrive at
-$F_{N}(a,b)\ =\ b-a+O\left(N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\right).$ (2.33)
-We have shown that if $\log(p_{1}/p_{2})$ has a finite irrationality exponent $\kappa$, then the logarithm of the stick lengths is equidistributed mod 1, and thus the stick lengths converge to the Benford distribution.
-If $\log(p_{1}/p_{2})$ has infinite irrationality exponent, the calculation is exactly the same, except that we start with the error term $o(N^{\delta})$ instead of $O(N^{\delta\left(1-\frac{1}{\kappa}+\epsilon^{\prime}\right)})$ in Section 2.5, and so we end up with the $F_{N}(a,b)\ =\ b-a+o(1)$.
+$F_{N}(a,b)\ =\ b-a+O\left(N^{\delta\left(-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}\right).$ (47)
+Thus, we have shown that if $\log_{B}(p_{1}/p_{2})$ has a finite irrationality exponent $\kappa$, then the logarithm of the stick lengths is equidistributed mod 1, and thus the stick lengths converge to the Benford distribution.
+Moreover, the discrepancy between the distribution of the logarithm of the stick length and uniform distribution mod 1 is quantified in terms of the irrationality exponent of $\log_{B}(p_{1}/p_{2})$.
+Equation (47) tells us that the smaller the irrationality exponent, the smaller the discrepancy is (thanks to the guarantee that the error term is optimal), and the closer the distribution of the stick length is to the Benford distribution.
+If $\log(p_{1}/p_{2})$ has infinite irrationality exponent, the calculation is exactly the same, except that we start with the error term $o(N^{\delta})$ instead of $O(N^{\delta\left(1-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)})$ in Section 2.5, and so we end up with the $F_{N}(a,b)\ =\ b-a+o(1)$.
 Thus, this establishes the strong Benford behavior of fixed multi-proportion 1-dimensional stick fragmentation process when $y_{i}\not\in\mathbb{Q}$ for some $1\leq i\leq m-1$.
-∎
-3. Proof of Theorem 1.13
-3.1. Preliminary
-In this section, we prove Theorem 1.13, i.e., resolve Conjecture 4.1 of [BDMMM] that we mentioned in Section 1 under some mild conditions. Let us first recall the conjecture.
-Conjecture 3.1.
+On a final note, recall from Section 2.1 our choices of the order of $p_{1},p_{2},\dots,p_{m}$ and thus factorization are arbitrary.
+Hence, to obtain an optimal error term, we simply need to choose an order of $p_{1},p_{2},\dots,p_{m}$ such that $\log_{B}(p_{1}/p_{2})$ is irrational with the least irrationality exponent.
+Let $\kappa_{0}$ be such irrationality exponent.
+Then we have
+$F_{N}(a,b)\ =\ b-a+O\left(N^{\delta\left(-\frac{1}{\kappa_{0}-1}+\epsilon^{\prime}\right)}\right).$ (48)
+3 Proof of Theorem 5
+3.1 Preliminary
+In this section, we prove Theorem 5, i.e., Conjecture 4.1 of [6], under some conditions. Let us first recall the conjecture.
+Conjecture 1.
 Every linear fragmentation process satisfies the Maximum Criterion in all dimensions $1\leq d\leq m$.
-[BDMMM] has verified the conjecture in the case of $d=1$ and arbitrary $m$ for i.i.d.
-log-uniform distributions, i.e., the proportion at which each dimension is cut is $P_{i}$, where $\log P_{i}\sim\textup{Uniform}(a,b)$ and $a<b\leq 0$.
-They decide to normalize $P_{i}$ to $\textup{Uniform}(-\sqrt{3},\sqrt{3})$ with mean 0 and variance 1, which simplifies the calculation later on and does not affect the nature or the solution to the problem.
-After stage $N$, the side length of the $i$th dimension of the box, after taking logarithm and normalizing, is
-$Z_{i}^{(N)}\ :=\ \frac{\log_{B}(P_{i}^{(1)}\cdots P_{i}^{(N)})-N\mu_{P}}{\sqrt{N}\sigma_{P}}\ =\ \frac{\log_{B}(P_{i}^{(1)}\cdots P_{i}^{(N)})}{\sqrt{N}}.$ (3.1)
-By the CLT, $Z_{i}^{(N)}$ converges to $N(0,1)$.
-Moreover, Lemma 3.1 states that if $f_{Z_{i}^{(N)}}(z)$ is the probability density function of $Z_{i}^{(N)}$ and $\varphi(z)$ is the probability density function of $N(0,1)$, then for all $z\in[-\sqrt{3}N,\sqrt{3}N]$ and any $\epsilon>0$, we have
-$f_{Z_{i}^{(N)}}(z_{i})\ =\ \varphi(z_{i})+O(N^{-1+4\epsilon}),$ (3.2)
-where
-$\varphi(x)\ :=\ \frac{1}{\sqrt{2\pi}}e^{-x^{2}/2}.$ (3.3)
-Due to the presence of the error term, to establish the strong Benford behavior of the process, they need to rely on the fact that the distributions for the cuts are finitely supported with uniformly bounded supports.
-They also need to rely on the fact that the exponent of the error term is less than 1/2-1/2 for sufficiently small $\epsilon>0$, which allows the error term from integrating $f_{Z^{(N)}_{i}}(z)$ over the support to be asymptotically 0.
-For the rest of this paper, we prove Conjecture 4.1 of [BDMMM] for any $m\geq 2$ and $d\leq m$.
-Physically, this means that we consider any arbitrary dimensional volume of faces of that dimension of the $m$-dimensional box.
-We assume that the $\log(P_{i})$ are i.i.d., have mean 0 and variance 1, are finitely supported with support [C,C][-C,C], and that the probability density function (PDF) of $Z_{i}^{(N)}$ is $f_{Z_{i}^{(N)}}(z_{i})=\varphi(z_{i})+A(z_{i})$, where $A(z_{i})=O(N^{-1/2-\delta})$ for some $\delta>0$.
-We also assume that the cumulative density function (CDF) of $Z_{i}^{(N)}$ is given by $F_{Z_{i}^{(N)}}(z_{i})=\Phi(z_{i})+B(z_{i})$, where $\Phi$ is the CDF of $N(0,1)$ and $B(z_{i})=O(N^{-\delta})$.
-We prove that the maximum $d$-dimensional volume $\mathfrak{m}_{d}^{(N)}$ of a $d$-dimensional face of the $m$-dimensional box converges to strong Benford behavior.
-Once we prove this statement, then by the maximum criterion established in [BDMMM], we have that the $d$-dimensional volume $V_{d}^{(N)}$ of the $d$-dimensional faces of the $m$-dimensional box also converges to strong Benford behavior.
-The key idea behind our proof is the use of the Mean Value Theorem and a recursive method of change of variables and consistent cancellation that provide upper bounds on the error terms in terms of Gaussian densities.
-We shall explore this more in Section 3.6.
-To begin, we use knowledge from order statistics to find the PDF of $\mathfrak{m}_{d}^{(N)}$.
-Definition 3.2.
+Betti et al.
+[6] were able to prove the Conjecture for any arbitrary $m\geq 1$ and $d=1$.
+That is, they showed that under any linear fragmentation process, the length of the longest side of a fragmented box of any arbitrary dimension converges to strong Benford behavior.
+For the rest of this paper, we prove Theorem 5 for any $m\geq 1$ and $d\leq m$, i.e., we consider the volume of the largest $d$-dimensional face of a fragmented box of any arbitrary dimension $m\geq d$ converges to strong Benford behavior.
+Our proof relies on the following conditions on the distribution of $P_{i}^{(j)}$:
+  1. 1.
+    Condition 1: $P_{i}^{(j)}$’s are i.i.d.
+for all $1\leq i\leq m$ and $1\leq j\leq N$.
+That is, $P_{i}^{(j)}$’s are all independent and there exists a random variable $P$ such that $P_{i}^{(j)}\sim P$ for all $i,j$;
+  2. 2.
+    Condition 2: Let the mean and the variance of $\log_{B}(P)$ be $\mu_{P}$ and $\sigma_{P}$.
+Then $\log_{B}(P)$ has finite centered moments, i.e., $\mathbb{E}[(\log_{B}(P)-\mu_{P})^{k}]<\infty$ for all $k\geq 1$ and the characteristic function of $(P-\mu_{P})/\sigma_{P}$ admits Taylor expansion.
+Moreover, the variance is nonzero;
+  3. 3.
+    Condition 3: Let $f(x)$ be the probability density function of $\log_{B}(P)$, then we have $\textup{ess sup}_{x}f(x)<\infty$;
+  4. 4.
+    Condition 4: $f(x)$ is differentiable with $\textup{sup}_{x}|f^{\prime}(x)|<\infty$.
+These conditions ensure that we are working with “nice” proportion cuts and are satisfied by a wide range of random variables.
+Moreover, we believe that Theorem 5 should hold in a far more general setting, particular when the $P_{i}^{(j)}$’s have independent but non-identical distributions.
+For example, as long as $\{P_{i}^{(j)}\}_{j=1}^{N}$ satisfies the conditions for CLT for every $i$, then the distribution of the normalized side lengths of the box are still approximately Gaussian, and the proof techniques developed in this paper should extend without major difficulty.
+The only technical caveat is perhaps that we would need to use [21] instead of (53) to obtain a formula for the volume of the largest lower dimensional face of the box, since the side lengths are no longer i.i.d..
+The formula in [21] is rather complicated–it involves the permanent of a matrix whose entries are simple expressions in terms of the PDF and the CDF of random variables.
+However, after permanent expansion, it is essentially a linear sum with terms resembling (53).
+Hence, we believe that our analysis should extend to the general setting where $P_{i}^{(j)}$ are independent but non-identical without too much trouble, but for the clarity of exposition, we decide to restrict ourselves to the i.i.d.
+setting.
+Under these conditions, our goal is then to prove that the volume $\mathfrak{m}_{d}^{(N)}$ of the largest $d$-dimensional face of a $m$-dimensional fragmented box converges to strong Benford behavior.
+Once we prove this, then by the Maximum Criterion, the volume $V_{d}^{(N)}$ of the $d$-dimensional faces of a $m$-dimensional fragmented box also converges to strong Benford behavior.
+To begin, let us first introduce the notion of order statistics.
+Definition 11.
 Suppose that $X_{1},X_{2},\dots,X_{k}$ are random variables, and $_{i}$ returns the $i$th largest number among a list $(a_{1},a_{2},\dots,a_{k})$ of real numbers.
 For each outcome $\omega$ in the sample space $\Omega$, we define
-$X_{(k-i+1)}(w)\ :=\ [(X_{1}(w),X_{2}(w),\dots,X_{k}(w))]_{k-i+1}.$ (3.4)
+$X_{(k-i+1)}(w)\ :=\ [(X_{1}(w),X_{2}(w),\dots,X_{k}(w))]_{k-i+1}.$ (49)
 We say that the random variable $X_{(k-i+1)}$ is the $i$th order statistics, or the $i$th largest random variable among $X_{1},X_{2},\dots,X_{k}$.
 Hence, as random variables, $X_{(1)}\leq X_{(2)}\leq\dots\leq X_{(k)}$.
-Let $S_{i}^{(N)}$ be the side length of the $i\textsuperscript{th}$ dimension of the $m$-dimensional box after stage $N$.
-Then the maximum volume of a $d$-dimensional face is the product of the longest $d$ sides of the box, i.e., $\mathfrak{m}_{d}^{(N)}=\prod_{i=1}^{d}S_{(m+1-i)}^{(N)}$.
-We define
-$Y^{(N)}\ :=\ \frac{\log_{B}(\mathfrak{m}_{d}^{(N)})}{\sqrt{N}}\ =\ \sum_{i=1}^{d}\frac{\log_{B}(S^{(N)}_{(m+1-i)})}{\sqrt{N}}\ =\ \sum_{i=1}^{d}Z_{(m+1-i)}^{(N)},$ (3.5)
-To find the PDF of $Y^{(N)}$, we need the joint PDF of $Z_{(m+1-i)}^{(N)}$ for all $1\leq i\leq d$. This is a standard question in order statistics.
-Proposition 3.3.
-[DL] For i.i.d. random variables $X_{1},\dots,X_{k}$ with PDF $f(x)$ and CDF $F(x)$, the joint PDF of $X_{(i)},\dots,X_{(k)}$ where $1\leq i\leq k$ is given by
-$f_{X_{(i)},\dots,X_{(k)}}(x_{i},\dots,x_{k})\ =\ C_{k}^{i}f(x_{i})\cdots f(x_{k})(F(x_{i}))^{i-1},$ (3.6)
-where $C_{k}^{i}:=k!/(i-1)!$.
-Thus, in our case, the joint PDF of $Z_{(m+1-i)}^{(N)}$ for all $1\leq i\leq d$ is
-$f_{Z_{(m+1-d)},\dots,Z_{(m)}}(z_{m+1-d},\dots,z_{m})$
-$\ =\ C_{m}^{m+1-d}\left(\Phi(z_{m+1-d})+B(z_{m+1-d})\right)^{m-d}\prod_{i=1}^{d}\left(\varphi(z_{m+i-d})+A(z_{m+i-d})\right).$ (3.7)
-By binomial expansion,
-$\left(\Phi(z_{m+1-d})+B(z_{m+1-d})\right)^{m-d}\ =\ \sum_{j=0}^{m-d}\binom{m-d}{j}(\Phi(z_{m+1-d}))^{m-d-j}(B(z_{m+1-d}))^{j}.$ (3.8)
-Note that $B(z_{m+1-d})=O(N^{-\delta})$ and $0\leq\Phi(x)\leq 1$, which we shall constantly use without reference for the remainder of the paper.
-Then any term where $j>1$ is at most of the same magnitude as $B(z_{m+1-d})$.
-We have
-$\left(\Phi(z_{m+1-d})+B(z_{m+1-d})\right)^{m-d}\ =\ \left(\Phi(z_{m+1-d})\right)^{m-d}+O(B(z_{m+1-d})).$ (3.9)
-We make a slight abuse of notation here to write the error term above as $B(z_{m+1-d})$, which does not change the order of magnitude of the error term. Hence,
-$f_{Z_{(m+1-d)},\dots,Z_{(m)}}(z_{m+1-d},\dots,z_{m})$
-$\ =\ C_{m}^{m+1-d}\left(\Phi(z_{m+1-d})^{m-d}+B(z_{m+1-d})\right)\prod_{i=1}^{d}\left(\varphi(z_{m+i-d})+A(z_{m+i-d})\right).$ (3.10)
-3.2. PDF of $Y^{(N)}$
-In this section, we find the PDF of $Y^{(N)}$ as defined in (3.1).
+Let $S_{1}^{(N)},\cdots,S_{m}^{(N)}$ be the side lengths of the $m$-dimensional box after stage $N$.
+Then the volume of the largest $d$-dimensional face is the product of the longest $d$ side lengths of the box, i.e., $\mathfrak{m}_{d}^{(N)}=\prod_{i=1}^{d}S_{(m+1-i)}^{(N)}$.
+Since $S_{i}^{(N)}=P_{i}^{(1)}\cdots P_{i}^{(N)}$ for all $i$ and the $P_{i}^{(j)}$’s are between 0 and $1$, then $\mathfrak{m}_{d}^{(N)}$ is asymptotically very small.
+To make this quantity more convenient to work with, we normalize it as follows.
+By Condition 1, $P_{i}^{(j)}$’s are i.i.d., which means that $S_{i}^{(N)}$’s are i.i.d.
+as well.
+Since $\mathbb{E}[\log_{B}(P)]=\mu_{P}$ and $\textup{Var}[\log_{B}(P)]=\sigma_{P}^{2}$, then for all $i$,
+$\mathbb{E}[\log_{B}(S_{i}^{(N)})]\ =\ N\mu_{P}\quad\textup{and}\quad\textup{Var}[\log_{B}(S_{i}^{(N)})]\ =\ N\sigma_{P}^{2}.$ (50)
+Now consider
+$Z_{i}^{(N)}\ :=\ \frac{\log_{B}(S_{i}^{(N)})-N\mu_{P}}{\sqrt{N}\sigma_{P}}.$ (51)
+It is a well-known fact from order statistics that if the order statistics of $(X_{1},\dots,X_{k})$ are $(X_{(1)},\dots,X_{(k)})$, then the order statistics of $(g(X_{1}),\dots,g(X_{k}))$ are $(g(X_{(1)}),\dots,g(X_{(k)}))$ for any monotonically increasing function $g$.
+Since the order statistics of $(S_{1}^{(N)},\dots,S_{m}^{(N)})$ are $(S_{(1)}^{(N)},\dots,S_{(m)}^{(N)})$ and $g(x):=(\log_{B}(x)-N\mu_{P})/\sqrt{N}\sigma_{P}$ is monotonically increasing, then it follows that the order statistics of $(g(S_{1}^{(N)}),\dots,g(S_{m}^{(N)}))$ are $(g(S_{(1)}^{(N)}),\dots,g(S_{(m)}^{(N)}))$, which are just $(Z^{(N)}_{(1)},\dots,Z^{(N)}_{(m)})$.
+We are now ready to define the normalized form of $\mathfrak{m}_{d}^{(N)}$ with which we shall use to establish strong Benford behavior:
+$Y^{(N)}\ :=\ \frac{\log_{B}(\mathfrak{m}_{d}^{(N)})-d\cdot N\mu_{P}}{\sqrt{N}\sigma_{P}}\ =\ \sum_{i=1}^{d}\frac{\log_{B}(S^{(N)}_{(m+1-i)})-N\mu_{P}}{\sqrt{N}\sigma_{P}}\ =\ \sum_{i=1}^{d}Z_{(m+1-i)}^{(N)}.$ (52)
+To find the PDF of $Y^{(N)}$, we need the joint PDF of $Z_{(m+1-i)}^{(N)}$ for all $1\leq i\leq d$, which is a standard question in order statistics.
+Proposition 5.
+[9] For i.i.d. random variables $X_{1},\dots,X_{n}$ with PDF $f(x)$ and CDF $F(x)$, the joint PDF of $X_{(i)},\dots,X_{(n)}$ is given by
+$f_{X_{(i)},\dots,X_{(n)}}(x_{i},\dots,x_{n})\ =\ C_{n}^{i}f(x_{i})\cdots f(x_{n})(F(x_{i}))^{i-1},$ (53)
+where $1\leq i\leq n$ and $C_{n}^{i}:=n!/(i-1)!$.
+Since $Z_{i}^{(N)}$’s are i.i.d., then to find the joint PDF of $Z_{(m+-i)}^{(N)}$, it suffices to find the PDF and CDF of $Z_{i}^{(N)}$’s, which converge to that of the standard Gaussian distribution by CLT.
+However, to gain a sharp estimate on the distribution of $Y^{(N)}$, we would need to quantify the error terms of the PDF and the CDF, as follows.
+To begin, we prove the following Riemann-Lebesgue lemma type bound on the characteristic function of a random variable.
+Lemma 3.
+Suppose that $f(x)$ is the PDF of a random variable $X$, where $f(x)$ is differentiable with $\sup_{x\in\mathbb{R}}|f^{\prime}(x)|<\infty$ and $X$ has finite mean and nonzero, finite variance.
+Then the characteristic function
+$\psi(t)\ :=\ \int_{-\infty}^{\infty}f(x)e^{itx}dx$ (54)
+satisfies $\psi(t)=O(t^{-2/5})$ for any $\epsilon>0$.
+Proof.
+We know that $e^{itx}=\cos(tx)+i\sin(tx)$. Hence, it suffices to show that
+$\int_{-\infty}^{\infty}f(x)\cos(tx)dx\ =\ O(t^{-2/5})\quad\textup{and}\quad\int_{-\infty}^{\infty}f(x)\sin(tx)dx\ =\ O(t^{-2/5}).$ (55)
+The proof for the two bounds are similar, and we only show it for $\int_{-\infty}^{\infty}f(x)\cos(tx)dx$. We first estimate the tail of the integral.
+$\left|\int_{|x|\geq t^{1/5}}f(x)\cos(tx)dx\right|$ $\ \leq\ \int_{|x|\geq t^{1/5}}f(x)dx$
+$\ =\ \mathbb{P}\left(|X|\geq t^{1/5}\right)$
+$\ \leq\ \mathbb{P}\left(|X-\mathbb{E}[X]|\geq t^{1/5}-\mathbb{E}[X]\right)+\mathbb{P}\left(|X-\mathbb{E}[X]|\geq t^{1/5}+\mathbb{E}[X]\right)$
+$\ \leq\ \frac{\textup{Var}[X]}{(t^{1/5}-\mathbb{E}[X])^{2}}+\frac{\textup{Var}[X]}{(t^{1/5}+\mathbb{E}[X])^{2}}\ =\ O(t^{-2/5}),$ (56)
+where the last line follows from Chebyshev’s inequality, which requires $X$ to have finite mean and nonzero, finite variance.
+Hence, we can restrict the domain of integration to $t^{1/5},t^{1/5}]$.
+We know that if $g(x)$ is differentiable on $[a,b]$ with $\sup_{x\in[a,b]}|g^{\prime}(x)|=M<\infty$ [14], then for any Riemann sum $S=\sum_{k=1}^{n}g(x_{k}^{*})(x_{k}-x_{k-1})$,
+$\left|\int_{-\infty}^{\infty}g(x)dx-S\right|\ \leq\ \frac{M(b-a)\delta_{\textup{max}}}{2},$ (57)
+where $\delta_{max}=\max_{k}(x_{k}-x_{k-1})$.
+Now consider a Riemann sum $S_{t}=\sum_{k=1}^{n}f(x_{k}^{*})(x_{k}-x_{k-1})$ with the partition $x_{k}-x_{k-1}=t^{-3/5}$ for all $k$.
+Define the function $f_{t}$ on $t^{1/5},t^{1/5}]$ such that $f_{t}(x):=f(x_{k}^{*})$ for all $x\in[x_{k-1},x_{k})$ and for every $k$.
+Since $f$ is nonnegative, then $f_{t}$ is also nonnegative.
+Hence,
+$\left|\int_{-t^{1/5}}^{t^{1/5}}f(x)\cos(tx)dx-\int_{-t^{1/5}}^{t^{1/5}}f_{t}(x)\cos(tx)dx\right|$ $\ \leq\ \left|\int_{-t^{1/5}}^{t^{1/5}}f(x)dx-S_{t}\right|$
+$\ \leq\ \frac{M_{t}\cdot(2t^{1/5})\cdot t^{-3/5}}{2}\ =\ O(t^{-2/5}),$ (58)
+where $M_{t}:=\sup_{x\in[-t^{1/5},t^{1/5}]}|f^{\prime}(x)|<\infty$. Moreover, we have
+$\int_{-t^{1/5}}^{t^{1/5}}f_{t}(x)\cos(tx)dx$ $\ =\ \int_{-t^{1/5}}^{t^{1/5}}\sum_{k=1}^{n}f(x_{k}^{*})\chi_{[x_{k-1},x_{k})}(x)\cos(tx)dx$
+$\ =\ \sum_{k=1}^{n}f(x_{k}^{*})\int_{x_{k-1}}^{x_{k}}\cos(tx)dx$
+$\ =\ \frac{1}{t}\sum_{k=1}^{n}f(x_{k}^{*})(\sin(tx_{k})-\sin(tx_{k-1}))$
+$\ =\ O\left(\frac{1}{t}\sum_{k=1}^{n}f(x_{k}^{*})\right).$ (59)
+Since $S_{t}=t^{-3/5}\sum_{k=1}^{n}f(x_{k}^{*})$ and that the Riemann sum is finite for large $t$, then
+$\int_{-t^{1/5}}^{t^{1/5}}f_{t}(x)\cos(tx)dx$ $\ =\ O(t^{-2/5}).$ (60)
+Thus, combining (3.1), (3.1), and (60), we have
+$\left|\int_{-\infty}^{\infty}f(x)\cos(tx)dx\right|$
+$\ \leq\ \left|\int_{|x|\geq t^{1/5}}f(x)\cos(tx)dx\right|+\left|\int_{-t^{1/5}}^{t^{1}/5}f(x)\cos(tx)dx-\int_{-t^{1/5}}^{t^{1/5}}f_{t}(x)\cos(tx)dx\right|$
+$\quad\quad+\left|\int_{-t^{1/5}}^{t^{1/5}}f_{t}(x)\cos(tx)dx\right|\ =\ O(t^{-2/5}).$ (61)
+Lemma 4.
+Let $Z_{i}^{(N)}\sim Z^{(N)}$. For any $0<\epsilon<1/2$, the PDF $f_{Z^{(N)}}$ and the CDF $F_{Z^{(N)}}$ of $Z^{(N)}$ satisfy
+$f_{Z^{(N)}}(z)\ =\ \varphi(z)+O(N^{-3/2+3\epsilon})$ (62)
+$F_{Z^{(N)}}(z)\ =\ \Phi(z)+O(N^{-1/2}),$ (63)
+where $\varphi$ and $\Phi$ are the PDF and the CDF of the standard Gaussian distribution.
+Proof.
+We first prove (62). Let the characteristic function of $W:=(\log_{B}(P)-\mu_{P})/\sigma_{P}$ be $\psi_{W}(t)$, which has mean 0, variance $1$, and finite moments by Condition 2. Hence,
+$\psi_{W}(t)$ $\ =\ \sum_{k=0}^{\infty}\frac{\mathbb{E}[W^{k}](it)^{k}}{k!}$
+= $\ =\ 1-\frac{t^{2}}{2}+O(t^{3})$ (64)
+for sufficiently small $t$. Let the characteristic function of $Z^{(N)}$ be $\psi_{Z^{(N)}}(t)$. By properties of characteristic functions,
+$\psi_{Z^{(N)}}(t)$ $\ =\ (\psi_{W}(t/\sqrt{N}))^{N}$
+$\ =\ \left(1-\frac{t^{2}}{2N}+O\left(\frac{t^{3}}{N^{3/2}}\right)\right)^{N}$
+$\ =\ \textup{exp}\left[N\log\left(1-\frac{t^{2}}{2N}+O\left(\frac{t^{3}}{N^{3/2}}\right)\right)\right]$
+$\ =\ \textup{exp}\left[N\left(\log\left(1-\frac{t^{2}}{2N}\right)+O\left(\frac{1}{1-t^{3}/N^{3/2}}\cdot\frac{t^{3}}{N^{3/2}}\right)\right)\right]$
+$\ =\ \textup{exp}\left[N\left(-\frac{t^{2}}{2N}+O\left(\frac{t^{4}}{N^{2}}\right)+O\left(\frac{t^{3}}{N^{3/2}}\right)\right)\right]$
+$\ =\ e^{-t^{2}/2}\left(1+O\left(\frac{t^{3}}{N^{3/2}}\right)\right),$ (65)
+where in the last two lines, we make use of the Taylor series expansions of $\log(1+x)$ and $e^{x}$.
+Note that the above estimate holds for $|t|\leq N^{\epsilon}$ for any $0<\epsilon<1/2$.
+By Fourier inversion, the PDF of $Z^{(N)}$ is
+$\frac{1}{2\pi}\int_{-\infty}^{\infty}\psi_{Z^{(N)}}(t)e^{-itx}dt.$ (66)
+We shall split the integral above into several parts and provide an estimate for each. We first estimate the integral for $|t|\leq N^{\epsilon}$. By (3.1),
+$\frac{1}{2\pi}\int_{|t|\leq N^{\epsilon}}\psi_{Z^{(N)}}(t)e^{-itx}dt\ =\ \varphi(x)+O(N^{-3/2+3\epsilon}).$ (67)
+By Lemma 3, we know that $\psi_{W}(t)=O(t^{-2/5})$.
+Let $E$ be a constant such that $E\geq(\pi/4)^{2/5}$ and $|\psi_{W}(t)|\leq Et^{-2/5}$ for sufficiently large $t$.
+Then, for any $\epsilon_{0}>0$, we have
+$\left|\frac{1}{2\pi}\int_{|t|\geq(E+\epsilon_{0})^{5/2}\sqrt{N}}\psi_{Z^{(N)}}(t)e^{-itx}dt\right|$ $\ \ll\ \int_{t\geq(E+\epsilon_{0})^{5/2}\sqrt{N}}\left(\frac{E}{(t/\sqrt{N})^{2/5}}\right)^{N}dt$
+$\quad\ \ll\ \sqrt{N}E^{N}\int_{t\geq(E+\epsilon_{0})^{5/2}}t^{-2N/5}dt$
+$\quad\ \ll\ \frac{\sqrt{N}}{N}\cdot\frac{E^{N}}{(E+\epsilon_{0})^{N}},$ (68)
+which is of exponential decay in $N$. By [7] and Condition 3, for $|t|\geq\pi/4$, we have
+$|\psi_{W}(t)|\ <\ 1-\frac{c_{1}}{m^{2}},$ (69)
+where $m$ is the median of $W$’s which is finite by Condition 3, and $c_{1}>0$ is some constant. Hence,
+$\left|\frac{1}{2\pi}\int_{\pi\sqrt{N}/4\leq|t|\leq(E+\epsilon_{0})^{5/2}\sqrt{N}}\psi_{Z^{(N)}}(t)e^{-itx}dt\right|\ \ll\ \sqrt{N}\left(1-\frac{c_{1}}{m^{2}}\right)^{N},$ (70)
+which is also of exponential decay in $N$. By [7], when $0<t<\pi/4$, we have
+$|\psi_{W}(t)|\ <\ 1-\frac{c_{2}t^{2}}{m},$ (71)
+where $m$ is again the median of $W_{i}$’s and $c_{2}>0$ is some constant. Hence,
+$\left|\frac{1}{2\pi}\int_{N^{\epsilon}\leq|t|\leq\pi\sqrt{N}/4}\psi_{Z^{(N)}}(t)e^{-itx}dt\right|$ $\ \ll\ \sqrt{N}\left(1-\frac{c_{2}}{m^{2}}N^{2\epsilon-1}\right)^{N}$
+$\ \ll\ \sqrt{N}e^{-2c_{2}N^{2\epsilon}/m^{2}},$ (72)
+which also decays exponential in $N$. Combining (67), (3.1), (70), (3.1), we have
+$f_{Z^{(N)}}(x)\ =\ \frac{1}{2\pi}\int_{-\infty}^{\infty}\psi_{Z^{(N)}}(t)e^{-itx}dt\ =\ \varphi(x)+O(N^{-3/2+3\epsilon}),$ (73)
+which is the desired PDF for $Z^{(N)}$. Moreover, by Berry-Esseen theorem [4, 11], we have that for all $N$ and $x$,
+$F_{Z^{(N)}}(x)\ =\ \Phi(x)+O(N^{-1/2}),$ (74)
+which is the desired estimate for the CDF of $Z^{(N)}$. ∎
+As we shall see for the rest of this section, it suffices for the error term of $f_{Z^{(N)}}$ to be $O(N^{-1/2-\delta})$ and the error term of $F_{Z^{(N)}}$ to be $O(N^{-\delta})$ for some $\delta>0$.
+Hence, the estimates of Lemma 4 are more than enough for this purpose.
+Now, we go back to (53) given in Proposition 5.
+In our case, the joint PDF of $Z_{(m+1-i)}^{(N)}$ for all $1\leq i\leq d$ is
+$f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)}}(z_{m+1-d},\dots,z_{m})$
+$\quad\ =\ C_{m}^{m+1-d}\left(\Phi(z_{m+1-d})+B(z_{m+1-d})\right)^{m-d}\prod_{i=1}^{d}\left(\varphi(z_{m+i-d})+A(z_{m+i-d})\right),$ (75)
+where $A(x)=O(N^{-1/2-\delta})$ and $B(x)=O(N^{-\delta})$. By binomial expansion,
+$\left(\Phi(z_{m+1-d})+B(z_{m+1-d})\right)^{m-d}\ =\ \sum_{j=0}^{m-d}\binom{m-d}{j}(\Phi(z_{m+1-d}))^{m-d-j}(B(z_{m+1-d}))^{j}.$ (76)
+Since $\Phi(x)$ is a CDF, then $0\leq\Phi(x)\leq 1$, a fact which we shall use without explicit mention for the remainder of the paper.
+Then, any term in (76) with $j>1$ is at most of the same order as $B(z_{m+1-d})$.
+Hence,
+$\left(\Phi(z_{m+1-d})+B(z_{m+1-d})\right)^{m-d}\ =\ \left(\Phi(z_{m+1-d})\right)^{m-d}+O(B(z_{m+1-d})).$ (77)
+We make a slight abuse of notation here to write the error term above as $B(z_{m+1-d})$, which does not change the order of the error term. So then we have
+$f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)}}(z_{m+1-d},\dots,z_{m})$
+$\quad\ =\ C_{m}^{m+1-d}\left(\Phi(z_{m+1-d})^{m-d}+B(z_{m+1-d})\right)\prod_{i=1}^{d}\left(\varphi(z_{m+i-d})+A(z_{m+i-d})\right).$ (78)
+3.2 PDF of $Y^{(N)}$
+In this section, we find the PDF of $Y^{(N)}$.
 The CDF $F_{Y^{(N)}}(y)$ of $Y^{(N)}$ is given by integrating over the appropriate region in $\mathbb{R}^{d}$, that is, over all the values $(z_{m+1-d},\dots,z_{m})$ of $(Z_{(m+1-d)}^{(N)},\dots,Z_{(m)}^{(N)})$ that sums to $y$.
@@
 In general, we know that any $1\leq j\leq d$, $z_{m+j-d}$ can take value from $z_{m+(j-1)-d}$ to $(y-\sum_{i=1}^{j-1}z_{m+i-d})/(d-j+1)$.
-The only exception is $z_{m+1-d}$, whose upper bound is already correctly stated but the lower bound is $C\sqrt{N}$.
+The only exception is $z_{m+1-d}$, whose upper bound is already correctly stated but the lower bound is -$\infty$.
 Hence, we have
 $F_{Y^{(N)}}(y)$
-$\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z_{(m+1-d)},\dots,Z_{(m)}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d},$ (3.11)
+$\quad\ =\ \int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d},$ (79)
 where we use the product of integrals to denote
-$\prod_{j=k}^{d}\int_{z_{j}}^{y_{j}}\ :=\ \int_{z_{k}}^{y_{k}}\cdots\int_{z_{j}}^{y_{j}}\cdots\int_{z_{d}}^{y_{d}}.$ (3.12)
+$\prod_{j=k}^{d}\int_{z_{j}}^{y_{j}}\ :=\ \int_{z_{k}}^{y_{k}}\cdots\int_{z_{j}}^{y_{j}}\cdots\int_{z_{d}}^{y_{d}}.$ (80)
 We separate out the main term $C_{m}^{m+1-d}\Phi(z_{m+1-d})^{m-d}\prod_{i=1}^{d}\varphi(z_{m+i-d})$ of the joint PDF (3.1) and denote it by
-$m_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})\ :=\ C_{m}^{m+1-d}\Phi(z_{m+1-d})^{m-d}\prod_{i=1}^{d}\varphi(z_{m+i-d}).$ (3.13)
-We also denote the error term of the joint PDF (3.1) by $e_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})$.
-Let $\mathcal{M}_{Y^{(N)}}(y)$ and $\mathcal{E}_{Y^{(N)}}(y)$ be the main term and the error term of $F_{Y^{(N)}}(y)$.
+$f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ main}}(z_{m+1-d},\dots,z_{m})\ :=\ C_{m}^{m+1-d}\Phi(z_{m+1-d})^{m-d}\prod_{i=1}^{d}\varphi(z_{m+i-d}).$ (81)
+We also denote the error term of the joint PDF (3.1) by $f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ error}}(z_{m+1-d},\dots,z_{m})$.
+Let $F_{Y^{(N)},\textup{ main}}(y)$ and $F_{Y^{(N)},\textup{ error}}(y)$ be the main term and the error term of $F_{Y^{(N)}}(y)$.
 Then
-$F_{Y^{(N)}}(y)$ $\ =\ \mathcal{M}_{Y^{(N)}}(y)+\mathcal{E}_{Y^{(N)}}(y)$
-$\mathcal{M}_{Y^{(N)}}(y)$ $\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}m_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}$
-$\mathcal{E}_{Y^{(N)}}(y)$ $\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}e_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}.$ (3.14)
+$F_{Y^{(N)}}(y)$ $\ =\ F_{Y^{(N)},\textup{ main}}(y)+F_{Y^{(N)},\textup{ error}}(y)$
+$F_{Y^{(N)},\textup{ main}}(y)$ $\ =\ \int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ main}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}$
+$F_{Y^{(N)},\textup{ error}}(y)$ $\ =\ \int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ error}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}.$ (82)
 We are now ready to find the PDF $f_{Y^{(N)}}$ of $Y^{(N)}$. We know that
-$f_{Y^{(N)}}(y)\ =\ \frac{d}{dy}F_{Y^{(N)}}(y)\ =\ \frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)+\frac{d}{dy}\mathcal{E}_{Y^{(N)}}(y).$ (3.15)
-To differentiate $\mathcal{M}_{Y^{(N)}}(y)$ and $\mathcal{E}_{Y^{(N)}}(y)$, we apply the Leibniz integral rule for differentiate under the integral sign.
-Lemma 3.4.
-[F] Suppose that $f(x,t)$ is a function such that both $f(x,t)$ and $df(x,t)/dx$ are continuous in $t$ and $x$ in the $xt$-plane, and $a(x)$ and $b(x)$ are also continuously differentiable. Then
-$\frac{d}{dx}\int_{a(x)}^{b(x)}f(x,t)dt\ =\ f(x,b(x))\cdot\frac{d}{dx}b(x)-f(x,a(x))\cdot\frac{d}{dx}a(x)+\int_{a(x)}^{b(x)}\frac{\partial}{\partial x}f(x,t)dt.$ (3.16)
-For the rest of this paper, one can easily check the regularity conditions on $f(x,t)$, $a(x)$, and $b(x)$ each time we apply the Leibniz integral rule, so there will not be explicit mention of the regularity conditions again.
+$f_{Y^{(N)}}(y)\ =\ \frac{d}{dy}F_{Y^{(N)}}(y)\ =\ \frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y)+\frac{d}{dy}F_{Y^{(N)},\textup{ error}}(y).$ (83)
+To differentiate $F_{Y^{(N)}(y),\textup{ main}}$ and $F_{Y^{(N)},\textup{ error}}(y)$, we apply the Leibniz integral rule for differentiate under the integral sign.
+Lemma 5.
+[12] Suppose that $f(x,t)$ is a function such that both $f(x,t)$ and $df(x,t)/dx$ are continuous in $t$ and $x$ in the $xt$-plane, and $a(x)$ and $b(x)$ are also continuously differentiable. Then
+$\frac{d}{dx}\int_{a(x)}^{b(x)}f(x,t)dt\ =\ f(x,b(x))\cdot\frac{d}{dx}b(x)-f(x,a(x))\cdot\frac{d}{dx}a(x)+\int_{a(x)}^{b(x)}\frac{\partial}{\partial x}f(x,t)dt.$ (84)
+For the rest of this paper, one can easily check the regularity conditions on $f(x,t)$, $a(x)$, and $b(x)$ each time we apply the Leibniz integral rule, so there will not be any explicit mention of the regularity conditions again.
 We first make some observations on the effect of $d/dy$ on
-$\int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}g(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}$ (3.17)
-for any Riemann integrable function $g(z_{m+1-d},\dots,z_{m})$.
-One can check that each application of the Leibniz integral rule to (3.17) gives two integrals, one of which has the lower bound of the integration coinciding with the upper bound of the integration and thus equals to 0.
-As a result, we are always left with one integral that interchanges the order of the differentiation and the integration signs.
-Eventually, (3.17) becomes
-$\int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\frac{\partial}{\partial y}\left(\int_{z_{m-1}}^{y-\sum_{i=1}^{d-1}z_{m+i-d}}g(z_{m+1-d},\dots,z_{m})dz_{m}\right)dz_{m-1}\cdots dz_{m+1-d}$
-$\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}g\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\cdots dz_{m+1-d}.$ (3.18)
-Applying (3.2) to $d\mathcal{M}_{Y^{(N)}}(y)/dy$ and $d\mathcal{E}_{Y^{(N)}}(y)/dy$, where $\mathcal{M}_{Y^{(N)}}(y)$ and $\mathcal{E}_{Y^{(N)}}(y)$ are as defined in (3.2), we have
-$\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)$ $\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}m_{Y^{(N)}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}$
-$\quad\quad\cdots dz_{m+1-d}$
-$\frac{d}{dy}\mathcal{E}_{Y^{(N)}}(y)$ $\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}e_{Y^{(N)}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}$
-$\quad\quad\cdots dz_{m+1-d}.$ (3.19)
-3.3. Outline of problem
-We are ready to formulate our problem.
+$\int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}g(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}$ (85)
+for any continuous function $g(z_{m+1-d},\dots,z_{m})$.
+One can check that each application of the Leibniz integral rule to (85) would only yield one integral, because for the other integral, the lower bound of the integration always coincides with the upper bound of the integration and thus equals to 0.
+As a result, each application of $d/dy$ here merely interchanges the order of the differentiation and the integral signs.
+Eventually, (85) becomes
+$\int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\frac{\partial}{\partial y}\left(\int_{z_{m-1}}^{y-\sum_{i=1}^{d-1}z_{m+i-d}}g(z_{m+1-d},\dots,z_{m})dz_{m}\right)dz_{m-1}\cdots dz_{m+1-d}$
+$\quad\ =\ \int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}g\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\cdots dz_{m+1-d}.$ (86)
+Applying (3.2) to $dF_{Y^{(N)}}/dy$, we have that the PDF of $Y^{(N)}$ is
+$\frac{d}{dy}F_{Y^{(N)}}(y)\ =\ \int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
+$\quad\quad\quad\quad\quad\quad\quad dz_{m-1}\cdots dz_{m+1-d}.$ (87)
+Applying (3.2) to $dF_{Y^{(N)},\textup{ main}}(y)/dy$ and $dF_{Y^{(N)},\textup{ error}}(y)/dy$, we have
+$\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y)$
+$\quad\ =\ \int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ main}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
+$\quad\quad\quad dz_{m-1}\cdots dz_{m+1-d}$
+$\frac{d}{dy}F_{Y^{(N)},\textup{ error}}(y)$
+$\quad\ =\ \int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ error}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
+$\quad\quad\quad dz_{m-1}\cdots dz_{m+1-d}$
+We would like to make some truncation to the tails of integrals in (3.2) and (3.2).
+The truncation method we now describe applies to both $dF_{Y^{(N)}}(y)/dy$ and $dF_{Y^{(N)},\textup{ main}}(y)/dy$, so we only demonstrate it for $dF_{Y^{(N)}}(y)/dy$.
+We want to show that
+$\int_{-\infty}^{-N^{\delta^{\prime}}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
+$dz_{m-1}\cdots dz_{m+1-d}$ (89)
+is asymptotically small for any $\delta^{\prime}$ satisfying $0<\delta^{\prime}<\delta$. By Proposition 5,
+$f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
+$\quad\ =\ C_{m}^{m+1-d}f_{Z^{(N)}}(z_{m+1-d})\cdots f_{Z^{(N)}}(z_{m-1})f_{Z^{(N)}}\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)(F_{Z^{(N)}}(z_{m+1-d}))^{m-d}.$ (90)
+By Lemma 4, we know that $f_{Z^{(N)}}(z)=\varphi(z)+O(N^{-1/2-\delta})$, and $F_{Z^{(N)}}(z)=\Phi(z)+O(N^{-\delta})$, where $\varphi(z)=1/\sqrt{2\pi}e^{-z^{2}/2}$ is bounded between 0 and $1/\sqrt{2\pi}$ and $0\leq\Phi(z)\leq 1$, both of which are $O(1)$.
+Then,
+$f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)\ \ll\ f_{Z^{(N)}}(z_{m+1-d})\cdots f_{Z^{(N)}}(z_{m-1}).$ (91)
+Returning to (3.2), we have that
+$\int_{-\infty}^{-N^{\delta^{\prime}}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
+$dz_{m-1}\cdots dz_{m+1-d}$
+$\quad\ \ll\ \int_{-\infty}^{-N^{\delta^{\prime}}}\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}f_{Z^{(N)}}(z_{m+1-d})f_{Z^{(N)}}(z_{m+2-d})\cdots f_{Z^{(N)}}(z_{m-1})$
+$\quad\quad\quad dz_{m-1}\cdots dz_{m+2-d}dz_{m+1-d}$
+$\quad\ =\ \int_{-\infty}^{-N^{\delta^{\prime}}}f_{Z^{(N)}}(z_{m+1-d})dz_{m+1-d}\prod_{i=2}^{d-1}\left(\int_{-\infty}^{\infty}f_{Z^{(N)}}(z_{m+i-d})dz_{m+i-d}\right).$ (92)
+Since $f_{Z^{(N)}}(z)$ is a PDF, then $\int_{-\infty}^{\infty}f_{Z^{(N)}}(z)dz=1$. Moreover,
+$\int_{-\infty}^{-N^{\delta^{\prime}}}f_{Z^{(N)}}(z_{m+1-d})dz_{m+1-d}$ $\ =\ \mathbb{P}\left(Z^{(N)}\leq-N^{\delta^{\prime}}\right)$
+$\ \leq\ \mathbb{P}\left(|Z^{(N)}-\mathbb{E}[Z^{(N)}]|\geq N^{\delta^{\prime}}+\mathbb{E}[Z^{(N)}]\right)$
+$\ \leq\ \frac{\textup{Var}[Z^{(N)}]}{(N^{\delta^{\prime}}+\mathbb{E}[Z^{(N)}])^{2}}\ =\ O(N^{-2\delta^{\prime}}),$ (93)
+where the last line follows from Chebyshev’s inequality and the fact that the mean and the variance of $Z^{(N)}$ are 0 and $1$ respectively by construction (50) and (51).
+Hence, we have established that the tail of $dF_{Y^{(N)}}(y)/dy$ is $O(N^{-2\delta^{\prime}})$, i.e.,
+$\frac{d}{dy}F_{Y^{(N)}}(y)$
+$\quad\ =\ \int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
+$\quad\quad\quad dz_{m-1}\cdots dz_{m+1-d}+O(N^{-2\delta^{\prime}}),$ (94)
+By the same truncation method, one can show that the tail of $dF_{Y^{(N)},\textup{ main}}(y)/dy$ is also $O(N^{-2\delta^{\prime}})$.
+The tail of $dF_{Y^{(N)},\textup{ error}}(y)/dy$ is the difference between the tail of $dF_{Y^{(N)}}(y)/dy$ and that of $dF_{Y^{(N)},\textup{ main}}(y)/dy$ by definition, then it is also $O(N^{-2\delta^{\prime}})$.
+Thus,
+$\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y)$
+$\quad\ =\ \int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ main}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
+$\quad\quad\quad dz_{m-1}\cdots dz_{m+1-d}+O(N^{-2\delta^{\prime}}),$ (95)
+$\frac{d}{dy}F_{Y^{(N)},\textup{ error}}(y)$
+$\quad\ =\ \int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ error}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
+$\quad\quad\quad dz_{m-1}\cdots dz_{m+1-d}+O(N^{-2\delta^{\prime}}).$ (96)
+3.3 Outline of problem
+We are ready to formulate Conjecture 1 in the language that we have set forth in Section 3.1.
 We want to show that $\mathfrak{m}_{d}^{(N)}$ converges to strong Benford behavior, which is equivalent to showing that $\log_{B}(\mathfrak{m}_{d}^{(N)})$ converges to being equidistributed mod 1.
-Hence, in terms of the probability density function $f_{Y^{(N)}}(y)$ of $Y^{(N)}:=\log_{B}(\mathfrak{m}_{d}^{(N)})/\sqrt{N}$, we can formulate our problem as to showing
-$F_{N}(a,b)\ :=\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}f_{Y^{(N)}}(y)dy\ \approx\ b-a,$ (3.20)
-for all $(a,b)\subset(0,1)$ for sufficiently large $N$. Let the main term and the error term of $F_{N}(a,b)$ be defined respectively as
-$\mathcal{M}_{N}(a,b)$ $\ :=\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)dy;$ (3.21)
-$\mathcal{E}_{N}(a,b)$ $\ :=\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}\mathcal{E}_{Y^{(N)}}(y)dy.$ (3.22)
+We know that $Y^{(N)}:=(\log_{B}(\mathfrak{m}_{d}^{(N)})-d\cdot N\mu_{P})/\sqrt{N}\sigma_{P}$.
+Since equidistribution mod 1 is invariant under translation [17], then it suffices to show that $\log_{B}(\mathfrak{m}_{d}^{(N)})-d\cdot N\mu_{P}$ converges to being equidistributed mod 1.
+Moreover, without loss of generality, we can assume $\sigma_{P}=1$ from now on, as the choice of renormalization would not change our final results.
+Hence, in terms of the PDF $f_{Y^{(N)}}(y)$, we can formulate our problem as to showing
+$F_{N}(a,b)\ :=\ \mathbb{P}\left(Y^{(N)}\textup{ mod 1}\in(a,b)\right)\ =\ \sum_{n=-\infty}^{\infty}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}f_{Y^{(N)}}(y)dy\ \approx\ b-a,$ (97)
+for all $(a,b)\subset(0,1)$ for sufficiently large $N$.
+Now, recall that we have truncated the outermost integral of $f_{Y^{(N)}}(y)=dF_{Y^{(N)}}(y)/dy$ at $N^{\delta^{\prime}}$ in Section 3.1, which is exactly the lower bound for $Z_{(m+1-d)}$.
+Since $Z_{(m+1-d)}\leq Z_{(m+2-d)}\leq\cdots\leq Z_{(m)}$, then the lower bound for their sum, $Y^{(N)}$, is $dN^{\delta^{\prime}}$.
+Hence, it is reasonable to truncate the sum of $F_{N}(a,b)$ at $n=\pm dN^{1/2+\delta^{\prime}}$ and instead show that
+$\sum_{n=-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}f_{Y^{(N)}}(y)dy\ \approx\ b-a.$ (98)
+We can make the truncation because the tails of the sum is negligible:
+$\int_{\left|y\right|\geq dN^{\delta^{\prime}}}f_{Y^{(N)}}(y)dy$ $\ =\ \mathbb{P}\left(\left|Y^{(N)}\right|\geq dN^{\delta^{\prime}}\right)$
+$\ \leq\ \mathbb{P}\left(\left|Y^{(N)}-\mathbb{E}[Y^{(N)}]\right|\geq dN^{\delta^{\prime}}-\mathbb{E}[Y^{(N)}]\right)$
+$\quad\quad+\mathbb{P}\left(\left|Y^{(N)}-\mathbb{E}[Y^{(N)}]\right|\geq dN^{\delta^{\prime}}+\mathbb{E}[Y^{(N)}]\right)$
+$\ \leq\ \frac{\textup{Var}[Y^{(N)}]}{(dN^{\delta^{\prime}}-\mathbb{E}[Y^{(N)}])^{2}}+\frac{\textup{Var}[Y^{(N)}]}{(dN^{\delta^{\prime}}+\mathbb{E}[Y^{(N)}])^{2}}\ =\ O(N^{-2\delta^{\prime}}),$ (99)
+where the last line follows from Chebyshev’s inequality and the fact that the mean and the variance of $Y^{(N)}$ are finite, as it is the sum of $d$ random variables whose means and variances are finite.
+From now on, we shall only consider the bulk part of $F_{N}(a,b)$.
+Let the main term and the error term of $F_{N}(a,b)$ be defined respectively as
+$\mathcal{M}_{N}(a,b)$ $\ :=\ \sum_{n=-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y)dy;$ (100)
+$\mathcal{E}_{N}(a,b)$ $\ :=\ \sum_{n=-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}F_{Y^{(N)},\textup{ error}}(y)dy.$ (101)
 Our goal is to show
-$\mathcal{M}_{N}(a,b)\ \approx\ b-a,\quad\mathcal{E}_{N}(a,b)\ \approx\ 0.$ (3.23)
-3.4. Upper bound on error term $\mathcal{E}_{N}(a,b)$
-We want to show that the error term $\mathcal{E}_{N}(a,b)$ is negligible.
-We start by giving some bounds on $d\mathcal{E}_{Y^{(N)}}(y)/dy$ from (3.2).
-Recall that $e_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})$ is the error term of the joint density function (3.1).
+$\mathcal{M}_{N}(a,b)\ \approx\ b-a,\quad\mathcal{E}_{N}(a,b)\ \approx\ 0.$ (102)
+3.4 Upper bound on error term $\mathcal{E}_{N}(a,b)$
+We first want to show that the error term $\mathcal{E}_{N}(a,b)$ is negligible.
+We start by giving some bounds on $dF_{Y^{(N)},\textup{error}}(y)/dy$ from (3.2).
+Recall that $f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ error}}(z_{m+1-d},\dots,z_{m})$ is the error term of the joint density function (3.1).
 Let $\ell$] be short hand for $\{1,\dots,\ell\}$.
 Then
-$\frac{1}{C_{m}^{m+1-d}}e_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})$
+$\frac{1}{C_{m}^{m+1-d}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ error}}(z_{m+1-d},\dots,z_{m})$
 $\quad\ =\ \Phi(z_{m+1-d})^{m-d}\left(\prod_{i=1}^{d}(\varphi(z_{m+i-d})+A(z_{m+i-d}))-\prod_{i=1}^{d}\varphi(z_{m+i-d})\right)$
 $\quad\quad\quad+B(z_{m+1-d})\prod_{i=1}^{d}(\varphi(z_{m+i-d})+A(z_{m+i-d}))$
-$\quad\ =\ \left(\Phi(z_{m+1-d})^{m-d}+B(z_{m+1-d})\right)\left(\sum_{\begin{subarray}{c}S\subset[d]\\ S\neq d\end{subarray}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\right)$
+$\quad\ =\ \left(\Phi(z_{m+1-d})^{m-d}+B(z_{m+1-d})\right)\left(\sum_{S\subsetneq[d]}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\right)$
 $\quad\quad\quad+B(z_{m+1-d})\prod_{i=1}^{d}\varphi(z_{m+i-d})$
-$\ \ll\ \sum_{\begin{subarray}{c}S\subset[d]\\ S\neq d\end{subarray}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)+N^{-\delta}\prod_{i=1}^{d}\varphi(z_{m+i-d}).$ (3.24)
-We know from (3.2) that to bound $d\mathcal{E}_{Y^{(N)}}(y)/dy$, it suffices to bound
-$D_{1}(y)$ $\ :=\ \sum_{\begin{subarray}{c}S\subset[d]\\ S\neq[d]\end{subarray}}\int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\Bigg{|}_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}$
-$\quad\quad dz_{m-1}\cdots dz_{m+1-d}$ (3.25)
+$\quad\ \ll\ \sum_{S\subsetneq[d]}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)+N^{-\delta}\prod_{i=1}^{d}\varphi(z_{m+i-d}).$ (103)
+We know from (3.2) that to bound $dF_{Y^{(N)},\textup{ error}}(y)/dy$, it suffices to bound
+$D_{1}(y)$ $\ :=\ \sum_{S\subsetneq[d]}\int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)$
+$\quad\quad\cdot\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\Bigg|_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}dz_{m-1}\cdots dz_{m+1-d}$ (104)
 as well as
-$D_{2}(y)\ :=\ N^{-\delta}\int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\prod_{i=1}^{d}\varphi(z_{m+i-d})\Bigg{|}_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}dz_{m-1}\cdots dz_{m+1-d},$ (3.26)
-because then we have $d\mathcal{E}_{Y^{(N)}}dy\ll D_{1}(y)+D_{2}(y)$.
+$D_{2}(y)\ :=\ N^{-\delta}\int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\prod_{i=1}^{d}\varphi(z_{m+i-d})\Bigg|_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}dz_{m-1}\cdots dz_{m+1-d},$ (105)
+because then by (3.2), we have
+$\frac{d}{dy}F_{Y^{(N)},\textup{ error}}(y)\ \ll\ D_{1}(y)+D_{2}(y)+O(N^{-2\delta^{\prime}}).$ (106)
 We first look at $D_{1}(y)$.
@@
 Such an index $k$ exists, since $S\neq[d]$.
-First, assume that we have $S$’s where $k_{S}=d$.
+First, consider $S$’s with $k_{S}=d$.
 Then the integrand of (3.4) becomes
@@
 $\quad\ =\ \left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d-1]\setminus S}A(z_{m+i-d})\right)A\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
-$\quad\ \ll\ N^{-\frac{1}{2}-\delta}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d-1]\setminus S}A(z_{m+i-d})\right),$ (3.27)
+$\quad\ \ll\ N^{-\frac{1}{2}-\delta}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d-1]\setminus S}A(z_{m+i-d})\right),$ (107)
 where in the last line we use the fact that $A(x)=O(N^{-1/2-\delta})$.
-Note that the lower and upper bounds on the interval of integration for each integral in (3.4) are $O(\sqrt{N})$, since $y,z_{m+1-d},\dots,z_{m-1}$ are all $O(\sqrt{N})$.
-Choose a constant $C^{\prime}$ such that $C^{\prime}\sqrt{N}$ is an upper bound on the upper bounds on the interval of integration for all the integrals in (3.4) and $C^{\prime}\sqrt{N}$ is a lower bound on the interval of integration for all the integrals in (3.4).
+Now, recall the two truncations we performed in (3.2) and in (3.3).
+These truncations ensure that $z_{m+1-d}$ is at least $N^{\delta^{\prime}}$ and that $y$ is bounded above by $dN^{\delta^{\prime}}$ and below by $dN^{\delta^{\prime}}$.
+Since $y=z_{m+1-d}+\cdots+z_{m}$ and $z_{m+1-d}\leq\cdots\leq z_{m}$, then $z_{m}$ is at most $(2d-1)N^{\delta^{\prime}}$, which occurs when $y=dN^{\delta^{\prime}}$ and $z_{m+1-d}=\cdots=z_{m-1}=-N^{\delta^{\prime}}$.
+Hence, $y,z_{m+1-d},\dots,z_{m-1}$ are all $O(N^{\delta^{\prime}})$, and the lower and upper bounds on the domain of integration for each integral in (3.4) are $O(N^{\delta^{\prime}})$.
+Choose a constant $C$ such that $\pm CN^{\delta^{\prime}}$ are uniform upper and lower bounds on the domains of integration for all the integrals in (3.4).
 Hence,
-$\int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\Bigg{|}_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}$
+$\int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\Bigg|_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}$
 $dz_{m-1}\cdots dz_{m+1-d}$
-$\quad\ \ll\ N^{-1/2-\delta}\underbrace{\int_{-C^{\prime}\sqrt{N}}^{C^{\prime}\sqrt{N}}\cdots\int_{-C^{\prime}\sqrt{N}}^{C\sqrt{N}}}_{d-1}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d-1]\setminus S}N^{-1/2-\delta}\right)dz_{m-1}\cdots dz_{m+1-d}$
-$\quad\ \ll\ N^{-1/2-\delta-|[d-1]\setminus S|\delta}\prod_{i\in S}\left(\int_{-C^{\prime}\sqrt{N}}^{C^{\prime}\sqrt{N}}\varphi(z_{m+i-d})dz_{m+i-d}\right)\ \ll\ N^{-1/2-(1+|[d-1]\setminus S|)\delta}\ \ll\ N^{-1/2-\delta},$ (3.28)
-where in the last line we use the fact that
-$\int_{-C^{\prime}\sqrt{N}}^{C^{\prime}\sqrt{N}}\varphi(x)dx\ \ll\ \int_{-\infty}^{\infty}\varphi(x)dx\ =\ 1,$ (3.29)
-which we shall use frequently without reference from now on. Since the number of subsets $S$ of $[d]$ is finite, then (3.4) gives us
-$D_{1}(y)\ \ll\ N^{-1/2-\delta}.$ (3.30)
-The case when $k_{S}<d$ is similar, though more care is needed to identify the $O(N^{-1/2-\delta})$ decay, which essentially comes from the $A(z_{m+k_{S}-d})$ term.
+$\quad\ \ll\ N^{-1/2-\delta}\underbrace{\int_{-CN^{\delta^{\prime}}}^{CN^{\delta^{\prime}}}\cdots\int_{-CN^{\delta^{\prime}}}^{CN^{\delta^{\prime}}}}_{d-1}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d-1]\setminus S}N^{-1/2-\delta}\right)dz_{m-1}\cdots dz_{m+1-d}$
+$\quad\ \ll\ N^{-1/2-\delta+|[d-1]\setminus S|(-1/2-\delta+\delta^{\prime})}\prod_{i\in S}\left(\int_{-CN^{\delta^{\prime}}}^{CN^{\delta^{\prime}}}\varphi(z_{m+i-d})dz_{m+i-d}\right)$
+$\quad\ \ll\ N^{-1/2-\delta+|[d-1]\setminus S|(-1/2-\delta+\delta^{\prime})}\ \ll\ N^{-1/2-\delta},$ (108)
+where in the last line we use the fact that $0<\delta^{\prime}<\delta$, as well as
+$\int_{-CN^{\delta^{\prime}}}^{CN^{\delta^{\prime}}}\varphi(x)dx\ \ll\ \int_{-\infty}^{\infty}\varphi(x)dx\ =\ 1,$ (109)
+which we shall use without reference for the remainder of the paper. Since there is only a finite number of subsets $S$ of $[d]$, then (3.4) has the estimate
+$D_{1}(y)\ \ll\ N^{-1/2-\delta}.$ (110)
+We move on to the case when $k_{S}<d$, which is similar to the $k_{S}=d$ case.
+Now, the $O(N^{-1/2-\delta})$ decay that we need comes from the $A(z_{m+k_{S}-d})$ term.
 With this in mind, using the same method for the case when $k_{S}=d$ as well as the convolution formula for Gaussian PDF, we can again show that $D_{1}(y)\ll N^{-1/2-\delta}$.
-We leave the details to Appendix B.
-We now turn to $D_{2}(y)$ as defined in (3.26), which is relatively more straightforward to bound. We have
-$D_{2}(y)\ \ll\ N^{-\delta}\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-1}\prod_{i=1}^{d-1}\varphi(z_{m+i-d})\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\cdots dz_{m+1-d}.$ (3.31)
+We leave the details to B.
+We now turn to $D_{2}(y)$ as defined in (105), which is relatively more straightforward to bound. We have
+$D_{2}(y)\ \ll\ N^{-\delta}\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-1}\prod_{i=1}^{d-1}\varphi(z_{m+i-d})\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\cdots dz_{m+1-d}.$ (111)
 We recognize that the integral is the convolution of $d$ standard Gaussian density functions.
@@
 Hence,
-$D_{2}(y)\ \ll\ N^{-\delta}\frac{1}{\sqrt{2\pi d}}e^{-y^{2}/(2d)}.$ (3.32)
-Thus,
-$\frac{d}{dy}\mathcal{E}_{Y^{(N)}}(y)$ $\ \ll\ D_{1}(y)+D_{2}(y)\ \ll\ N^{-\delta}\frac{1}{\sqrt{2\pi d}}e^{-y^{2}/(2d)}+N^{-1/2-\delta}.$ (3.33)
-We are now ready to bound the error term $\mathcal{E}_{N}(a,b)$ from (3.22):
-$\mathcal{E}_{N}(a,b)$ $\ =\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}\mathcal{E}_{Y^{(N)}}(y)dy$
-$\ \ll\ N^{-\delta}\int_{-dC\sqrt{N}}^{dC\sqrt{N}}\frac{1}{\sqrt{2\pi d}}e^{-y^{2}/(2d)}dy+\int_{-dC\cdot\sqrt{N}}^{dC\cdot\sqrt{N}}N^{-1/2-\delta}dy\ \ll\ N^{-\delta}.$ (3.34)
-3.5. Strategy for remainder of proof
-Our next task is to show that the main term $\mathcal{M}_{N}(a,b)$ defined in (3.21) satisfies
-$\mathcal{M}_{N}(a,b)\ :=\ \sum_{-dC\cdot N}^{dC\cdot N-1}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)dy\ \approx\ b-a.$ (3.35)
+$D_{2}(y)\ \ll\ N^{-\delta}\frac{1}{\sqrt{2\pi d}}e^{-y^{2}/(2d)}.$ (112)
+Thus, combining the estimates (110), (B) on $D_{1}(y)$, the estimate (112) on $D_{2}(y)$, as well as (106), we have
+$\frac{d}{dy}F_{Y^{(N)},\textup{ error}}(y)$ $\ \ll\ D_{1}(y)+D_{2}(y)+N^{-2\delta^{\prime}}\ \ll\ N^{-1/2-\delta}+N^{-\delta}\frac{1}{\sqrt{2\pi d}}e^{-y^{2}/(2d)}+N^{-2\delta^{\prime}}.$ (113)
+We are now ready to bound the error term $\mathcal{E}_{N}(a,b)$ from (101):
+$\mathcal{E}_{N}(a,b)$ $\ =\ \sum_{n=-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}F_{Y^{(N)},\textup{ error}}(y)dy$
+$\ \ll\ N^{-\delta}\int_{-dN^{\delta^{\prime}}}^{dN^{\delta^{\prime}}}\frac{1}{\sqrt{2\pi d}}e^{-y^{2}/(2d)}dy+\int_{-dN^{\delta^{\prime}}}^{dN^{\delta^{\prime}}}N^{-1/2-\delta}+N^{-2\delta^{\prime}}dy\ \ll\ N^{-\delta^{\prime}},$ (114)
+where we use the fact that $0<\delta^{\prime}<\delta<1/2$.
+3.5 Strategy for remainder of proof
+Our next task is to show that the main term $\mathcal{M}_{N}(a,b)$ defined in (100) satisfies
+$\mathcal{M}_{N}(a,b)\ :=\ \sum_{-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y)dy\ \approx\ b-a.$ (115)
 Our strategy is the following.
-First, we want to write the integral of $d\mathcal{M}_{Y^{(N)}}(y)/dy$ over each interval $(a+n)/\sqrt{N},(b+n)/\sqrt{N}]$ as the sum of a main term $m_{n}(a,b)$ that is constant with respect to $y$ within the interval and an error term $e_{n}(a,b)$ for each $n$.
-Essentially, this shows that $d\mathcal{M}_{Y^{(N)}}(y)/dy$ is almost constant within a small interval, incurring an negligible cost that, as we shall prove, does not accumulate, i.e.,
-$\mathcal{M}_{\textup{err},N}(a,b)\ =\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}e_{n}(a,b)\ \approx\ 0.$ (3.36)
+First, we want to write the integral of $dF_{Y^{(N)},\textup{ main}}(y)/dy$ over each interval $(a+n)/\sqrt{N},(b+n)/\sqrt{N}]$ as the sum of a constant main term $m_{n}(a,b)$ together with an error term $e_{n}(a,b)$ which, as we shall prove, does not accumulate, i.e.,
+$\mathcal{M}_{N,\textup{ error}}(a,b)\ =\ \sum_{n=-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}e_{n}(a,b)\ \approx\ 0.$ (116)
 Finally, we prove that the sum of the main term
-$\mathcal{M}_{\textup{main},N}(a,b)\ =\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}m_{n}(a,b)$ (3.37)
+$\mathcal{M}_{N,\textup{ main}}(a,b)\ =\ \sum_{n=-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}m_{n}(a,b)$ (117)
 is a Riemann sum that converges to bab-a. It then follows that
-$\mathcal{M}_{N}(a,b)\ =\ \mathcal{M}_{\textup{main},N}(a,b)+\mathcal{M}_{\textup{err},N}(a,b)\ \approx\ b-a.$ (3.38)
-Thus, based on the estimate on the error term in (3.22),
-$F_{N}(a,b)\ =\ \mathcal{M}_{N}(a,b)+\mathcal{E}_{N}(a,b)\ \approx\ b-a,$ (3.39)
+$\mathcal{M}_{N}(a,b)\ =\ \mathcal{M}_{N,\textup{ main}}(a,b)+\mathcal{M}_{N,\textup{ error}}(a,b)\ \approx\ b-a.$ (118)
+Thus, based on the estimate of the error term in (101),
+$F_{N}(a,b)\ =\ \mathcal{M}_{N}(a,b)+\mathcal{E}_{N}(a,b)\ \approx\ b-a,$ (119)
 which establishes the equidistribution result.
-3.6. Equidistribution within small interval
-In this section, we show that $d\mathcal{M}_{Y^{(N)}}(y)/dy$ is equidistributed within each small interval $(a+n)/\sqrt{N},(b+n)/\sqrt{N}]$.
-For each $n$, we write the integral over $(a+n)/\sqrt{N},(b+n)/\sqrt{N}]$ in terms of a main term $m_{n}(a,b)$ that is constant with respect to $y$ and an error term $e_{n}(a,b)$, i.e.,
-$\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)dy\ =\ m_{n}(a,b)+e_{n}(a,b),$ (3.40)
+3.6 Equidistribution within small interval
+In this section, we show that $dF_{Y^{(N)},\textup{ main}}(y)/dy$ is equidistributed within each small interval $(a+n)/\sqrt{N},(b+n)/\sqrt{N}]$.
+For each $n$, we write the integral over $(a+n)/\sqrt{N},(b+n)/\sqrt{N}]$ as the sum of a constant main term $m_{n}(a,b)$ and an error term $e_{n}(a,b)$, i.e.,
+$\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y)dy\ =\ m_{n}(a,b)+e_{n}(a,b),$ (120)
 where
-$m_{n}(a,b)$ $\ :=\ \int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}\left(\frac{n}{\sqrt{N}}\right)dy,$
-$e_{n}(a,b)$ $\ :=\ \int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)-\frac{d}{dy}\mathcal{M}_{Y^{(N)}}\left(\frac{n}{\sqrt{N}}\right)dy.$ (3.41)
-Since $d\mathcal{M}_{Y^{(N)}}(n/\sqrt{N})/dy$ is constant, then the main term $m_{n}(a,b)$ is
-$m_{n}(a,b)\ =\ \frac{b-a}{\sqrt{N}}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}\left(\frac{n}{\sqrt{N}}\right).$ (3.42)
+$m_{n}(a,b)$ $\ :=\ \int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}F_{Y^{(N)},\textup{ main}}\left(\frac{n}{\sqrt{N}}\right)dy,$
+$e_{n}(a,b)$ $\ :=\ \int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y)-\frac{d}{dy}F_{Y^{(N)},\textup{ main}}\left(\frac{n}{\sqrt{N}}\right)dy.$ (121)
+Since $dF_{Y^{(N)},\textup{ main}}(n/\sqrt{N})/dy$ is constant, then the main term $m_{n}(a,b)$ is
+$m_{n}(a,b)\ =\ \frac{b-a}{\sqrt{N}}\frac{d}{dy}F_{Y^{(N)},\textup{ main}}\left(\frac{n}{\sqrt{N}}\right).$ (122)
 We now quantify the error term. We start by providing an upper bound on
-$\left|\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)-\frac{d}{dy}\mathcal{M}_{Y^{(N)}}\left(\frac{n}{\sqrt{N}}\right)\right|,$ (3.43)
-for $y\in[n/\sqrt{N},(n+1)/\sqrt{N}]\supset[(a+n)/\sqrt{N},(b+n)/\sqrt{N}]$.
-As we shall see, $d\mathcal{M}_{Y^{(N)}}(y)/dy$ is a continuous function over $n/\sqrt{N},(n+1)/\sqrt{N}]$ and a differentiable function over $(n/\sqrt{N},(n+1)/\sqrt{N})$.
+$\left|\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y)-\frac{d}{dy}F_{Y^{(N)},\textup{ main}}\left(\frac{n}{\sqrt{N}}\right)\right|,$ (123)
+for $y\in[n/\sqrt{N},(n+1)/\sqrt{N}]\supset[(a+n)/\sqrt{N},(b+n)/\sqrt{N}]$. Recall from (3.2) that
+$\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y)$
+$\quad\ =\ \int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ main}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
+$\quad\quad\quad dz_{m-1}\cdots dz_{m+1-d}+O(N^{-2\delta^{\prime}})$
+$\quad\ =\ \ \int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\Phi(z_{m+1-d})\right)^{m-d}\prod_{i=1}^{d-1}\varphi(z_{m+i-d})$
+$\quad\quad\quad\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\cdots dz_{m+1-d}+O(N^{-2\delta^{\prime}}).$ (124)
+We let
+$I_{N}(y)$
+$\quad\ :=\ \int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ main}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
+$\quad\quad\quad dz_{m-1}\cdots dz_{m+1-d}.$ (125)
+We see that $I_{N}(y)$ is a continuous function over $n/\sqrt{N},(n+1)/\sqrt{N}]$ and a differentiable function over $(n/\sqrt{N},(n+1)/\sqrt{N})$.
 By the Mean Value Theorem, for any $y_{1},y_{2}\in[n/\sqrt{N},(n+1)/\sqrt{N}]$ with $y_{1}<y_{2}$, we have
-$\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y_{1})-\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y_{2})\ =\ \frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(c_{n})(y_{1}-y_{2})$ (3.44)
-for some $c_{n}\in(y_{1},y_{2})$. Hence
-$\left|\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y_{1})-\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y_{2})\right|\ \leq\ \left|\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(c_{n})\right||y_{1}-y_{2}|\ \leq\ \frac{1}{\sqrt{N}}\left|\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(c_{n})\right|.$ (3.45)
-We want to provide an upper on $|d^{2}\mathcal{M}_{Y^{(N)}}(y)/dy^{2}|$ over the interval $n/\sqrt{N},(n+1)/\sqrt{N}]$. Recall from (3.2) that
-$\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)$ $\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}m_{Y^{(N)}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}$
-$\quad\quad\cdots dz_{m+1-d}$
-$\ =\ \ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\Phi(z_{m+1-d})\right)^{m-d}\prod_{i=1}^{d-1}\varphi(z_{m+i-d})\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
-$\quad\quad dz_{m-1}\cdots dz_{m+1-d}.$ (3.46)
-Following the same procedure as shown in Section 3.2, we can again exchange the order of differentiation and integration.
+$I_{N}(y_{1})-I_{N}(y_{2})\ =\ \frac{d}{dy}I_{N}(c_{n})(y_{1}-y_{2})$ (126)
+for some $c_{n}\in(y_{1},y_{2})$. Hence,
+$\left|\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y_{1})-\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y_{2})\right|$ $\ \leq\ \left|I_{N}(y_{1})-I_{N}(y_{2})\right|+O(N^{-2\delta^{\prime}})$
+$\ \leq\ \left|\frac{d}{dy}I_{N}(c_{n})\right||y_{1}-y_{2}|+O(N^{-2\delta^{\prime}})$
+$\ \ll\ \frac{1}{\sqrt{N}}\left|\frac{d}{dy}I_{N}(c_{n})\right|+O(N^{-2\delta^{\prime}}).$ (127)
+To obtain an estimate on $dI_{N}(y)/dy$, we differentiate $I_{N}(y)$ by following the same procedure as shown in Section 3.2 to exchange the order of differentiation and integration.
 Since the product in the integrand in (3.6) goes from $j=2$ to d1d-1, we need to discuss the cases when $d=2$ and $d>2$ separately.
 If $d=2$,
-$\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)\ =\ \int_{-C\sqrt{N}}^{\frac{y}{2}}m_{Y^{{(N)}}}\left(z_{m-1},y-z_{m-1}\right)dz_{m-1}.$ (3.47)
-By the Leibniz integral rule and (3.13),
-$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)$
-$\ =\ \frac{1}{2}m_{Y^{(N)}}\left(\frac{y}{2},\frac{y}{2}\right)+\int_{-C\sqrt{N}}^{\frac{y}{2}}\frac{\partial}{\partial y}\left(m_{Y^{(N)}}(z_{m-1},y-z_{m-1})\right)dz_{m-1}$
-$\ \ll\ \varphi\left(\frac{y}{2}\right)+\int_{-C\sqrt{N}}^{\frac{y}{2}}(\Phi(z_{m-1}))^{m-2}\varphi(z_{m-1})\cdot(-1)\cdot(y-z_{m-1})\varphi(y-z_{m-1})dz_{m-1}.$ (3.48)
-We know that $y$ is an upper bound on the value of $z_{m-1}+z_{m}$, then $y\geq-2C\sqrt{N}$, and $y/2\geq-C\sqrt{N}$.
+$I_{N}(y)\ =\ \int_{-N^{\delta^{\prime}}}^{\frac{y}{2}}f_{Z^{(N)}_{(m-1)},Z^{(N)}_{(m)},\textup{ main}}\left(z_{m-1},y-z_{m-1}\right)dz_{m-1}.$ (128)
+By the Leibniz integral rule and (81),
+$\frac{d}{dy}I_{N}(y)$
+$\quad\ =\ \frac{1}{2}f_{Z^{(N)}_{(m-1)},Z^{(N)}_{(m)},\textup{ main}}\left(\frac{y}{2},\frac{y}{2}\right)+\int_{-N^{\delta^{\prime}}}^{\frac{y}{2}}\frac{\partial}{\partial y}\left(f_{Z^{(N)}_{(m-1)},Z^{(N)}_{(m)},\textup{ main}}(z_{m-1},y-z_{m-1})\right)dz_{m-1}$
+$\quad\ \ll\ \varphi\left(\frac{y}{2}\right)+\int_{-N^{\delta^{\prime}}}^{\frac{y}{2}}(\Phi(z_{m-1}))^{m-2}\varphi(z_{m-1})\cdot(-1)\cdot(y-z_{m-1})\varphi(y-z_{m-1})dz_{m-1}.$ (129)
+We know that $y\geq-2N^{\delta^{\prime}}$, so $y/2\geq-N^{\delta^{\prime}}$.
 We now break into cases when $y\leq 0$ and $y\geq 0$.
-When $y\leq 0$, $\varphi(z_{m-1})$ is at most $\varphi(y/2)$ on $C\sqrt{N},y/2]$.
+When $y\leq 0$, $\varphi(z_{m-1})$ is at most $\varphi(y/2)$ on $N^{\delta^{\prime}},y/2]$.
 So (3.6) becomes
-$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)$ $\ \ll\ \varphi\left(\frac{y}{2}\right)+\varphi\left(\frac{y}{2}\right)\int_{-\infty}^{\infty}|y-z_{m-1}|\varphi(y-z_{m-1})dz_{m-1}$
-$\ =\ \varphi\left(\frac{y}{2}\right)+\varphi\left(\frac{y}{2}\right)\int_{-\infty}^{\infty}|z_{m-1}|\varphi(z_{m-1})dz_{m-1}\ \ll\ \varphi\left(\frac{y}{2}\right),$ (3.49)
-where in the last line we use the fact that the integral $\int_{-\infty}^{\infty}|x|\varphi(x)dx$ is the expected value of the absolute value of a standard normal random variable, which is finite.
+$\frac{d}{dy}I_{N}(y)$ $\ \ll\ \varphi\left(\frac{y}{2}\right)+\varphi\left(\frac{y}{2}\right)\int_{-\infty}^{\infty}|y-z_{m-1}|\varphi(y-z_{m-1})dz_{m-1}$
+$\ =\ \varphi\left(\frac{y}{2}\right)+\varphi\left(\frac{y}{2}\right)\int_{-\infty}^{\infty}|z_{m-1}|\varphi(z_{m-1})dz_{m-1}\ \ll\ \varphi\left(\frac{y}{2}\right),$ (130)
+where in the last line we use the fact that the integral $\int_{-\infty}^{\infty}|x|\varphi(x)dx$ is the expected value of the absolute value of a standard Gaussian random variable, which is finite.
 When $y\geq 0$, we have
-$y-z_{m-1}$ $\ \geq\ y-\frac{y}{2}\ =\ \frac{y}{2}\ \geq\ 0.$ (3.50)
+$y-z_{m-1}$ $\ \geq\ y-\frac{y}{2}\ =\ \frac{y}{2}\ \geq\ 0.$ (131)
 Hence, (3.6) becomes
-$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)$ $\ \ll\ \varphi\left(\frac{y}{2}\right)+\int_{-C\sqrt{N}}^{\frac{y}{2}}(y-z_{m-1})\varphi(y-z_{m-1})dz_{m-1}$
-$\ =\ \varphi\left(\frac{y}{2}\right)+\varphi(y-z_{m-1})\Bigg{|}_{z_{m-1}=-C\sqrt{N}}^{z_{m-1}=\frac{y}{2}}\ \ll\ \varphi\left(\frac{y}{2}\right)+\varphi\left(y+C\sqrt{N}\right),$ (3.51)
+$\frac{d}{dy}I_{N}(y)$ $\ \ll\ \varphi\left(\frac{y}{2}\right)+\int_{-N^{\delta^{\prime}}}^{\frac{y}{2}}(y-z_{m-1})\varphi(y-z_{m-1})dz_{m-1}$
+$\ =\ \varphi\left(\frac{y}{2}\right)+\varphi(y-z_{m-1})\Bigg|_{z_{m-1}=-N^{\delta^{\prime}}}^{z_{m-1}=\frac{y}{2}}\ \ll\ \varphi\left(\frac{y}{2}\right)+\varphi\left(y+N^{\delta^{\prime}}\right),$ (132)
 where in the second line we employ the fact that $\int(y-x)\varphi(y-x)dx=\varphi(y-x)$.
-Thus, if $d=2$, regardless of whether $y\leq 0$ or $y\geq 0$, $d^{2}\mathcal{M}_{Y^{(N)}}(y)/dy^{2}$ is on the order of $\varphi(y/2)+\varphi(y+C\sqrt{N})$.
+Thus, if $d=2$, regardless of whether $y\leq 0$ or $y\geq 0$, $d^{2}\mathcal{M}_{Y^{(N)}}(y)/dy^{2}$ is on the order of $\varphi(y/2)+\varphi(y+N^{\delta^{\prime}})$.
 This concludes the estimate of $d^{2}\mathcal{M}_{Y^{(N)}}(y)/dy^{2}$ for $d=2$.
-If $d>2$, we repeatedly apply Leibniz integral rule as we have demonstrated in Subsection 3.4 to obtain
-$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)$
-$\quad\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-2}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\frac{\partial}{\partial y}\Bigg{(}\int_{z_{m-2}}^{\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}}m_{Y^{(N)}}\Bigg{(}z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\Bigg{)}$
-$\quad\quad\quad\quad dz_{m-1}\Bigg{)}dz_{m-2}\cdots dz_{m+1-d}$
-$\quad\ \ll\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-2}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\Phi(z_{m+1-d})\right)^{m-d}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\right)^{2}$
+If $d>2$, we repeatedly apply Leibniz integral rule as we have demonstrated in Subsection 3.4 to (3.6) and also use (81) to obtain
+$\frac{d}{dy}I_{N}(y)$
+$\ =\ \int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-2}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\frac{\partial}{\partial y}\Bigg(\int_{z_{m-2}}^{\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ main}}\Bigg(z_{m+1-d},\dots,z_{m-1},$
+$\quad\quad y-\sum_{i=1}^{d-1}z_{m+i-d}\Bigg)dz_{m-1}\Bigg)dz_{m-2}\cdots dz_{m+1-d}$
+$\ \ll\ \int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-2}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\Phi(z_{m+1-d})\right)^{m-d}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\right)^{2}$
 $\quad\quad+\Bigg(\int_{z_{m-2}}^{\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}}(\Phi(z_{m+1-d}))^{m-d}\left(\prod_{i=1}^{d-1}\varphi(z_{m+i-d})\right)(-1)\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
-$\quad\quad\quad\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\Bigg{)}dz_{m-2}\cdots dz_{m+1-d}.$ (3.52)
+$\quad\quad\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\Bigg)dz_{m-2}\cdots dz_{m+1-d}.$ (133)
 Now, similar to the case when $d=2$, we estimate the second integrand in (3.6) by breaking down into cases when $(y-\sum_{i=1}^{d-2}z_{m+i-d})/2\leq 0$ and $(y-\sum_{i=1}^{d-2}z_{m+i-d})\geq 0$.
-We leave the details to Appendix C.
-Hence, we have $d^{2}\mathcal{M}_{Y^{(N)}}(y)/dy^{2}$ in (3.6) is bounded above by
+We leave the details to C.
+Hence, we have $dI_{N}(y)/dy$ in (3.6) is bounded above by
 $\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)dz_{m-2}\cdots dz_{m+1-d}$
-$\quad\quad+\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)dz_{m-2}\cdots dz_{m+1-d}.$ (3.53)
-We first obtain some estimate on the first integral of (3.6).
-The idea behind estimating this integral is recursive uses of change of variables that lead to consistent cancellation.
-We first do the change of variable $z_{m-2}\mapsto z_{m-2}+(y-\sum_{i=1}^{d-3}z_{m+i-d})/5$.
-We have
-$\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)dz_{m-2}\cdots dz_{m+1-d}$
-$\quad\ \ll\ \underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}e^{-\frac{\sum_{i=1}^{d-3}z^{2}_{m+i-d}}{2}}\cdot e^{-\frac{\left(z_{m-2}+\left(y-\sum_{i=1}^{d-3}z_{m+i-d}\right)/5\right)^{2}}{2}}$
-$\quad\quad\quad\cdot e^{-\frac{\frac{1}{4}\left(-z_{m-2}+4\left(y-\sum_{i=1}^{d-3}z_{m+i-d}\right)/5\right)^{2}}{2}}dz_{m-2}\cdots dz_{m+1-d}.$ (3.54)
-Notice that in the expansion of the exponents, we get a cancellation of the term $2z_{m-2}((y-\sum_{i=1}^{d-3}z_{m+i-d})/5)$. Hence, (3.6) is bounded above by
-$\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-3}e^{-\frac{\sum_{i=1}^{d-3}z_{m+i-d}^{2}}{2}}\cdot e^{-\frac{\frac{1}{5}\left(y-\sum_{i=1}^{d-3}z_{m+i-d}\right)^{2}}{2}}\left(\int_{-\infty}^{\infty}e^{-\frac{\frac{5}{4}z_{m-2}^{2}}{2}}dz_{m-2}\right)dz_{m-3}\cdots dz_{m+1-d}$
-$\quad\ \ll\ \underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-3}e^{-\frac{\sum_{i=1}^{d-3}z_{m+i-d}^{2}}{2}}\cdot e^{-\frac{\frac{1}{5}\left(y-\sum_{i=1}^{d-3}z_{m+i-d}\right)^{2}}{2}}dz_{m-3}\cdots dz_{m+1-d}$ (3.55)
-We see that the estimate in (3.6) has a similar structure, which allows us to use induction to estimate the interval. For an arbitrary $k$ such that $1\leq k\leq d-2$, we want to look at
-$\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{k}e^{-\frac{\sum_{i=1}^{k}z^{2}_{m+i-d}}{2}}\cdot e^{-\frac{A_{k}(y-\sum_{i=1}^{k}z_{m+i-d})^{2}}{2}}dz_{m+k-d}\cdots dz_{m+1-d},$ (3.56)
-where $A_{k}>0$. Using the change of variable $z_{m+k-d}\mapsto z_{m+k-d}+\frac{A_{k}}{1+A_{k}}(y-\sum_{i=1}^{k-1}z_{m+i-d})$, (3.56) is bounded above by
-$\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{k}e^{-\frac{\sum_{i=1}^{k-1}z^{2}_{m+i-d}}{2}}\cdot e^{-\frac{\left(z_{m+k-d}+\frac{A_{k}}{1+A_{k}}\left(y-\sum_{i=1}^{k-1}z_{m+i-d}\right)\right)^{2}}{2}}$
-$\quad\quad\quad\cdot e^{-\frac{A_{k}\left(-z_{m+k-d}+\frac{1}{1+A_{k}}\left(y-\sum_{i=1}^{k-1}z_{m+i-d}\right)\right)^{2}}{2}}dz_{m+k-d}\cdots dz_{m+1-d}.$ (3.57)
-Expanding the exponents also gives a cancellation of the term
-$\frac{2A_{k}}{1+A_{k}}z_{m+k-d}\left(y-\sum_{i=1}^{k-1}z_{m+i-d}\right).$ (3.58)
-Hence, (3.56) is bounded above by
-$\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{k}e^{-\frac{\sum_{i=1}^{k-1}z^{2}_{m+i-d}}{2}}\cdot e^{-\frac{A_{k-1}\left(y-\sum_{i=1}^{k-1}z_{m+i-d}\right)^{2}}{2}}e^{-\frac{(1+A_{k})z^{2}_{m+k-d}}{2}}dz_{m+k-d}\cdots dz_{m+1-d},$ (3.59)
-where
-$A_{k-1}\ :=\ \frac{A_{k}^{2}}{1+A_{k}^{2}}+\frac{A_{k}}{(1+A_{k})^{2}}.$ (3.60)
-Note that since $A_{k}>0$, then $A_{k+1}>0$. Hence, (3.56) becomes bounded above by
-$\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{k-1}e^{-\frac{\sum_{i=1}^{k-1}z^{2}_{m+i-d}}{2}}\cdot e^{-\frac{A_{k-1}\left(y-\sum_{i=1}^{k-1}z_{m+i-d}\right)^{2}}{2}}\left(\int_{-\infty}^{\infty}e^{-\frac{(1+A_{k})z_{m+k-d}^{2}}{2}}dz_{m+k-d}\right)$
-$\quad\quad\quad dz_{m+(k-1)-d}\cdots dz_{m+1-d}$
-$\quad\ \ll\ \underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{k-1}e^{-\frac{\sum_{i=1}^{k-1}z^{2}_{m+i-d}}{2}}\cdot e^{-\frac{A_{k-1}\left(y-\sum_{i=1}^{k-1}z_{m+i-d}\right)^{2}}{2}}dz_{m+(k-1)-d}\cdots dz_{m+1-d},$ (3.61)
-Thus, by induction, (3.6) becomes bounded above by
-$e^{-(A_{0}y^{2})/2},$ (3.62)
-where $A_{0}>0$ is recursively defined by (3.60) with $A_{d-2}=1/4$. Similarly, we can show by induction that the second integral in (3.6) has the estimate
-$\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)\right)dz_{m-2}\cdots dz_{m+1-d}\ \ll\ e^{-(B_{0}y^{2})/2}$ (3.63)
-for some $B_{0}>0$. Hence, if $d>2$, (3.6) becomes
-$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)$ $\ \ll\ e^{-(A_{0}y^{2})/2}+e^{-(B_{0}y^{2})/2}\ \ll\ e^{-(C_{0}y^{2})/2},$ (3.64)
-where $C_{0}:=\min\{A_{0},B_{0}\}$. Recall from (3.6) that when $d=2$,
-$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)\ \ll\ \varphi\left(\frac{y}{2}\right)+\varphi\left(y+C\sqrt{N}\right).$ (3.65)
+$\quad\quad+\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)dz_{m-2}\cdots dz_{m+1-d}.$ (134)
+Notice that the first integral in (3.6) is the PDF of $W_{1}+\cdots+W_{d-2}+2W_{d-1}$, where $W_{i}$’s are i.i.d.
+standard Gaussians, and the second integral in (3.6) is the PDF of $V_{1}+\cdots+V_{d-2}+3/2V_{d-1}$, where the $V_{i}$’s are also i.i.d.
+standard Gaussians.
+Hence, there exists a constant $C_{0}>0$ such that
+$\frac{d}{dy}I_{N}(y)\ \ll\ e^{-(C_{0}y^{2})/2}.$ (135)
+Moreover, recall from (3.6) that when $d=2$,
+$\frac{d}{dy}I_{N}(y)\ \ll\ \varphi\left(\frac{y}{2}\right)+\varphi\left(y+N^{\delta^{\prime}}\right).$ (136)
 Thus, regardless of whether $d=2$ or $d>2$,
-$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)\ \ll\ e^{-\frac{D_{0}y^{2}}{2}}+e^{-\frac{D_{0}\left(y+C\sqrt{N}\right)^{2}}{2}},$ (3.66)
+$\frac{d}{dy}I_{N}(y)\ \ll\ e^{-\frac{D_{0}y^{2}}{2}}+e^{-\frac{D_{0}\left(y+N^{\delta^{\prime}}\right)^{2}}{2}},$ (137)
 where $D_{0}:=\min\{C_{0},1/4\}$.
-Note that $e^{-D_{0}y^{2}/2}$ and $\varphi(\sqrt{D_{0}}(y+C\sqrt{N}))$ each has only one global extreme, at $y=0$ and $y=-C\sqrt{N}$ respectively.
+Note that $e^{-D_{0}y^{2}/2}$ and $\varphi(\sqrt{D_{0}}(y+N^{\delta^{\prime}}))$ each has only one global extreme, at $y=0$ and $y=-N^{\delta^{\prime}}$ respectively.
 Hence, locally on $n/\sqrt{N},(n+1)/\sqrt{N}]$ the functions are monotonic and can be bounded above by the sum of its values at the two end points.
 Hence, for all $n$,
-$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)\ \ll\ e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}+C\sqrt{N}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}+C\sqrt{N}\right)^{2}}{2}},$ (3.67)
-for all $y\in[n/\sqrt{N},(n+1)/\sqrt{N}]$. Thus, returning to (3.45), for all $y_{1},y_{2}\in[n/\sqrt{N},(n+1)/\sqrt{N}]$,
-$\left|\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y_{1})-\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y_{2})\right|$
-$\quad\ \ll\ \frac{1}{\sqrt{N}}\left(e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}+C\sqrt{N}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}+C\sqrt{N}\right)^{2}}{2}}\right).$ (3.68)
+$\frac{d}{dy}I_{N}(y)\ \ll\ e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}+N^{\delta^{\prime}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}+N^{\delta^{\prime}}\right)^{2}}{2}},$ (138)
+for all $y\in[n/\sqrt{N},(n+1)/\sqrt{N}]$. Now, we return to (3.6). For all $y_{1},y_{2}\in[n/\sqrt{N},(n+1)/\sqrt{N}]$,
+$\left|\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y_{1})-\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y_{2})\right|$
+$\quad\ \ll\ \frac{1}{\sqrt{N}}\left(e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}+N^{\delta^{\prime}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}+N^{\delta^{\prime}}\right)^{2}}{2}}\right)+O(N^{-2\delta^{\prime}}).$ (139)
 Thus, the error term $e_{n}(a,b)$ as defined in (3.6) is bounded by
-$e_{n}(a,b)\ \leq\ \int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\left|\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)-\frac{d}{dy}\mathcal{M}_{Y^{(N)}}\left(\frac{n}{\sqrt{N}}\right)\right|dy$
-$\quad\ \ll\ \left(\frac{1}{\sqrt{N}}\right)^{2}\left(e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}+C\sqrt{N}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}+C\sqrt{N}\right)^{2}}{2}}\right).$ (3.69)
-3.7. Upper bound on error term $\mathcal{M}_{\textup{err},N}(a,b)$
+$e_{n}(a,b)$ $\ \leq\ \int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\left|\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y)-\frac{d}{dy}F_{Y^{(N)},\textup{ main}}\left(\frac{n}{\sqrt{N}}\right)\right|dy$
+$\ \ll\ \left(\frac{1}{\sqrt{N}}\right)^{2}\left(e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}+N^{\delta^{\prime}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}+N^{\delta^{\prime}}\right)^{2}}{2}}\right)$
+$\quad\quad+O(N^{-1/2-2\delta^{\prime}}).$ (140)
+3.7 Upper bound on error term $\mathcal{M}_{N,\textup{ error}}(a,b)$
 In this section, we want to prove that the error term of the main term $\mathcal{M}_{N}(a,b)$ is small, i.e.,
-$\mathcal{M}_{\textup{err,N}}(a,b)\ =\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}e_{n}(a,b)\ \approx\ 0.$ (3.70)
-Based on the estimate $e_{n}(a,b)$ in (3.6), we can pull out one of the $1/\sqrt{N}$ factors in the bracket and obtain
-$\mathcal{M}_{\textup{err},N}(a,b)\ =\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}e_{n}(a,b)$
-$\ \ll\ \frac{1}{\sqrt{N}}\cdot\left(\frac{1}{\sqrt{N}}\sum_{n=-dC\cdot N}^{dC\cdot N-1}\left(e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}+C\sqrt{N}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}+C\sqrt{N}\right)^{2}}{2}}\right)\right).$ (3.71)
-We observe that the above in the parentheses are Riemann sums for $e^{-D_{0}x^{2}/2}$ and $e^{-D_{0}(x+C\sqrt{N})^{2}/2}$ on $(-\infty,\infty)$, which are both finite. Hence
-$\mathcal{M}_{\textup{err},N}(a,b)\ \ll\ \frac{1}{\sqrt{N}}.$ (3.72)
+$\mathcal{M}_{N,\textup{ error}}(a,b)\ =\ \sum_{n=-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}e_{n}(a,b)\ \approx\ 0.$ (141)
+Based on the estimate of $e_{n}(a,b)$ in (3.6), we can pull out one of the $1/\sqrt{N}$ factors and obtain
+$\mathcal{M}_{N,\textup{ error}}(a,b)\ =\ \sum_{n=-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}e_{n}(a,b)$
+$\ \ll\ \frac{1}{\sqrt{N}}\cdot\left(\frac{1}{\sqrt{N}}\sum_{n=-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}\left(e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n}{\sqrt{N}}+N^{\delta^{\prime}}\right)^{2}}{2}}+e^{-\frac{D_{0}\left(\frac{n+1}{\sqrt{N}}+N^{\delta^{\prime}}\right)^{2}}{2}}\right)\right)$
+$\quad\quad+O(N^{-\delta^{\prime}}).$ (142)
+We observe that item in the parentheses above is double of the sum of the Riemann sums for $\int_{-\infty}^{\infty}e^{-D_{0}x^{2}/2}dx$ and $\int_{-\infty}^{\infty}e^{-D_{0}(x+N^{\delta^{\prime}})^{2}/2}dx$, which are both finite.
+Hence,
+[…]
 We have thus established that the error term $\mathcal{M}_{\textup{err},N}(a,b)$ of the main term $\mathcal{M}_{N}(a,b)$ is negligible.
-3.8. Evaluation of main term $\mathcal{M}_{\textup{main},N}(a,b)$
-Finally, we establish the main term $\mathcal{M}_{\textup{main},N}(a,b)$ of the main term $\mathcal{M}_{N}(a,b)$. From (3.37) and (3.42), we see that
-$\mathcal{M}_{\textup{main},N}(a,b)$ $\ =\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}m_{n}(a,b)$
-$\ =\ (b-a)\cdot\left(\frac{1}{\sqrt{N}}\sum_{n=-dC\cdot N}^{dC\cdot N-1}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}\left(\frac{n}{\sqrt{N}}\right)\right).$ (3.73)
-Since the term in the parentheses above is a Riemann sum for $\int_{-\infty}^{\infty}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)dy$, we have
-$\frac{1}{\sqrt{N}}\sum_{n=-2C\cdot N}^{2C\cdot N-1}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}\left(\frac{n}{\sqrt{N}}\right)$ $\ =\ \int_{-\infty}^{\infty}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)dy+o(1)$
-$\ =\ \mathcal{M}_{Y^{(N)}}(\infty)-\mathcal{M}_{Y^{(N)}}(-\infty)+o(1).$ (3.74)
-Hence the main term $\mathcal{M}_{\textup{main},N}(a,b)$ becomes
-$\mathcal{M}_{\textup{main},N}(a,b)\ =\ (b-a)(\mathcal{M}_{Y^{(N)}}(\infty)-\mathcal{M}_{Y^{(N)}}(-\infty))+o(1).$ (3.75)
+3.8 Evaluation of main term $\mathcal{M}_{N,\textup{ main}}(a,b)$
+Finally, we establish the main term $\mathcal{M}_{N,\textup{ main}}(a,b)$ of $\mathcal{M}_{N}(a,b)$. From (117) and (122), we see that
+$\mathcal{M}_{N,\textup{ main}}(a,b)$ $\ =\ \sum_{n=-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}m_{n}(a,b)$
+$\ =\ (b-a)\cdot\left(\frac{1}{\sqrt{N}}\sum_{n=-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}\frac{d}{dy}F_{Y^{(N)},\textup{ main}}\left(\frac{n}{\sqrt{N}}\right)\right).$ (144)
+Since the item in the parentheses above is the Riemann sum for $\int_{-\infty}^{\infty}\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y)dy$, we have
+$\frac{1}{\sqrt{N}}\sum_{n=-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}\frac{d}{dy}F_{Y^{(N)},\textup{ main}}\left(\frac{n}{\sqrt{N}}\right)$ $\ =\ \int_{-\infty}^{\infty}\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y)dy+o(1)$
+$\ =\ F_{Y^{(N)},\textup{ main}}(\infty)-F_{Y^{(N)},\textup{ main}}(-\infty)+o(1).$ (145)
+Hence the main term $\mathcal{M}_{N,\textup{ main}}(a,b)$ becomes
+$\mathcal{M}_{N,\textup{ main}}(a,b)\ =\ (b-a)(F_{Y^{(N)},\textup{ main}}(\infty)-F_{Y^{(N),\textup{ main}}}(-\infty))+o(1).$ (146)
 Recall that
-$\mathcal{M}_{Y^{(N)}}(y)$ $\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}m_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}.$ (3.76)
-To simplify our calculation, we want to extend the interval of integration of the outer integral to -$\infty$. We do so by showing that the tail of the integral is small, i.e.,
-$T_{N}\ :=\ \int_{-\infty}^{-C\sqrt{N}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}m_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}\ \approx\ 0.$ (3.77)
-We have
-$T_{N}$ $\ \ll\ \int_{-\infty}^{-C\sqrt{N}}\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-1}(\Phi(z_{m+1-d}))^{m-d}\prod_{i=1}^{d}\varphi(z_{m+i-d})dz_{m}\cdots dz_{m+1-d}$
-$\ \ll\ \int_{-\infty}^{-C\sqrt{N}}\varphi(z_{m+1-d})dz_{m+1-d},$ (3.78)
-Hence, we have reduced the problem to showing the smallness of the Gaussian tail. We first make a definition.
-Definition 3.5.
-For functions $f(x)$ and $g(x)$, we say that $f(x)=\Theta(g(x))$ if $f(x)=O(g(x))$ and $g(x)=O(f(x))$.
-The following result is a straightforward calculation that provides an upper bound on the Gaussian tail.
-Proposition 3.6.
-[Mil1] Suppose that $g:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ is a function such that $g(N)=\Theta(N^{\epsilon^{\prime}})$ for some $\epsilon^{\prime}>0$. Then
-$\int_{-\infty}^{-g(N)}\frac{1}{\sqrt{2\pi}}e^{-x^{2}/2}dx\ =\ \int_{g(N)}^{\infty}\frac{1}{\sqrt{2\pi}}e^{-x^{2}/2}dx\ \ll\ e^{-g(N)/2}.$ (3.79)
-Back to (3.8), by Proposition 3.6
-$T_{N}\ \ll\ e^{-C\sqrt{N}}.$ (3.80)
-Thus, (3.76) becomes
-$\mathcal{M}_{Y^{(N)}}(y)$ $\ =\ \int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}m_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}+O\left(e^{-C\sqrt{N}}\right).$ (3.81)
-Now, by Proposition 3.3,
-$m_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})\ :=\ C_{m}^{m+1-d}\left(\Phi(z_{m+1-d})\right)^{m-d}\prod_{i=1}^{d}\varphi(z_{m+i-d})$ (3.82)
-is the joint PDF of $W_{(m+1-d)},\dots,W_{(m)}$, where $W_{1},\dots,W_{m}$ are i.i.d. $\sim N(0,1)$. Thus, following our derivation in Subsection 3.2,
-$H(y)\ :=\ \int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}m_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}$ (3.83)
-is the CDF of $\sum_{i=1}^{d}W_{(m+i-d)}$. We know that for a cumulative density function $H(y)$, $H(-\infty)=0$ and $H(\infty)=1$. Hence,
-$\mathcal{M}_{Y^{(N)}}(-\infty)$ $\ =\ H(-\infty)+O(e^{-C\sqrt{N}})\ =\ O\left(e^{-C\sqrt{N}}\right)$
-$\mathcal{M}_{Y^{(N)}}(\infty)$ $\ =\ H(\infty)+O(e^{-C\sqrt{N}})\ =\ 1+O\left(e^{-C\sqrt{N}}\right).$ (3.84)
-Thus, substituting these two estimates into (3.75) yields
-$\mathcal{M}_{\textup{main},N}(a,b)\ =\ (b-a)\left(1+O\left(e^{-C\sqrt{N}}\right)\right)+o(1))\ =\ (b-a)+o(1).$ (3.85)
-Returning to (3.39) and (3.38), combined with the estimate for the error term $\mathcal{E}_{N}(a,b)$ in (3.4) and the error term $\mathcal{M}_{\textup{err},N}(a,b)$ of $\mathcal{M}_{N}(a,b)$ in (3.72), we have
+$F_{Y^{(N)},\textup{ main}}(y)$ $\ =\ \int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ main}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}.$ (147)
+Now, by Proposition 5 and (81),
+$f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ main}}(z_{m+1-d},\dots,z_{m})\ :=\ C_{m}^{m+1-d}\left(\Phi(z_{m+1-d})\right)^{m-d}\prod_{i=1}^{d}\varphi(z_{m+i-d})$ (148)
+is the joint PDF of $W_{(m+1-d)},\dots,W_{(m)}$, where $W_{1},\dots,W_{m}$ are i.i.d. standard Gaussians. Thus, following our derivation in Subsection 3.2,
+$F_{Y^{(N)},\textup{ main}}(y)\ =\ \int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ main}}(z_{m+1-d},\dots,z_{m})dz_{m}\cdots dz_{m+1-d}$ (149)
+is the CDF of $\sum_{i=1}^{d}W_{(m+i-d)}$. We know that for a cumulative density function $F(y)$, $F(-\infty)=0$ and $F(\infty)=1$. Thus, substituting these two estimates into (146) yields
+$\mathcal{M}_{N,\textup{ main}}(a,b)\ =\ (b-a)+o(1).$ (150)
+Returning to (119) and (118), combined with the estimate for the error term $\mathcal{E}_{N}(a,b)$ in (3.4) and the error term $\mathcal{M}_{\textup{err},N}(a,b)$ of $\mathcal{M}_{N}(a,b)$ in (143), we have
 $F_{N}(a,b)$ $\ =\ \mathcal{M}_{N}(a,b)+\mathcal{E}_{N}(a,b)$
-$\ =\ \mathcal{M}_{\textup{main},N}(a,b)+\mathcal{M}_{\textup{err},N}(a,b)+\mathcal{E}_{N}(a,b)$
-$\ =\ (b-a)+o(1)+O\left(\frac{1}{\sqrt{N}}\right)+O(N^{-\delta})$
-$\ =\ (b-a)+o(1).$ (3.86)
+[…]
+$\ =\ (b-a)+o(1)+O(N^{-\delta^{\prime}})+O(N^{-\delta^{\prime}})$
+$\ =\ (b-a)+o(1).$ (151)
 We conclude that $\log_{B}(\mathfrak{m}_{d}^{(N)})$ converges to being equidistributed mod 1, and therefore by Uniform Distribution Characterization $\mathfrak{m}_{d}^{(N)}$ converges to strong Benford behavior.
 ∎
-Appendix A Proof of Proposition 2.5
-In this section, we prove Proposition 2.5 which provides a quantitative bound on the difference among probabilities within an interval. Let us first recall the statement of Proposition 2.5.
-Proposition 2.5.
+4 Funding
+This work was partially supported by Williams College Summer Science Program Research Fellowship, the Finnerty Fund, and NSF Grant DMS2241623.
+We would like to thank the anonymous referee for their valuable time and constructive comments.
+Appendix A Proof of Proposition 3
+In this section, we prove Proposition 3, which provides a quantitative bound on the difference among probabilities within an interval. Let us first recall the statement of Proposition 3.
+Proposition 3.
 For $\ell\leq\sqrt{k(N)}/2$,
-$\left|\binom{k(N)}{k_{1,\ell}}-\binom{k(N)}{k_{1,\ell+1}}\right|\ \leq\ O\left(\binom{k(N)}{k_{1,\ell}}\cdot N^{-\frac{3\epsilon}{10}}\right).$ (A.1)
+$\left|\binom{k(N)}{k_{1,\ell}}-\binom{k(N)}{k_{1,\ell+1}}\right|\ \leq\ O\left(\binom{k(N)}{k_{1,\ell}}\cdot N^{-\frac{3\epsilon}{10}}\right).$ (152)
 Proof.
@@
 $\ =\ \frac{k(N)!}{\left(\frac{k(N)}{2}+\ell N^{\delta}\right)!\left(\frac{k(N)}{2}-\ell N^{\delta}\right)!}-\frac{k(N)!}{\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)!\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)}$
-$\ =\ \frac{k(N)!\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)!\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!-k(N)!\left(\frac{k(N)}{2}+\ell N^{\delta}\right)!\left(\frac{k(N)}{2}-\ell N^{\delta}\right)!}{\left(\frac{k(N)}{2}+\ell N^{\delta}\right)!\left(\frac{k(N)}{2}-\ell N^{\delta}\right)!\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)!\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!}$
 $\ =\ \frac{k(N)!}{\left(\frac{k(N)}{2}+\ell N^{\delta}\right)!\left(\frac{k(N)}{2}-\ell N^{\delta}\right)!}$
-$\cdot\frac{\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!-\left(\frac{k(N)}{2}+\ell N^{\delta}\right)!\left(\frac{k(N)}{2}-\ell N^{\delta}\right)!}{\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)!\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!}$
-$\ =\ \binom{k(N)}{k_{1,\ell}}\left(1-\frac{\left(\frac{k(N)}{2}+\ell N^{\delta}\right)!\left(\frac{k(N)}{2}-\ell N^{\delta}\right)!}{\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)!\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!}\right).$ (A.2)
+$\cdot\frac{\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)!\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!-\left(\frac{k(N)}{2}+\ell N^{\delta}\right)!\left(\frac{k(N)}{2}-\ell N^{\delta}\right)!}{\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)!\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!}$
+$\ =\ \binom{k(N)}{k_{1,\ell}}\left(1-\frac{\left(\frac{k(N)}{2}+\ell N^{\delta}\right)!\left(\frac{k(N)}{2}-\ell N^{\delta}\right)!}{\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)!\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!}\right).$ (153)
 Now, we analyze the term
-$\alpha_{\ell,N}\ :=\ \frac{\left(\frac{k(N)}{2}+\ell N^{\delta}\right)!\left(\frac{k(N)}{2}-\ell N^{\delta}\right)!}{\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)!\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!}.$ (A.3)
+$\alpha_{\ell,N}\ :=\ \frac{\left(\frac{k(N)}{2}+\ell N^{\delta}\right)!\left(\frac{k(N)}{2}-\ell N^{\delta}\right)!}{\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)!\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)!}.$ (154)
 We want to show that $\alpha_{\ell,N}\rightarrow 1$ as $N\rightarrow\infty$, so that the difference in (A) is asymptotically much smaller than the main term $\binom{k(N)}{k_{1,\ell}}$. We have
 $\frac{\left(\frac{k(N)}{2}-(\ell+1)N^{\delta}\right)^{N^{\delta}}}{\left(\frac{k(N)}{2}+(\ell+1)N^{\delta}\right)^{N^{\delta}}}$ $\ \leq\ \alpha_{\ell,N}\ \leq\ \frac{\left(\frac{k(N)}{2}-\ell N^{\delta}\right)^{N^{\delta}}}{\left(\frac{k(N)}{2}+\ell N^{\delta}\right)^{N^{\delta}}}$
-$\left(1-\frac{4(\ell+1)N^{\delta}}{k(N)+2(\ell+1)N^{\delta}}\right)^{N^{\delta}}$ $\ \leq\ \alpha_{\ell,N}\ \leq\left(1-\frac{4\ell N^{\delta}}{k(N)+2\ell N^{\delta}}\right)^{N^{\delta}}.$ (A.4)
+$\left(1-\frac{4(\ell+1)N^{\delta}}{k(N)+2(\ell+1)N^{\delta}}\right)^{N^{\delta}}$ $\ \leq\ \alpha_{\ell,N}\ \leq\left(1-\frac{4\ell N^{\delta}}{k(N)+2\ell N^{\delta}}\right)^{N^{\delta}}.$ (155)
 Since $\ell\leq\sqrt{k(N)}/2$, then
-$0\ \leq\ \frac{4\ell N^{\delta}}{k(N)+2\ell N^{\delta}}\ \leq\ \frac{4\ell N^{\delta}}{k(N)}\ \leq\ \frac{2N^{\delta}}{\sqrt{k(N)}}.$ (A.5)
+$0\ \leq\ \frac{4\ell N^{\delta}}{k(N)+2\ell N^{\delta}}\ \leq\ \frac{4\ell N^{\delta}}{k(N)}\ \leq\ \frac{2N^{\delta}}{\sqrt{k(N)}}.$ (156)
 Since $k(N)\geq N^{\epsilon}$, and $\delta\in(0,\epsilon/10)$, then
-$0\ \leq\ \frac{2N^{2\delta}}{\sqrt{k(N)}}$ $\ \leq\ \frac{2N^{\epsilon/10}}{N^{\epsilon/2}}\ =\ 2N^{-2\epsilon/5}\ =\ O(N^{-2\epsilon/5}).$ (A.6)
+$0\ \leq\ \frac{2N^{2\delta}}{\sqrt{k(N)}}$ $\ \leq\ \frac{2N^{\epsilon/10}}{N^{\epsilon/2}}\ =\ 2N^{-2\epsilon/5}\ =\ O(N^{-2\epsilon/5}).$ (157)
 Similarly, we also have
-$\frac{4(\ell+1)N^{\delta}}{k(N)+2(\ell+1)N^{\delta}}\ =\ O(N^{-2\epsilon/5}).$ (A.7)
+$\frac{4(\ell+1)N^{\delta}}{k(N)+2(\ell+1)N^{\delta}}\ =\ O(N^{-2\epsilon/5}).$ (158)
 Hence, for sufficiently large $N$,
@@
 Returning to (A), given (A), we have
-$\left(1-\frac{4(\ell+1)N^{\delta}}{k(N)+2(\ell+1)N^{\delta}}\right)^{N^{\delta}}\ \leq\ \alpha_{\ell,N}\ \leq\ \left(1-\frac{4\ell N^{\delta}}{k(N)+2\ell N^{\delta}}\right)^{N^{\delta}}.$ (A.9)
+$\left(1-\frac{4(\ell+1)N^{\delta}}{k(N)+2(\ell+1)N^{\delta}}\right)^{N^{\delta}}\ \leq\ \alpha_{\ell,N}\ \leq\ \left(1-\frac{4\ell N^{\delta}}{k(N)+2\ell N^{\delta}}\right)^{N^{\delta}}.$ (160)
 Using binomial expansion,
-$\sum_{j=1}^{N^{\delta}}\binom{N^{\delta}}{j}(-1)^{j}\left(\frac{4(\ell+1)N^{\delta}}{k(N)+2(\ell+1)N^{\delta}}\right)^{j}$ $\ \leq\ \alpha_{\ell,N}-1\ \leq\ \sum_{j=1}^{N^{\delta}}\binom{N^{\delta}}{j}(-1)^{j}\left(\frac{4\ell N^{\delta}}{k(N)+2\ell N^{\delta}}\right)^{j}.$ (A.10)
+$\sum_{j=1}^{N^{\delta}}\binom{N^{\delta}}{j}(-1)^{j}\left(\frac{4(\ell+1)N^{\delta}}{k(N)+2(\ell+1)N^{\delta}}\right)^{j}$ $\ \leq\ \alpha_{\ell,N}-1\ \leq\ \sum_{j=1}^{N^{\delta}}\binom{N^{\delta}}{j}(-1)^{j}\left(\frac{4\ell N^{\delta}}{k(N)+2\ell N^{\delta}}\right)^{j}.$ (161)
 We first bound the right sum in (A). Using the assumption that $\ell\leq\sqrt{k(N)}/2$, we have
@@
 $\ =\ \sum_{j=1}^{N^{\delta}}\left(\frac{2N^{2\delta}}{\sqrt{k(N)}}\right)^{j}$
-$\ =\ \frac{2N^{2\delta}}{\sqrt{k(N)}}\cdot\frac{1-\left(\frac{2N^{2\delta}}{\sqrt{k(N)}}\right)^{N^{\delta}}}{1-\frac{2N^{2\delta}}{\sqrt{k(N)}}},$ (A.11)
+$\ =\ \frac{2N^{2\delta}}{\sqrt{k(N)}}\cdot\frac{1-\left(\frac{2N^{2\delta}}{\sqrt{k(N)}}\right)^{N^{\delta}}}{1-\frac{2N^{2\delta}}{\sqrt{k(N)}}},$ (162)
 where on the last line we use the geometric series formula. Since $k(N)\geq N^{\epsilon}$, and $\delta\in(0,\epsilon/10)$, then
-$0\ \leq\ \frac{2N^{2\delta}}{\sqrt{k(N)}}$ $\ \leq\ \frac{2N^{\epsilon/5}}{N^{\epsilon/2}}\ =\ 2N^{-3\epsilon/10}.$ (A.12)
+$0\ \leq\ \frac{2N^{2\delta}}{\sqrt{k(N)}}$ $\ \leq\ \frac{2N^{\epsilon/5}}{N^{\epsilon/2}}\ =\ 2N^{-3\epsilon/10}.$ (163)
 Hence
 $\left|\sum_{j=1}^{N^{\delta}}\binom{N^{\delta}}{j}(-1)^{j}\left(\frac{4\ell N^{\delta}}{k(N)+2\ell N^{\delta}}\right)^{j}\right|$ $\ \leq\ 2N^{-3\epsilon/10}\cdot\frac{1}{1-2N^{-3\epsilon/10}}$
-$\ =\ O\left(N^{-3\epsilon/10}\right).$ (A.13)
+$\ =\ O\left(N^{-3\epsilon/10}\right).$ (164)
 Similarly, for the left sum in (A), we also have
-$\left|\sum_{j=1}^{N^{\delta}}\binom{N^{\delta}}{j}(-1)^{j}\left(\frac{4(\ell+1)N^{\delta}}{k(N)+2(\ell+1)N^{\delta}}\right)^{j}\right|\ =\ O\left(N^{-3\epsilon/10}\right).$ (A.14)
-Applying (A) and (A.14) to (A), we get
-$\left|\alpha_{\ell,N}-1\right|\ =\ O\left(N^{-3\epsilon/10}\right).$ (A.15)
-Thus, substituting the estimate (A.15) back to (A) gives us
-$\left|\binom{k(N)}{k_{1,\ell}}-\binom{k(N)}{k_{1,\ell+1}}\right|\ \leq\ O\left(\binom{k(N)}{k_{1,\ell}}\cdot N^{-3\epsilon/10}\right).$ (A.16)
+$\left|\sum_{j=1}^{N^{\delta}}\binom{N^{\delta}}{j}(-1)^{j}\left(\frac{4(\ell+1)N^{\delta}}{k(N)+2(\ell+1)N^{\delta}}\right)^{j}\right|\ =\ O\left(N^{-3\epsilon/10}\right).$ (165)
+Applying (A) and (165) to (A), we get
+$\left|\alpha_{\ell,N}-1\right|\ =\ O\left(N^{-3\epsilon/10}\right).$ (166)
+Thus, substituting the estimate (166) back to (A) gives us
+$\left|\binom{k(N)}{k_{1,\ell}}-\binom{k(N)}{k_{1,\ell+1}}\right|\ \leq\ O\left(\binom{k(N)}{k_{1,\ell}}\cdot N^{-3\epsilon/10}\right).$ (167)
 Appendix B Case for $k_{S}<d$
@@
 Recall that $S$ is a proper subset of $[d]$ and $k_{S}$ is the largest index such that $k\in[d]\setminus S$, and that $D_{1}(y)$ is defined in (3.4) to be
-$D_{1}(y)$ $\ :=\ \sum_{\begin{subarray}{c}S\subset[d]\\ S\neq[d]\end{subarray}}\int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\Bigg{|}_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}$
-$\quad\quad dz_{m-1}\cdots dz_{m+1-d}.$ (B.1)
+$D_{1}(y)$ $\ :=\ \sum_{S\subsetneq[d]}\int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)$
+$\quad\quad\cdot\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\Bigg|_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}dz_{m-1}\cdots dz_{m+1-d}.$ (168)
 Let $\ell_{1};\ell_{2}]$ denote $\{\ell_{1},\dots,\ell_{2}\}$ if $\ell_{1},\ell_{2}$ are integers such that $\ell_{1}\leq\ell_{2}$, and let it be $\emptyset$ otherwise.
@@
 $\quad\ =\ \left(\prod_{i\in S\setminus[k_{S}+1;d]}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus(S\cup\{k_{S}\})}A(z_{m+i-d})\right)\cdot A(z_{m+k_{S}-d})$
-$\quad\quad\quad\cdot\left(\prod_{i\in[k_{S}+1;d-1]}\varphi(z_{m+i-d})\right)\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right).$ (B.2)
+$\quad\quad\quad\cdot\left(\prod_{i\in[k_{S}+1;d-1]}\varphi(z_{m+i-d})\right)\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right).$ (169)
 Hence,
-$\int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\Bigg{|}_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}$
+$\int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\Bigg|_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}$
 $dz_{m-1}\cdots dz_{m+1-d}$
-$\quad\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S\setminus[k_{S}+1;d]}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus(S\cup\{k_{S}\})}A(z_{m+i-d})\right)$
-$\quad\quad\quad\cdot A(z_{m+k_{S}-d})\cdot\left(\prod_{i=k_{S}+1}^{d-1}\varphi(z_{m+i-d})\right)\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\cdots dz_{m+1-d}.$ (B.3)
-To bound (B), we first give some estimate on the following
+$\quad\ =\ \int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S\setminus[k_{S}+1;d]}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus(S\cup\{k_{S}\})}A(z_{m+i-d})\right)$
+$\quad\quad\quad\cdot A(z_{m+k_{S}-d})\cdot\left(\prod_{i=k_{S}+1}^{d-1}\varphi(z_{m+i-d})\right)\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\cdots dz_{m+1-d}.$ (170)
+To bound (B), we first give an estimate on the following
 $\prod_{j=k_{S}+1}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i=k_{S}+1}^{d-1}\varphi(z_{m+i-d})\right)\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\cdots dz_{m+(k_{S}+1)-d}$
@@
 $\quad\quad\quad\cdots dz_{m+(k_{S}+1)-d}$
-$\quad\ \ll\ \frac{1}{\sqrt{2\pi(d-k_{S})}}e^{-\frac{\left(y-\sum_{i=1}^{k_{S}}z_{m+i-d}\right)^{2}}{2(d-k_{S})}},$ (B.4)
+$\quad\ \ll\ \frac{1}{\sqrt{2\pi(d-k_{S})}}e^{-\frac{\left(y-\sum_{i=1}^{k_{S}}z_{m+i-d}\right)^{2}}{2(d-k_{S})}},$ (171)
 where in the last line we use the fact that the second last line is exactly the convolution of $d-k_{S}$ standard Gaussian density function evaluated at $y-\sum_{i=1}^{k_{S}}z_{m+i-d}$, which is exactly the probability density function of sum of $d-k_{S}$ independent standard Gaussian random variables and thus is itself also Gaussian with mean 0 and variance $d-k_{S}$, evaluated at $y-\sum_{i=1}^{k_{S}}z_{m+i-d}$.
 Hence, (B) is bounded above by
-$\int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{k_{S}-1}\int_{z_{m}+(j-1)-d}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S\setminus[k_{S}+1;d]}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus(S\cup\{k_{S}\})}A(z_{m+i-d})\right)$
+$\int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{k_{S}-1}\int_{z_{m}+(j-1)-d}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S\setminus[k_{S}+1;d]}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus(S\cup\{k_{S}\})}A(z_{m+i-d})\right)$
 $\quad\quad\int_{z_{m}+(k_{S}-1)-d}^{\frac{y-\sum_{i=1}^{k_{S}-1}z_{m+i-d}}{d-k_{S}+1}}A(z_{m+k_{S}-d})\cdot\frac{1}{\sqrt{2\pi(d-k_{S})}}e^{-\frac{\left(y-\sum_{i=1}^{k_{S}}z_{m+i-d}\right)^{2}}{2(d-k_{S})}}dz_{m+k_{S}-d}\cdots dz_{m+1-d}$
-$\quad\ \ll\ N^{-\frac{1}{2}-\delta}\underbrace{\int_{-C^{\prime}\sqrt{N}}^{C^{\prime}\sqrt{N}}\cdots\int_{-C^{\prime}\sqrt{N}}^{C^{\prime}\sqrt{N}}}_{k_{S}-1}\left(\prod_{i\in S\setminus[k_{S}+1;d]}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus(S\cup\{k_{S}\})}A(z_{m+i-d})\right)$
+$\quad\ \ll\ N^{-1/2-\delta}\underbrace{\int_{-CN^{\delta^{\prime}}}^{CN^{\delta^{\prime}}}\cdots\int_{-CN^{\delta^{\prime}}}^{CN^{\delta^{\prime}}}}_{k_{S}-1}\left(\prod_{i\in S\setminus[k_{S}+1;d]}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus(S\cup\{k_{S}\})}A(z_{m+i-d})\right)$
 $\quad\quad\quad dz_{m+(k_{S}-1)-d}\cdots dz_{m+1-d}$
-$\quad\ \ll\ N^{-\frac{1}{2}-\delta}\prod_{i\in S\setminus[k_{S}+1;d]}\left(\int_{-C^{\prime}\sqrt{N}}^{C^{\prime}\sqrt{N}}\varphi(z_{m+i-d})dz_{m+i-d}\right)\cdot\prod_{i\in[d]\setminus(S\cup\{k_{S}\})}\left(\int_{-C^{\prime}\sqrt{N}}^{C^{\prime}\sqrt{N}}A(z_{m+i-d})dz_{m+i-d}\right)$
-$\quad\ \ll\ N^{-1/2-\left(1+\left|[d]\setminus(S\cup\{k_{S}\})\right|\right)\delta}\ \ll\ N^{-1/2-\delta},$ (B.5)
+$\quad\ \ll\ N^{-1/2-\delta}\prod_{i\in S\setminus[k_{S}+1;d]}\left(\int_{-CN^{\delta^{\prime}}}^{CN^{\delta^{\prime}}}\varphi(z_{m+i-d})dz_{m+i-d}\right)$
+$\quad\quad\quad\cdot\prod_{i\in[d]\setminus(S\cup\{k_{S}\})}\left(\int_{-CN^{\delta^{\prime}}}^{CN^{\delta^{\prime}}}A(z_{m+i-d})dz_{m+i-d}\right)$
+$\quad\ \ll\ N^{-1/2-\delta+\left|[d]\setminus(S\cup\{k_{S}\})\right|(-1/2-\delta+\delta^{\prime})}\ \ll\ N^{-1/2-\delta},$ (172)
 where in the second line, we use the change of variable $z_{m+k_{S}-d}\mapsto z_{m+k_{S}-d}-(\sum_{i=1}^{k_{S}-1}z_{m+i-d}+y)$ and the fact that
 […]
-Since the number of proper subsets $S$ of $[d]$ is finite, then when $k_{S}<d$, we have that $D_{1}(y)\ll N^{-1/2-\delta}$ by definition of $D_{1}(y)$ in (B).
+Since there is only a finite number of proper subsets $S$ of $[d]$, then when $k_{S}<d$, we have that $D_{1}(y)\ll N^{-1/2-\delta}$ by definition of $D_{1}(y)$ in (B).
 Appendix C Case for $d>2$
-In this section, we want to obtain the following estimate on $d^{2}\mathcal{M}_{Y^{(N)}}(y)/dy^{2}$ when $d>2$
-$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)$
+In this section, we want to obtain the following estimate on $dI_{N}(y)/dy$ when $d>2$
+$\frac{d}{dy}I_{N}(y)$
 $\quad\ \ll\ \underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)dz_{m-2}\cdots dz_{m+1-d}$
-$\quad\quad+\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)dz_{m-2}\cdots dz_{m+1-d}.$ (C.1)
+$\quad\quad\quad+\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)dz_{m-2}\cdots dz_{m+1-d}.$ (174)
 First, recall from (3.6) that
-$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)$
-$\quad\ \ll\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-2}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\Phi(z_{m+1-d})\right)^{m-d}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\right)^{2}$
-$\quad\quad+\Bigg{(}\int_{z_{m-2}}^{\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}}(\Phi(z_{m+1-d}))^{m-d}\left(\prod_{i=1}^{d-1}\varphi(z_{m+i-d})\right)(-1)\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
-$\quad\quad\quad\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\Bigg{)}dz_{m-2}\cdots dz_{m+1-d}.$ (C.2)
-Since $y$ is an upper bound on $\sum_{i=1}^{d}z_{m+i-d}$ and $z_{m-2}\leq z_{m-1},z_{m}$, then we have
-$z_{m-2}\ \leq\ \frac{z_{m-1}+z_{m}}{2}\ \leq\ \frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}.$ (C.3)
-When $(y-\sum_{i=1}^{d-2}z_{m+i-d})/2\leq 0$, $\varphi(z_{m-1})$ is at most $\varphi((y-\sum_{i=1}^{d-2}z_{m+i-d})/2)$. Hence,
+$\frac{d}{dy}I_{N}(y)$
+$\quad\ \ll\ \int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-2}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\Phi(z_{m+1-d})\right)^{m-d}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\right)^{2}$
+$\quad\quad\quad+\Bigg(\int_{z_{m-2}}^{\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}}(\Phi(z_{m+1-d}))^{m-d}\left(\prod_{i=1}^{d-1}\varphi(z_{m+i-d})\right)(-1)\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
+$\quad\quad\quad\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\Bigg)dz_{m-2}\cdots dz_{m+1-d}.$ (175)
+Since $y$ is an upper bound on $\sum_{i=1}^{d}z_{m+i-d}$ and $z_{m-2}\leq z_{m-1}\leq z_{m}$, then we have $z_{m-2}\leq z_{m-1}\leq(z_{m-1}+z_{m})/2\leq(y-\sum_{i=1}^{d-2}z_{m+i-d})/2$.
+When $(y-\sum_{i=1}^{d-2}z_{m+i-d})/2\leq 0$, $\varphi(z_{m-1})$ is at most $\varphi((y-\sum_{i=1}^{d-2}z_{m+i-d})/2)$.
+Hence,
 $\int_{z_{m-2}}^{\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}}(\Phi(z_{m+1-d}))^{m-d}\left(\prod_{i=1}^{d-1}\varphi(z_{m+i-d})\right)(-1)\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
@@
 $\quad\quad\quad\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}$
-$\quad\ll\ \left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\int_{-\infty}^{\infty}|z_{m-1}|\varphi(z_{m-1})dz_{m-1},$ (C.4)
+$\quad\ll\ \left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\int_{-\infty}^{\infty}|z_{m-1}|\varphi(z_{m-1})dz_{m-1},$ (176)
 where in the last line we use the change of variable $z_{m-1}\mapsto-z_{m-1}+y-\sum_{i=1}^{d-2}z_{m+i-d}$.
-The integral $\int_{-\infty}^{\infty}|x|\varphi(x)dx$ is the expected value of the absolute value of a standard normal random variable, which is finite.
+The integral $\int_{-\infty}^{\infty}|x|\varphi(x)dx$ is the expected value of the absolute value of a standard Gaussian random variable, which is finite.
 Hence, when $(y-\sum_{i=1}^{d-2}z_{m+i-d})/2\leq 0$,
@@
 $\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}$
-$\quad\ \ll\ \left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right).$ (C.5)
+$\quad\ \ll\ \left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right).$ (177)
 On the other hand, when $(y-\sum_{i=1}^{d-2}z_{m+i-d})/2\geq 0$,
-$y-\sum_{i=1}^{d-1}z_{m+i-d}\ \geq\ y-\sum_{i=1}^{d-2}z_{m+i-d}-\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\ =\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\ \geq\ 0.$ (C.6)
+$y-\sum_{i=1}^{d-1}z_{m+i-d}\ \geq\ y-\sum_{i=1}^{d-2}z_{m+i-d}-\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\ =\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\ \geq\ 0.$ (178)
 Hence, we find
@@
 $\quad\ =\ \left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)\Bigg|_{z_{m-1}=z_{m-2}}^{z_{m-1}=\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}}$
-$\quad\ =\ \left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)-\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)\right).$ (C.7)
+$\quad\ =\ \left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)-\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)\right).$ (179)
 Thus, if $d>2$, regardless of whether $(y-\sum_{i=1}^{d-2}z_{m+i-d})/2\leq 0$ or $(y-\sum_{i=1}^{d-2}z_{m+i-d})/2\geq 0$, we have by combining (C) and (C) that
@@
 $\quad\cdot\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}$
-$\quad\ \ll\ \left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)+\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)\right).$ (C.8)
+$\quad\ \ll\ \left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)+\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)\right).$ (180)
 Hence, (3.6) becomes
-$\frac{d^{2}}{dy^{2}}\mathcal{M}_{Y^{(N)}}(y)$
-$\quad\ \ll\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-2}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\right)^{2}$
+$\frac{d}{dy}I_{N}(y)$
+$\quad\ \ll\ \int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-2}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)\right)^{2}$
 $\quad\quad\quad+\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\left(\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)+\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)\right)$
@@
 $\quad\ \ll\ \underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(\frac{y-\sum_{i=1}^{d-2}z_{m+i-d}}{2}\right)dz_{m-2}\cdots dz_{m+1-d}$
-$\quad\quad+\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)dz_{m-2}\cdots dz_{m+1-d},$ (C.9)
+$\quad\quad+\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}\left(\prod_{i=1}^{d-2}\varphi(z_{m+i-d})\right)\varphi\left(y-\left(\sum_{i=1}^{d-2}z_{m+i-d}\right)-z_{m-2}\right)dz_{m-2}\cdots dz_{m+1-d},$ (181)
 which is exactly the estimate we need.